### v1

# Let’s Get to the Point: LLM-Supported Planning, Drafting, and Revising of Research-Paper Blog Posts

Abstract.
Research-paper blog posts help scientists disseminate their work to a larger audience, but translating papers into this format requires substantial additional effort. Blog post creation is not simply transforming a long-form article into a short output, as studied in most prior work on human-AI summarization. In contrast, blog posts are typically full-length articles that require a combination of strategic planning grounded in the source document, well-organized drafting, and thoughtful revisions. Can tools powered by large language models (LLMs) assist scientists in writing research-paper blog posts? To investigate this question, we conducted a formative study (N=6) to understand the main challenges of writing such blog posts with an LLM: high interaction costs for 1) reviewing and utilizing the paper content and 2) recurrent sub-tasks of generating and modifying the long-form output. To address these challenges, we developed Papers-to-Posts, an LLM-powered tool that implements a new Plan-Draft-Revise workflow, which 1) leverages an LLM to generate bullet points from the full paper to help users find and select content to include (Plan) and 2) provides default yet customizable LLM instructions for generating and modifying text (Draft, Revise). Through a within-subjects lab study (N=20) and between-subjects deployment study (N=37 blog posts, 26 participants) in which participants wrote blog posts about their papers, we compared Papers-to-Posts to a strong baseline tool that provides an LLM-generated draft and access to free-form LLM prompting. Results show that Papers-to-Posts helped researchers to 1) write significantly more satisfying blog posts and make significantly more changes to their blog posts in a fixed amount of time without a significant change in cognitive load (lab) and 2) make more changes to their blog posts for a fixed number of writing actions (deployment).
1. Introduction
Sharing scientific ideas outside of research papers is important for increasing awareness and impact (Jucan and Jucan, 2014; Burns et al., 2003). Scientists often seek different ways to communicate their work to broader audiences, such as through tweets or blog posts (Gero et al., 2021; Long et al., 2023; Williams et al., 2022; Holmberg and Thelwall, 2014); indeed, tweeting about research papers has been shown to raise their citation count (Luc et al., 2021). While some tools have been created to help author short-form derivatives like tweets (Long et al., 2023), significant challenges remain for scientists seeking to create longer-form derivatives like blog posts, which require more extensive planning, drafting, and revising writing processes (Flower and Hayes, 1981). In the planning step, the blog post author must review the paper, itself a long document, to find and select several pieces of information to include in the blog post. In the drafting and revising steps, there are many sub-tasks for completing the full blog post, such as creating an engaging hook or reporting results with sufficient detail but in simpler terms than the paper.
Integrating large language models (LLMs) in writing tools offers a promising avenue for improving writing outcomes; however, existing approaches are not designed to help create long-form derivatives like blog posts from long documents like academic papers. Prior work in human-LLM text summarization has focused on supporting short-form summaries, a sentence or paragraph in length (Zhang et al., 2023b; Lai et al., 2022). In the broader area of human-LLM writing, prior work has looked into providing assistance for writing long-form creative texts such as stories (Chung et al., 2022; Goldfarb-Tarrant et al., 2019; Singh et al., 2022) and screenplays (Mirowski et al., 2023), but these tasks do not take a long-form text input. Prior work in human-LLM scientific writing has investigated how LLM tools can assist scientists in writing short-form texts such as tweetorials (Long et al., 2023) and extended metaphors (Kim et al., 2023), but to the best of our knowledge, no work has sought to develop an LLM tool for creating long-form scientific texts, such as research-paper blog posts.
Still, prior work has shown that LLM tools can support planning, drafting, and revising for different forms of writing, such as creative stories (Goldfarb-Tarrant et al., 2019), short summaries (Zhang et al., 2023b), and short scientific texts (Long et al., 2023). Thus, in this paper we ask: how best can an LLM tool provide assistance in dealing with the challenges of writing long-form summaries of scientific long documents? We designed a technology probe of a simple writing interface with preliminary LLM-based features to better understand the costs and benefits of interacting with LLMs for the purpose of planning, drafting, and revising a research-paper blog post. The technology probe provides the paper text, an LLM-generated bullet point outline for the paper, a drafting space to write the blog post in sections, and access to an LLM prompting interface for writing tasks. Through a formative study (N=6), we gathered feedback on how users wish to leverage LLMs in this context and common interaction challenges. For one, the participants valued having a paper outline for finding and selecting relevant information from the paper to include in the blog post, but they struggled with understanding and utilizing it (e.g., understanding how the bullet points were connected to the paper text). Also, participants desired to provide the LLM recurring instructions for generating and modifying text that were not well-supported by the tool (e.g., rewrite a paragraph to better capture readers’ attention, continue writing from a particular sentence).
Based on the identified challenges, we iterated on the technology probe to develop the prototype tool Papers-to-Posts for LLM-assisted paper blog post writing. The tool follows our Plan-Draft-Revise workflow (Figure 1). In the Plan phase, the user may control text generation by selecting a subset of LLM-generated bullet points corresponding to key points in the source; Papers-to-Posts is the first LLM-powered long-document summarization tool to support direct control of content selection through a source outline, maintaining the source’s narrative structure. In the Draft phase, the user may supply additional instructions and preview LLM generations that use the selected bullet points, instructions, and blog post text thus far. In the Revise phase, they may use an LLM to further refine text with macros that encapsulate frequent editing transformations (e.g., to have less jargon, to be more dramatic).
We evaluated Papers-to-Posts through two studies in which participants wrote blog posts for their own papers, and we compared our tool to a baseline tool that provided an LLM-generated draft and integrated prompting. In a within-subjects lab study (N=20) in which researchers wrote blog posts for two of their own papers, one per tool, we found that, by helping authors to plan, draft, and revise their blog post, Papers-to-Posts led to significantly higher writer satisfaction with the blog post output alongside significantly greater change to the provided draft, without increasing task completion time or cognitive load. To study Papers-to-Posts in a more realistic setting, we also conducted an unmonitored, between-subjects deployment study (N=37), in which participants using Papers-to-Posts again made greater change to the provided drafts for a given amount of time spent or number of writing actions. Furthermore, through quantitative survey ratings and behavioral analysis as well as qualitative short answers and semi-structured interviews across the two studies, we observe that an LLM-generated paper outline and affordances for streamlining common long-form text generation and modification sub-tasks contributed to this increased satisfaction. In discussing our results, we also reflect on design implications for other human-LLM scientific long-document summarization tools with long-form output.
In summary, we make the following main contributions:
A formative study (N=6) producing two primary design goals for an LLM-powered tool for writing research-paper blog posts: (1) help users review and select content from a long source document during the planning stage, and (2) transparently streamline the process of instructing the LLM to generate and modify text during the drafting and revising stages.
Papers-to-Posts, a prototype tool for writing research-paper blog posts, which addresses the design goals identified in the formative study and implements the Plan-Draft-Revise workflow (Figure 1).
A within-subjects lab study (N=20 participants) providing evidence that Papers-to-Posts improves upon a powerful baseline (LLM draft and free-form prompting). Papers-to-Posts led to significantly higher writer satisfaction with the final blog post (with comparable task completion time and cognitive load) alongside significantly greater changes to the initial draft.
A field deployment study (N=37 blog posts, 26 participants) comparing Papers-to-Posts to the same strong baseline show that the same benefits hold in realistic scenarios where participants were writing blog posts on their own and sharing them publicly at the end of the study.
Design implications for other human-LLM scientific long-document summarization tools with long-form output: avoid solely post-editing interactions, present long-document outlines to support content selection, and include transparent affordances to streamline sub-tasks in generating and modifying long-form output text.
2. Related Work
2.1. Human-LLM Summarization
Although there is much work on automatic text summarization (Koh et al., 2022; El-Kassas et al., 2021), issues remain with the output summaries, such as inaccuracies and style misalignment (Lu et al., 2023; Koh et al., 2022). To address these issues, other works have looked into having a human and LLM work together to generate a summary (Lai et al., 2022; Moramarco et al., 2021). According to Cheng et al., there are five common methods of human-AI summarization: guiding AI output, selecting or rating AI output, post-editing of AI-generated summaries, interactive editing, and writing with AI assistance (Cheng et al., 2022). Our tool Papers-to-Posts provides affordances related to guiding and post-editing AI output.
Prior work in human-AI summarization has mainly focused on direct model guidance (e.g., utilizing concepts/queries (Zhang et al., 2023b; Bayatmakou et al., 2022; Avinesh and Meyer, 2017; Avinesh et al., 2018; Ding et al., 2023; Wang et al., 2023b), typing (Xie et al., 2023), ratings (Gao et al., 2020; Bohn and Ling, 2021), gaze (Bohn and Ling, 2021)) or simply post-editing (e.g., (Lai et al., 2022; Moramarco et al., 2021)). A few works have supported summarization based on user-selected points from the source (Bayatmakou et al., 2022; Yan et al., 2011; Zhang et al., 2011; Wang et al., 2023a, b), but the points are ranked or clustered by the system without the same narrative structure as the source document. Another work proposed but did not implement or evaluate model guidance through in-context content selection (Gehrmann et al., 2019). Our approach allows model guidance directly grounded in the input source while maintaining the source narrative structure.
Providing more direct control of input content is useful when the writer has significant knowledge of the source and can use reminders of its contents to select what to include in the summary. We focus on summarizing one’s own content, unlike most prior work, with the exception of work on summarizing medical notes that used the post-editing technique (Moramarco et al., 2021; Cai et al., 2022). Thus, we take advantage of the writer’s knowledge of the source document by having them select content in context of the source’s layout.
There are several works on human-LLM summarization of long documents (Zhang et al., 2011; Gehrmann et al., 2019) and some for scientific long documents in particular (Zarinbal et al., 2019; Bayatmakou et al., 2022; Zhang et al., 2023b). Long documents provide a unique challenge for human-LLM summarization, as there is a great deal of information for the human to digest. As writing research-paper blog posts entails summarizing portions of a research paper, this work seeks to support human-LLM long-document summarization.
2.2. Human-LLM Writing
While LLMs have impressive capabilities with respect to quickly generating large amounts of text, interacting with swaths of text can be difficult to manage without additional affordances. Popular chatbot LLMs such as chatGPT (OpenAI, [n. d.]) and BARD (Google, [n. d.]) offer some options for the user to guide their output beyond general prompting. Both provide example inputs for the user to select. chatGPT includes the option to provide custom instructions for the questions ”What would you like chatGPT to know about you to provide better responses?” and ”How would you like chatGPT to respond?”. BARD allows the user to modify its generated text to be shorter, longer, simpler, more casual, more professional, or modified based on a custom prompt. Tools from prior research also provide users with means to guide LLM writing beyond general prompting, such as keyword specifications (Zhang et al., 2023b), style specifications (Ghazvininejad et al., 2017; Coenen et al., 2021), and input modification (Zhang et al., 2011).
With respect to planning text, a unique aspect of our work is providing an LLM-generated outline for source content selection. There are a couple works in human-LLM writing that have utilized outlines, but they focused on outlining output content rather than input content. Zhang et al. introduce the tool VISAR, a human-AI argumentative writing assistant that allows users to iterate on their draft with adjustments to a visual outline of the draft (Zhang et al., 2023a). Dang et al. introduce a text editor that generates summaries of each paragraph written in the editor to comprise a reverse outline of the writing (Dang et al., 2022). The summaries then become affordances to help writers revise their writing (e.g., restructuring the writing). Our work also generates a reverse outline but as an affordance to guide summary generation based on the input content. For summarizing a long source document, an outline is particularly useful, as reviewing the entire source may require substantial time and effort.
With respect to revising text, like BARD (Google, [n. d.]) and other LLM-powered writing tools (NotionAI, [n. d.]), our final tool allows writers to utilize preset modifications, but some of these are set to be more specific to blog post writing. Closely related to our work, Laban et al. present an editing interface to help users revise their writing with LLM edit suggestions that can be integrated into their writing (Laban et al., 2023). This editing interface focuses solely on the revising step and on smaller edits than our tool.
2.3. Human-LLM Scientific Writing
Several works have investigated how to support human-LLM writing. Most of these works focus on tools for creative writing (Singh et al., 2022; Chung et al., 2022; Mirowski et al., 2023), but a few cater to scientific writing. A few works have investigated human-LLM scientific writing for broader science communication tasks that do not involve a source document. Gero et al. present a system for generating scientific ”sparks”, or inspiring sentences (Gero et al., 2022). Long et al. examine how LLM scaffolding can help people generate relatable hooks for complex scientific topics (Long et al., 2023), and Kim et al. investigate how science writers can generate extended metaphors for scientific ideas with the help of an LLM (Kim et al., 2023). The couple works that have explored human-AI scientific summarization do not allow for in-context selection of source content but rely on selection of concepts (Zhang et al., 2023b) or out-of-context sentences extracted based on queries (Bayatmakou et al., 2022).
These previously studied forms of human-LLM scientific writing are the length of a sentence or paragraph. Based on prior work (Mehlenbacher, 2019), popular resources (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]; Medium, [n. d.]), and feedback from three communications experts, we determined a set of guidelines for writing research-paper blog posts, which indicate that the shortest blog posts are recommended to be around 500 words in length (Medium, [n. d.]; Kudos, [n. d.]; Patrick Dunleavy, [n. d.]), making blog posts a long-form output. Because research-paper blog posts are longer than a sentence or a paragraph, they require coordinating more writing sub-tasks, such as rewriting a sentence about a method to be easier for a broader audience to understand or summarizing a paper’s results section.
2.4. Scientific Blog Posts
Scientists can engage with broad community through various channels, including press releases, magazines, journals, and tweets (August et al., 2020; Gero et al., 2022). This work focuses on scientific blog posts, which offers increased recognition and transparency of one’s work, writing practice, editorial freedom, and public engagement (Jarreau, 2015; Blanchard, 2011; Zou and Hyland, 2019). There are several forms of scientific blog posts, including academic commentary, free commentary, political commentary, and mediation of research to laypersons (Mahrt and Puschmann, 2014). Here, we focus on blog posts with the primary goal of communicating information from a research paper to scientists in the broad area of the paper (e.g., computer science, psychology).
3. Formative Study and System Iteration
We conducted a formative study to identify challenges that people encounter in working with an LLM-powered technology probe for writing blog posts about their research papers, and corresponding design goals for our final LLM tool.
3.1. Participants
Via academic social networks, we recruited 6 participants for the formative study and compensated them with $25 for the 45-minute study. Participants were PhD students, a population frequently tasked with writing blog posts for papers they have authored. Five worked in computer science and one in information science. To circumvent initial time required for getting used to LLMs themselves, we recruited people who already had some experience of working with LLMs in general (6-10 sittings: 1; ¿10 sittings: 5). We also required participants to have previously written research papers (3-10 papers: 3; ¿10 papers: 3), since the purpose of such a tool is to write blog posts for one’s own research papers. Participants had a mixture of experience with writing research-paper blog posts (0 posts: 4; 1-2 posts: 2). Also, for sufficient participant motivation, we made sure that all participants were more interested than not (at least 5 on a 7-point Likert scale where 7 indicates the highest interest) in writing blog posts for their research papers.
3.2. Study Procedure
The study consisted of 3 iterations, each of which involved sessions with 2 participants. Based on participants’ interactions and feedback, we updated the probe after each iteration for better insight into remaining challenges. Prior to each study session, the participant completed a participation agreement. The session coordinator recorded and transcribed each session in Google Hangouts. The coordinator asked the participant to assume their audience was laypeople, as a common advantage of writing scientific blog posts is to communicate with a broader audience than the paper does (Jarreau, 2015; Blanchard, 2011). The participant then used a link to access the technology probe initialized for one of their research papers that they provided. In order to see how participants would intuitively use the probe, the coordinator purposefully did not initially explain how to use it. For 30 minutes, the participant used the probe to write as much as possible for a blog post about their research paper. We asked participants to think aloud as they worked (Van Someren et al., 1994). The coordinator occasionally asked follow-up questions about what the participant said aloud. Participants could ask for help when they did not understand an aspect of the tool. Once 30 minutes had passed, the coordinator engaged the participant in a semi-structured interview. They asked the participant about the difficulties and benefits of using the probe to 1) navigate the paper and outline, 2) generate and edit text for a blog post section, and 3) edit the blog post layout. Finally, they asked for the participant’s thoughts on receiving different types of feedback on their blog post, but this topic was left for future work.
3.3. Initial Technology Probe
In the first iteration of the technology probe interface (Appendix 28), the left side consisted of blog post sections pre-populated with an LLM-generated draft for drafting and revising the blog post. The right side contained the paper text with selectable paragraphs and an LLM-generated paper outline with bullet points for planning. Under the currently selected blog post section, there was an LLM-supported text generator that contained buttons to generate section text from the selected paragraphs and bullet points as well as to expand and condense the text in the blog post section editor.
3.4. Design Goals and Tool Iteration
After each iteration of the formative study, the first author conducted inductive thematic analysis (Braun and Clarke, 2006) of the participants’ think-aloud and interview comments on their experience with the probe. Participants described difficulties in working with the probe that, aside from general usability issues, may be classified under two overarching challenges.
3.4.1. Challenge 1: Planning Paper Content to Use
An academic paper is a long document with dense information, so we expected that identifying information to include in the blog post would be a challenge. For this reason, our technology probe parsed the paper text and generated a paper outline with an LLM in order to support navigating and utilizing the paper. Still, participants faced difficulties with finding and selecting information from the paper. Our associated design goal was to help writers plan their writing by facilitating review of a long document’s contents and selection of essential information to include in the writing (DG1).
Although participants faced issues with the paper outline, four participants expressed that they liked having the paper outline, and four appreciated the ability to select paper paragraphs or bullet points from which to generate text. Because participants overall liked and used the paper outline for identifying and using essential information from the paper, we iterated on this affordance to address DG1. We identified the following requested functionality associated with this design goal:
Transparent Connection Between Paragraphs and Bullet Points. In the first iteration, P2 noted confusion around how the bullet points in the outline corresponded to the original paper text, despite the existence of a button to the left of each set of bullet points that revealed the associated paragraph when clicked. We therefore added a bright line to the left of each set of bullet points to better visualize its grouping and its connection to the corresponding paragraph (see Figure 2). In the second iteration, P3 expressed uncertainty about how the selection of paragraphs and bullet points were related. For example, they were surprised that a paragraph was selected but only one of three corresponding bullet points was selected. During the formative study, the selection logic was as follows. If a bullet point is selected, the associated paragraph needs to be selected, but if a paragraph is selected, no bullet point needs to be selected. This was done in an attempt to make it clearer that the associated paragraph was given to the model as context in addition to any bullet points selected. However, this logic seemed to confuse the participants.
Filter for Selected Content. In the third iteration, P6 said that they were not sure if they had selected everything that they wanted to select. Throughout all the iterations, the probe only communicated the number of selected paragraphs and bullet points. P6 desired a way to view only the selected paragraphs and bullet points.
3.4.2. Challenge 2: Drafting and Revising with Recurring LLM Instructions
The second challenge was attempting to provide the LLM instructions for generating and modifying text that were not well-supported by the probe. Some of these instructions were frequently desired, but the participant had no reasonable way to input them or, once directly prompting the LLM was available, had to determine a prompt to fit their needs. Participants also faced confusion regarding the prompts being used in the back-end. Our corresponding design goal was therefore to help writers draft and revise their writing by transparently streamlining commonly desired instructions to the LLM for generating and modifying text (DG2). We identified the following requested functionalities under this design goal:
Transparent Prompt Input. In the second iteration, P3 liked a default blog post section and was confused that it appeared significantly different from the new version they generated without adjusting the pre-selected bullet points. For one, there was a difference in the requested length in the prompt for the initial-draft section versus the live-generation section. Also, the participant was not aware of the section-specific instructions that were provided to the model for the initial-draft section. To provide access to default settings, which may often be desirable, we updated the probe to show these section-specific instructions as preset (but adjustable) custom instructions. While the model was always prompted to add 1 to 2 sentences in the third iteration, P5 was confused regarding if and how much the text would expand after clicking the ”expand” button multiple times. Meanwhile, P6 commented that they were unsure if the model had context for the blog post sections above the one for which they were generating text.
Custom Instructions. In the first iteration, our design probe did not support direct custom instructions. Users could only generate or modify the text using the pre-defined operations. P1 tried to work around this limitation by giving specific instructions to the LLM in the blog post section editor, custom bullet points, and custom instructions. The participant seemed unaware that the model was not taking these instructions into account. To provide users with more control over the generated text, we added a field for custom instructions to the generator box.
New Section Generation. In the second iteration, P4 noted wanting to generate entire new sections with the model, perhaps based on the section header.
Flexible Input for Text Modification. In the first iteration, P2 expected and wanted the ”condense” button to iterate on the generated text rather than the editor text, so we updated the ”expand” and ”condense” buttons to do that. However, in later iterations, P4 expressed interest in the model expanding on something they wrote without reference to the paper, and P6 expressed interest in rewriting portions of the original paper in different manners.
Preset Modification Actions. We observed participants attempting to rewrite text in manners that seemed applicable to many research-paper blog posts. For instance, P1 wanted a button to make text more approachable for their audience of laypeople, while P3 rewrote text that seemed ”too lofty” in its claims. For the third iteration, we therefore added buttons for such modifying actions, which included ”simpler terms”, ”less dramatic”, and ”more dramatic.” When the participant clicked the button, the modification would be applied to the last generation.
4. The Papers-to-Posts System
4.1. Paper Outline and Initial Draft Generation
The paper text was extracted from the paper PDF using a public tool for converting scientific PDFs to HTML.111https://papertohtml.org We manually adjusted the parsed HTML so that the section headers were correct, no section or subsection was completely missing, and only the main paper text was included (i.e., no appendix, references, or footnotes). Within the paper text box, participants could click a button to see a paper outline generated for the paper text. To obtain this paper outline, we used an LLM to generate 1 to 3 bullet points to summarize each paragraph in the paper, depending on the length of the paragraph. To generate the initial blog post draft for the tool, all of the paper’s bullet points were given to the model along with the request to select 10 bullet points relevant to the initial four blog post sections: introduction, methods, results, and conclusion. To generate each draft section, the model was provided the selected bullet points, their associated paragraphs, the portion of the blog post that had already been generated, and section-specific guidelines based on prior work (Mehlenbacher, 2019), popular resources (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]; Medium, [n. d.]), and feedback from three communications experts. We found that this process of generating one draft section at a time with the context of prior sections helped make the draft more coherent while avoiding overwhelming the model with instructions for many sections. For both the initial draft and user generations, the model was provided the associated paragraphs for any bullet points in order to ground the model in the actual paper. We used GPT-4 for all LLM tasks except for extracting bullet points from a paragraph, which we produced using GPT-3.5-Turbo after observing that it produced shorter bullet points that were quicker to read. It was also many times cheaper for this token-intensive generation task.
4.2. User Interface
Developed over the course of the formative study, the Papers-to-Posts system consists of two main panes split vertically. On the left is the pane for writing the blog post (Figure 2a), and on the right is the area for generating and modifying text. The blog post is divided into sections, which can be named, reorganized, and deleted. Each section is connected to a unique LLM state, so the history of all inputs and outputs for text generation and modification are stored separately for each section. The initial blog post draft is generated as described in Section 3.3. The blog post section currently being edited is highlighted in purple to indicate that it is the section for which text is generated. Using the controls on the top of the left pane, users can view the blog post’s word count, a preview of the full blog post, and tips for writing blog posts based on prior work (Mehlenbacher, 2019) and popular resources (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]; Medium, [n. d.]), as well as feedback from 3 communications experts. There is a button to submit the blog post at the bottom. The writer can also add a new section to the blog post using the plus-sign button next to each blog post section. When the writer clicks this button, they can either create a blank section or generate a section based on the section header they provide. To generate a new section, we prompt GPT-4 to select 10 bullet points for the section based on the provided header and to generate text for the section focused on those bullet points (Appendix A).
The pane for generating and modifying text on the right has three areas. The first is the planning-support area (Figure 2b), which contains the Paper Outline to address DG1. The Paper Outline supports writers in planning their blog post, through review of the paper’s contents and selection of essential information to include in the post. The paper view at the top of the pane offers the writer two views of the paper in two tabs: Paper Text and Paper Outline. Under the Paper Text tab, the writer may view the original paper text, divided into paragraphs. The paper is parsed into collapsible sections. The figures and tables are presented in a separate final section of the paper. Their captions can be copied for pasting in the editor so that the writer knows where they should go in the final blog post. Under the Paper Outline tab, the writer may view the paper outline, divided into sets of bullet points associated with the original text’s paragraphs. The writer may view a bullet point set’s corresponding paragraph by clicking the button to the left, and the reverse is true under the Paper Text tab. The outline is generated, and the pre-selected bullet points for each blog post section are displayed. The writer may select or deselect any bullet points or paragraphs that they want the LLM to use in generating text for the blog post section on which they are working. Selecting a paragraph does not affect the selection of associated bullet points, and vice versa. On the back-end, the model still receives the context of the associated paragraph for each bullet point selected. There are buttons below the paper view to clear all selected bullet points and paragraphs, filter for only selected bullet points and paragraphs, and close or open all the paper sections. When the filter button is activated, if the writer is in the Paper Text tab, if a paragraph was deselected but a corresponding bullet point is still selected, the tool notes that ”A corresponding bulletpoint(s) is still selected,” and vice versa.
Below the planning-support area is the drafting-support area (Figure 2c). There are several inputs to address DG2 by transparently streamlining commonly desired instructions to the LLM for generating text. First, there is a menu to select the desired length for the generated text. The default length is labeled as ”auto,” and the other options are 1 sentence, 1 paragraph, or a few paragraphs. The ”auto” length is the same as for the initial LLM draft sections (around 125 to 250 words) There is next a toggle turned on by default for the LLM to use the context of blog post sections above the one currently being edited. The writer can optionally utilize text fields to provide custom bullet points, custom instructions, and starting text to the LLM. For the initial introduction, methods, results, and conclusion sections of the blog post, the custom instructions field is pre-filled with the section-specific instructions used to generate those sections. The text generator’s ”generate” button utilizes the inputs to generate text for the current blog post section using GPT-4.
The tool also provides a revising-support area (Figure 3d, located below the drafting-support area in the tool), which addresses DG2 by transparently streamlining commonly desired instructions to the LLM for modifying text. First, there is a text field for inputting the text to be modified. This can be anything, including text from the blog post draft, the paper, or a generation. There is then a menu to select the desired length for the modified text. The options are the same as for generated text, with ”auto” being unique to each modifying action (see Appendix A). Next, there is a toggle, on by default, for the LLM to use the context of blog post sections above the one currently being edited. There is also a toggle turned on by default for the LLM to base modified writing on selected paragraphs and bullet points. Underneath the toggles, there is a text field for providing any custom instructions for the modification action. At the bottom, there are 5 preset modification buttons: ”expand”, ”condense”, ”simpler terms”, ”less dramatic”, and ”more dramatic”. There is also a ”custom modification” button. Aside from the ”expand” and ”condense” buttons, all of the affordances for generating and modifying text were additions made to support recurring types of LLM instructions observed in the formative study. In the lab study, participants utilized each affordance to some degree.
5. Lab Study
We conducted a lab study to determine if and how the LLM tool resulting from the formative study may support researchers in writing blog posts about their research papers.
5.1. Hypotheses
We hypothesized that, compared to a baseline consisting of an LLM draft and the ability to prompt the LLM, Papers-to-Posts scaffolding and affordances would provide participants a better experience. Our individual hypotheses were as follows:
H1: Papers-to-Posts leads to higher satisfaction with the output blog post than the baseline.
H2: Papers-to-Posts leads to higher satisfaction with the tool than the baseline.
H3: Papers-to-Posts leads to lower cognitive load than the baseline.
H4: Papers-to-Posts leads to lower task completion time than the baseline in generating a blog post that the writer would be comfortable posting publicly.
5.2. Participants
We recruited 20 participants (M: 14, W: 6) via academic social networks and institutional mailing lists and compensated them with $83.33 USD for 2.5 hours of their time. Participants were predominantly early career researchers (PhD student: 17, master’s student: 1, postdoc: 1, industry researcher: 1), who are often lead authors tasked with writing blog posts. All participants conducted research in computer science. Eight participants noted being in a subarea related to human-computer interaction. Participants had a broad range of experience in terms of prior LLM use (¿10 sittings: 14, 6-10 sittings: 3, 2-5 sittings: 2, ¡2 sittings: 1) and number of research-paper blog posts authored (0 posts: 14, 1-2 posts: 5, 3-10 posts: 1). In order to simulate a naturalistic setting, participants were required to have authored at least 2 research papers for use in the study (2 papers: 1, 3-10 papers: 14, ¿10 papers: 5) and to be more interested than not (at least 5 on a 7-point Likert scale where 7 indicates the highest interest) in writing blog posts for their research papers.
5.3. Study Conditions
The treatment condition involved interacting with Papers-to-Posts (Figure 2), which is described in Section 4. The tool in the baseline condition was designed to be a simplified version of Papers-to-Posts without affordances developed to address the design goals (Figure 4). The baseline provided input paper text, sourced from the PDF and divided into paragraphs and collapsible sections as in Papers-to-Posts. The baseline also provided the ability to prompt GPT-4 with any instructions. A history of generations and their associated instructions was retained for the user’s perusal. Lastly, the baseline had an editor for writing the blog post, which was pre-populated with an LLM draft generated by GPT-4-32k using the entirety of the input paper. GPT-4-32k was used rather than GPT-4 because the context window for GPT-4 was insufficient for longer research papers. The prompt was designed to mirror the prompt for Papers-to-Posts’s initial draft sections, so the resulting draft also had introduction, methods, results, and conclusion sections. However, the prompt did not contain section-specific instructions. The full prompt can be found in Appendix A.
In each condition, the tool was set up for one of the participant’s research papers submitted in the recruitment survey. The paper was different for each condition. Each paper was required to have been published online between 2018 and the study period in 2023, with the participant as an author. Only one paper was published in 2018, and the rest were published during or after 2020. There was no significant difference in publication dates between conditions (Paired-Samples t-Test, t(19)=0.90, p=n.s.). The papers were parsed as described in Section 3.3. The average paper length across conditions differed by less than 10% (baseline: 8452 words, treatment: 7742 words).
5.4. Study Procedure
The study was within-subjects, so each participant completed two study sessions, one in the treatment condition and one in the baseline condition, in randomized order and following a counter-balanced design. The participants’ selected papers were assigned randomly to conditions. Each study session lasted up to about 75 minutes. To provide sufficient flexibility for recruited participants, they can select two study sessions within a week of each other, but the majority of sessions were a day apart (1 day: 13, 2 days: 3, 3 days: 3, 4 days: 1). The condition order was counter-balanced within the group of participants with sessions 1 day apart and within the group of participants with sessions over 1 day apart.
Participants completed a participation agreement before the study. We recorded and transcribed each session in Google Hangouts. To start, the session coordinator introduced participants to the task of writing a blog post with an AI tool. The coordinator then presented a tutorial on the tool. After, participants had 2 minutes to interact with the tool in the context of a sample paper that was not their own, and they could ask questions about the tool. Participants next opened a tool link initialized with one of their research papers. The coordinator asked them to assume that their audience was people in their broad area, which was computer science for all participants. We made this assumption because it was the most common audience for whom potential recruits wanted to write blog posts (52 of 60), according to the recruitment survey. The coordinator instructed participants to work as if they were not there, but to let them know if any questions about the tool arose. They also instructed participants not to use any tools or information outside of the tool provided. We made an exception when participants wanted to view the PDF of their paper due to issues working with its parsing in the tool. This only occurred twice. Lastly, the coordinator asked participants to let them know if they reached a point at which they would feel comfortable publicly publishing the blog post in its current state.
Each participant had up to one hour to work on the blog post. At the 20-minute and 40-minute marks, the coordinator reminded them to let them know if they reached the point of comfort with publication. If and when the participant reached that point, the coordinator asked them to save their draft and continue revising the blog post. After the first four study sessions, we noticed that participants were sometimes done early and did not desire to revise further. Thus, we added a rule so that the participant would let the coordinator know if they reached a point at which they were proud of the blog post and saw no reason to revise further. If and when the participant reached that point, or if an hour had passed, they were asked to submit and moved on to the follow-up survey regarding their experience with the tool. The survey consisted of Likert-type questions regarding perceived efficiency, cognitive load, ease of using the tool in relation to the design goals, satisfaction with the tool, satisfaction with the final blog post, perceived control, and perceived ownership. There were also short answer questions regarding the difficulties and benefits of using the tool in relation to the design goals. If it was the participant’s second session, the survey ended with a few questions directly comparing the two conditions. These were not for analysis but only to help us get a quick sense of the results. However, they were not allowed to change their first-session responses. At the end of the session, the coordinator told participants to close out of the tool and, if it was their first session, not to use the tool before the next session. Each time the participant prompted the LLM, saved, or submitted, a log of the user action along with the full state of the blog post draft was saved for analysis.
5.5. Results
We evaluated our 4 hypotheses (Section 5.1) for statistical significance and found that blog post satisfaction is significantly different between the two conditions, while tool satisfaction, cognitive load, and difference in task completion time are not. We also investigated how well the design goals were fulfilled, and the extent of participants’ editing. We saw evidence that our tool addressed the design goals and increased participants’ satisfaction with control over the tool. In addition, we observed that participants did significantly more editing of the initial LLM draft using Papers-to-Posts (we note that this significance test was not part of our initial hypotheses and is therefore exploratory in nature).
5.5.1. Blog Post Satisfaction
We utilized the Likert responses (see Figure 5) as a measure of participants’ satisfaction with their final blog post under both conditions. Given the Likert nature of the data, we employed a non-parametric test. Furthermore, we used a sign test rather than a Wilcoxon signed rank test, which is normally used to analyze non-parametric within-subjects data for two conditions of one factor, due to the violation of that test’s assumption of a symmetric distribution of the paired differences about the median. We observed that participants were significantly more satisfied with the final blog post generated with Papers-to-Posts (M=6.00,Q1=5.00,Q3=7.00) as compared to the baseline tool (M=5.00,Q1=5.00,Q3=6.00) (Dependent-Samples Sign Test, S=11, p¡.05). The majority of participants (11) were more satisfied in the treatment condition, 7 were equally satisfied in both conditions, and only 2 were more satisfied in the baseline condition.
Inspecting participants’ Likert responses regarding aspects of their blog post satisfaction, we see that participants particularly demonstrated more satisfaction with the content covered in the Papers-to-Posts blog post than the baseline blog post (Figure 5b). The majority of participants (14) were more satisfied with the content under the treatment condition (equally: 2, less: 4). Overall, participants were more satisfied with the blog post structure using Papers-to-Posts than the baseline (Figure 5d). The plurality of participants (9) were more satisfied with the structure under the treatment (equally: 6, less: 5). Participants did not show increased satisfaction with the blog post’s tone under the treatment compared to the baseline (Figure 5c).
Looking at participants’ perceived task performance (Figure 6c), we observe that participants had higher perceived performance with Papers-to-Posts than with the baseline. The plurality of participants (9) thought they were more successful in writing a research-paper blog post with Papers-to-Posts compared to the baseline (equally: 7, less: 4). This further corroborates participants’ blog post satisfaction under the treatment condition.
5.5.2. Tool Satisfaction
For measuring tool satisfaction, we used the 5-question Likert scale on tool acceptance from Kocielnik et al. (Kocielnik et al., 2019) (Figure 5). The 5 responses were averaged to obtain one score. We used the non-parametric sign test to evaluate the data, as it was Likert data and violated the Wilcoxon signed rank test’s assumption of a symmetric distribution of the paired differences about the median. While participants had a higher median satisfaction with Papers-to-Posts (M=5.9,Q1=5.2,Q3=6.5) than the baseline tool (M=5.6,Q1=4.8,Q3=6.0), this difference was not significant (Dependent-Samples Sign Test, S=13, p=n.s.).
In Figure 5e-h, we see that, under the treatment condition, participants were more likely to find the tool helpful, use it if it were available, recommend it, and perceive it as useful for productivity. For each of these questions, at least 4 more participants increased rather than decreased their Likert rating from the baseline to the treatment condition. However, for the question about the tool inducing annoyance, only 2 more participants indicated decreased rather than increased annoyance with Papers-to-Posts.
5.5.3. Cognitive Load
To measure cognitive load, we used the NASA TLX Index questionnaire (Hart and Staveland, 1988) (Figure 6a-d) and summed the responses for the mental demand, temporal demand, effort, and frustration questions (see 5.5.1 for discussion of the responses for performance). We utilized the non-parametric sign test to evaluate the data, as it was Likert data and violated the Wilcoxon signed rank test’s assumption of a symmetric distribution of the paired differences about the median. We saw no significant difference between the baseline and treatment conditions, although participants reported slightly lower cognitive load with Papers-to-Posts (M=13.0,Q1=11.5,Q3=16.3) than the baseline tool (M=14.5,Q1=12.0, Q3=18.3) (Dependent-Samples Sign Test, S=8, p=n.s.). Figure 6a-d presents little difference in mental demand, temporal demand, and frustration between the treatment and baseline conditions. However, with respect to effort, we note a divergence. The majority of participants (11) reported using less effort with Papers-to-Posts than with the baseline tool, whereas only 4 participants reported using more effort.
5.5.4. Task Completion Time
We recorded two metrics for task completion time. One was the amount of time participants needed during the session before they would be comfortable publicly publishing the blog post. The other was the perceived amount of time participants would need after the session before they would be comfortable publicly publishing the blog post. As participants reached that point of comfort during most sessions, we focused on the former metric. Because this data was numeric within-subjects data for two conditions of one factor, we used a paired-samples t-test. Participants worked slightly faster with the baseline tool (M=44.3 minutes,SD=12.5) in comparison to Papers-to-Posts (M=46.0 minutes,SD=13.9), but this difference was not significant (Paired-Samples t-Test, t(19)=0.55, p=n.s.).
5.5.5. Control and Ownership
Participants indicated higher perceived control (treatment: M=5.00,Q1=4.50,Q3=6.00, baseline: M=5.00, Q1=3.00,Q3=6.00) and satisfaction with their control (treatment: M=5.00,Q1=2.75,Q3=6.00, baseline: M=5.00,Q1=3.00,Q3=5.25) under the treatment condition only if we look within-subjects (Figure 7a-b). Eleven participants had higher perceived control with Papers-to-Posts than the baseline tool (equal: 2, lower: 7), and 10 participants had higher satisfaction with Papers-to-Posts (equal: 4, lower: 6). On the other hand, participants did not demonstrate increased perceived ownership (treatment: M=6.00,Q1=5.75,Q3=6.00, baseline: M=6.00,Q1=5.00,Q3=6.00) or satisfaction with ownership (treatment: M=6,Q1=4.00,Q3=6.25, baseline: M=6.00,Q1=5.00,Q3=6.00) under the treatment condition (Figure 7c-d). Although they felt that they had more control over Papers-to-Posts, participants may not necessarily have felt that the LLM tool contributed more or less to the creation of the blog post.
5.5.6. Fulfillment of Design Goal 1
Looking at Figure 8a-b, we see that participants were more confident that all essential information was in the blog post (treatment: M=6.00,Q1=3.00,Q3=6.25, baseline: M=5.00,Q1=3.00,Q3=6.00) and found it easier to include essential information from the paper (treatment: M=6.00,Q1=5.00,Q3=6.00, baseline: M=5.00,Q1=4.75,Q3=6.00) using Papers-to-Posts. In addition, the interaction logs show that participants utilized Papers-to-Posts’s DG1-inspired affordance to select relevant paper bullet points and paragraphs to incorporate in the blog post. All participants adjusted the LLM’s pre-selected bullet points to some degree. On average, participants added or removed 28 bullet points and 13 paragraphs. We note that some added bullet points come from automatic selection of bullet points used to generate a new section rather than manual selection. Participants also provided an average of 1.70 custom bullet points.
In their short answers regarding aspects of Papers-to-Posts that they found difficult and helpful, 10 participants noted the paper outline being useful in some manner. Four participants described the outline as helpful for generating text. For instance, P15 noted, ”I could rewrite entire sections or direct it with just a few bullets.” Three participants mentioned that the outline was useful for reviewing the information in or structure of the paper. P4 commented, ”The outline was really helpful in providing a quick overview of what content I had written.” For two participants, the bullet points provided transparency around the origins of the initial LLM draft and generations. P6 explained, ”The checked-marked bullet points helped me understand where the pre-filled content on the left was from,” and P12 shared, ”I liked the bullet point features to see what was being fed in vs not.” Five participants also specified a difficulty related to the paper outline. Two participants were overwhelmed with the number of bullet points. P9 described the outline as ”really bloated and hard to navigate.” Three more participants were confused or encumbered by aspects of utilizing the bullet points. For example, P3 ”felt stressed by choosing between the entire paragraph at hand and the (presumably) AI-generated summary bullets,” and P20 did not realize that bullet points could come from different sections of the paper.
5.5.7. Fulfillment of Design Goal 2
Participants’ Likert responses regarding DG2 indicate that they found it easier to produce appropriate blog post text (treatment: M=6.00,Q1=5.00,Q3=6.00, baseline: M=5.00,Q1=4.00,Q3=6.00) and iterate on the blog post text (treatment: M=6.00,Q1=5.00,Q3=6.00, baseline: M=5.00,Q1=4.00,Q3=6.00) in the treatment condition (Figure 8c-d). Furthermore, the interaction logs for the treatment condition indicate that participants took advantage of DG2-based affordances for generating and modifying text for sub-tasks (Appendix 27). For an average of 6.85 generations per treatment condition, participants often utilized custom instructions (5.70 times), selected paragraphs or bullet points (6.35 times), and the context of prior blog post sections (6.00 times), though we note that these features were toggled on by default. In addition, they modified text 6.85 times on average, with the ”condense” (1.85 times), ”custom” (1.80 times), and ”simpler terms” (1.40 times) modifications being the most common.
In their short answers, 5 participants noted the usefulness of being able to provide custom instructions for either generating or modifying text using Papers-to-Posts. P9 described the custom instructions as providing a greater sense of control: ”…being able to include explicit instructions for the model to generate text from was helpful in being able to control the information in the text that it generated.” With respect to revising writing in particular, 12 participants noted in their short answers that Papers-to-Posts’s affordances for modifying text were useful. All of the preset modification buttons (”simpler terms”, ”condense”, ”expand”, ”more dramatic”, ”less dramatic”) were specifically mentioned as helpful affordances at least once. The most commonly noted one was the ”condense” button. For example, P12 commented, ”I sometimes put the modify text through multiple rounds of expand and contract, focusing on different parts. I found that helpful.” Two participants commented that they desired more preset modification buttons. P16 wanted a ”more technical” button, while P7 simply shared, ”I wish there are more default buttons for similar functions.” P7 then added, ”Or even let user to add their favorite rewrite button.” Interestingly, P2 found the preset modification buttons useful for understanding the LLM tool’s capabilities. They reflected, ”It was nice to know what options the tool already knew of, and would perhaps have success with. For example, I would probably not have gone to a blank-box tool and said, ’please make the following more dramatic’. I would have stuck to more basic things like ’please summarize the following’.”
Participants also discussed difficulties that they encountered in drafting and revising writing with Papers-to-Posts. Eight participants commented that they appreciated the simplicity of the baseline tool or that the baseline tool gave them more freedom over the instructions sent to the tool. For instance, P9 noted, ”the ’generate writing’ panel felt overwhelming in terms of features,” and P19 explained, ”It’s very open end so I felt like I got more freedom what to ask it to do.” Meanwhile, in the treatment condition, 4 participants expressed wanting to take specific actions that were not explicitly supported by Papers-to-Posts. As an example, 2 participants wanted assistance with checking for redundancies in the blog post. Another 2 wanted feedback on the blog post writing. Regarding the affordances for modifying text in particular, 8 participants noted a relative difficulty. A common issue, raised by 4 participants, was confusion about how or when to use certain affordances. As an example, P3 ”didn’t really understand if the custom instructions only applied for the ’custom modification’ option,” when it in fact would be applied to any modification action. Meanwhile, P6 was unsure of ”when to use which one of the six [modification] options.” Another recurrent issue noted by 5 participants was a modification affordance not working as desired. For P20, setting the desired length of the modified writing to one paragraph still resulted in three, as the back-end prompts are not guaranteed to work as planned. For P12, modifications were sometimes too exaggerated, while for P16, a custom modification to make writing more technical was too subdued. Despite these comments about feeling restricted or confused by Papers-to-Posts, participants still reported higher perceived control and satisfaction with control in the treatment condition, which we discuss in Section 5.5.5.
5.5.8. Extent of Change to Blog Post
Interestingly, we found that participants made larger changes to the initial LLM draft blog post when using Papers-to-Posts as compared to the baseline tool. In Figure 9a, for each condition, we show the Levenshtein distance, a common metric for measuring the amount of editing done to text, between participants’ initial LLM draft blog posts and their blog posts at the point at which the writer would feel comfortable publishing it publicly. For exploratory statistical testing, we employed a paired-samples t-test and observed that this Levenshtein distance is significantly higher with Papers-to-Posts (M=3977.05,SD=1146.46) compared to the baseline (M=2764.15,SD=1042.16) (Paired-Samples t-Test, t(19)=4.31, p¡.0005).
In Figure 9a, we show how the Levenshtein distance progressed over time, with the removal of one outlier. The difference in Levenshtein distance does not appear attributable to the time spent on the task or the change in length of the blog post. Participants did not spend substantially more time to reach the point at which they were comfortable publicly publishing the blog post using Papers-to-Posts (M=49.37, Q1=35.43, Q3=59.54) as compared to the baseline (M=46.81, Q1=33.92,Q3=56.60) (Appendix 25). Furthermore, we saw that participants changed the length of their blog posts, measured in characters, less using Papers-to-Posts (M=-162.00, Q1=-1126.25, Q3=297.00) than using the baseline (M=1971.00, Q1=965.25,Q3=2514.25) (Appendix 25(b)). Thus, we suspected that Papers-to-Posts may have required a decreased number of writing actions (inserting, deleting, replacing text in text fields) in order to achieve a given increase in Levenshtein distance. We investigated this potential reduction in effort in the follow-up deployment study.
6. Deployment Study
Through the controlled lab study, we observed that participants produced final blog posts that were significantly more distant from their initial LLM-generated drafts when using Papers-to-Posts than the baseline tool. We wanted to further understand how participants in a more realistic setting perceive Papers-to-Posts versus the baseline tool (RQ2), and whether Papers-to-Posts would lead participants to change their provided drafts more in a given time period (RQ1). For this, we conducted a deployment study where participants wrote blog posts on their own time with the explicit goal of posting the article publicly and sharing it with others.
6.1. Participants
We recruited 26 participants (M: 17, W: 8, undisclosed: 1) through academic social networks and institutional mailing lists. They were compensated with $50 over PayPal for each blog post that they wrote, with a maximum of two blog posts. Participants were predominantly early career researchers (undergraduate student: 1, master’s student: 4, PhD student: 13, postdoc: 2, professor: 2, industry researcher: 4). All participants conducted research in an area of computer science, except for one who studied cognitive science. The majority of the participants had interacted with an LLM several times (¿10 sittings: 21, 6-10 sittings: 4, 2-5 sittings: 1). The participants had varied experience in writing blog posts (0 posts: 16, 1-2 posts: 9, 3-10 posts: 1). They were required to have authored at least one research paper (1-2 papers: 6, 3-10 papers: 11, ¿10 research papers: 9) and to be more interested than not (at least 5 on a 7-point Likert scale where 7 indicates the highest interest) in writing research-paper blog posts. We also required participants to be willing to share their final blog posts with others if they were able to reach the point at which they would feel comfortable doing so. Lastly, the papers that the participants want to write about had to be compatible with our tools’ PDF parser.
6.2. Study Conditions
While the treatment and baseline conditions were largely the same as in the lab study, several minor usability issues identified in the lab study were addressed before proceeding with the deployment study. For example, there were better instructions and less intrusive highlight colors (Appendix 29). We also replaced GPT-4 and GPT-4-32k with the newer GPT-4-0125-preview model with a larger context window that became available after the lab study. This enabled Papers-to-Poststo support longer inputs. In the deployment study, participants could freely close and reopen the tool in their browsers to work on their drafts when convenient. Participants could save their progress across sessions with a newly added save button. We note that, due to the manner in which text was logged in the deployment study, the tool in both conditions had a lag if participants pasted a chunk of text more then a couple paragraphs long. Lastly, the treatment tool now automatically parsed the input paper.
The majority of the 37 papers about which participants wrote were published in 2023 or 2024 (26 papers). The two oldest papers were from 2019 and 2008. The average parsed paper length across conditions differed by less than 5% (baseline: 37064 characters, treatment: 35845 characters).
6.3. Study Procedure
The task for the deployment study was to share a blog post with others, which can require significantly higher effort than the controlled lab study, which only required voicing comfort in publicly posting the blog post. To control for the study length and ease of recruitment, we opted for a between-subjects study so that each participant only had to write one blog post. At the same time, we allowed participants who wished to participate in the study twice (N=11/26 participants) to use both tools each time in randomized, counter-balanced order. When analyzing the results we treated them as non-paired data. Eighteen participants completed the baseline condition, and 19 participants completed the treatment condition. For participants who wrote two blog posts, their selected papers were assigned randomly to each condition. Participants had up to six days to complete a research-paper blog post using the assigned tool on their own time. We provided participants with a video tutorial on the assigned tool, a link to the paper about which they would be writing, and a link to the assigned tool. We noted that their aim should be to share the final blog post with others, but if they felt that they could not reach the point of being comfortable with sharing it with others, they could let us know and still be compensated. Participants were required to spend at least 20 minutes using the tool to write the blog post. To better understand how participants utilized the two systems, we recorded their interactions, including prompting, inserting, deleting and replacing text. In addition, we also took periodic snapshots of the drafts whenever they (re)opened the systems in their browsers and every 60 seconds during active usage.
Once participants were satisfied with their draft, participants shared a blog post based on their drafts. We provided instructions on how to create blog posts on the Medium platform,222https://medium.com/ but they were allowed to use other publishing platforms if they preferred. They then emailed us the blog post link and a screenshot or link showcasing evidence that they shared the blog post with others (e.g., tweet, email to co-authors, slack channel message, LinkedIn post). Only 2 of the 26 participants (one under both conditions and one under the treatment condition) alerted us to the fact that they could not reach the point at which they felt comfortable sharing the blog post with others in the allotted time.333We note that some participants indicated that the blog post was part of an experiment when sharing with others, which may have increased their willingness to share it; regardless, they were willing to share a post at the end of the study. Participants then completed the same survey from the lab study regarding their experience with the tool. We note that four participants who we realized had not completed 20 minutes of interaction were asked to do so after completing the survey. The only difference was the removal of a few questions irrelevant to this study: a question regarding how much additional time would be needed to feel comfortable posting the blog post publicly and questions regarding a direct comparison between the two tools. Lastly, for the 15 participants who were willing, the first author engaged them in a 15-minute semi-structured interview to discuss their experience with the tool.
6.4. Results
We did not plan to investigate statistical significance in this between-subjects study, given power analyses indicating that we would need a larger sample size than we could obtain to see a significant result for our metric of Levenshtein distance (power=0.95, alpha=0.05, effect size=0.45; requires 29 participants per condition) or blog post satisfaction (power=0.95, alpha=0.05, effect size=0.98; requires 83 participants per condition). Throughout our analysis of the deployment study logs, we exclude the three study runs in which the participant made no change to the initial draft, as there is no interaction in producing the blog post. We also exclude six study runs in which the participant claimed to have made edits outside the tool beyond small things like grammar, formatting, and adding images; this is to avoid drawing conclusions based on missing data about effort spent. The resulting log data used thus includes 28 study runs (treatment: 14, baseline: 14). The first author conducted inductive thematic analysis (Braun and Clarke, 2006) of the participants’ survey and interview comments.
6.4.1. RQ1: Blog Post Iteration
When given a time constraint in the lab study, we observed that participants made larger changes to the tool’s initial draft with Papers-to-Posts than with the baseline tool. The deployment study allowed us to investigate if this evident decrease in effort could be reflected in a realistic setting and in terms of other metrics of effort (i.e., number of writing actions). We found that participants using Papers-to-Posts took fewer writing actions in order to create one unit of change in Levenshtein distance to the blog post (Figure 9(c)). We also observed, in alignment with the lab study, that Papers-to-Posts led to greater changes in Levenshtein distance for a given amount of active time, which we define as time taking any action in the tool (Figure 9(b)). Therefore, when using Papers-to-Posts as compared to the baseline, participants experienced reduced effort in iterating on the blog post with respect to both writing actions needed and time needed.
Once more, this is despite the fact that participants did not spend substantially more active time using Papers-to-Posts (M=22.00, Q1=17.25,Q3=34.00) than the baseline (M=22.50,Q1=16.00,Q3=27.00) (Appendix 26(a)). Furthermore, we again saw that participants changed the length of their blog posts, measured in characters, less using Papers-to-Posts (M=-5.5, Q1=-525.25, Q3=961.00) than using the baseline (M=710.00, Q1=47.25, Q3=4261.25) (Appendix 26(b)). In fact, Figure 9 shows that the maximum active time and writing actions taken by a baseline participant is higher than that of a treatment participant. This provides some evidence that, for writers who tend to spend more time than average on the task, Papers-to-Posts may help them to reduce the time they spend on it.
Participants’ survey and interview comments reflected this discrepancy between the two tools. With respect to the baseline, participants most commonly cited the initial draft as a helpful aspect of the tool. Nine participants under the baseline condition specifically noted that it was useful as a starting point. Focusing on the initial draft, baseline participants often spoke of editing or adding to it rather than replacing it. For example, P20 explained, ”It’s easier to critique than it is to generate,” and P26 shared, ”It was very useful to have a straw-person of what the blog post could look like (generated by the LLM). I mostly used the draft as a place to edit and steer in the direction I wanted the post to take.”
On the other hand, 15 participants noted the selectable paper bullet points and paragraphs as a helpful feature of Papers-to-Posts. Focusing on the feature designed to generate full blog post sections, participants using Papers-to-Posts often reported doing just that. As an example, P12 commented that the ”outline [to] text generation process is very helpful to get the first draft of the post.” Furthermore, because the tool shows users what bullet points the tool used to generate the initial draft, participants may have been inclined to inspect and iterate upon this selection. This is illustrated by P20, who observed, ”The automatic selection of bullet points to include did not capture most of the information I wanted it to, and so required a lot of manual selection of bullet points, but once I selected those, the generation made it easy to write.” Participants also frequently mentioned the utility of the Papers-to-Posts’s modification abilities. P18 reflected, ”The ”MODIFY WRITING” section is helpful. I like that the interface contains this particular section to support iterating on existing text.” By facilitating modifications to the initial draft, Papers-to-Posts may have encouraged participants to iterate more heavily on the initial draft.
6.4.2. RQ2: Real-World Tool Perceptions
We evaluated participants’ survey results in the same manner as in the lab study (see Section 5.5). With respect to satisfaction with the final blog post quality, we observed no difference in the median satisfaction under Papers-to-Posts(M=6.00,Q1=5.00,Q3=6.00) versus the baseline (M=6.00,Q1=5.25, Q3=6.00) (Figure 5j-m). This discrepancy with the lab study may be related to the fact that participants were not writing under a time constraint in the deployment study. With sufficient time to write the blog post to their liking (only 3 blog posts were not shared with others), participants may not have noticed a pronounced difference in blog post quality between conditions.
Similarly to the lab study, we observed slightly greater satisfaction with the tool under the treatment condition (M=5.60,Q1=5.00, Q3=6.40) than the baseline condition (M=5.40,Q1=4.05,Q3=5.80) (Figure 5n-r). In alignment, four participants who participated twice expressed a preference for Papers-to-Posts, whereas no one shared a preference for the baseline tool. P19 remarked, ”I mean, honestly, I like the first tool better than the second tool… the second tool is a bit simplistic.” However, the complexity and larger associated learning curve of Papers-to-Posts may have decreased its chances of being widely preferred over the baseline. For example, P15 commented, ”It was a lot of information on the interface to deal with. I would prefer less moving things in front of me.” Nevertheless, we once again found slightly lower cognitive load under the treatment condition (M=15.00,Q1=9.50,Q3=17.50) in comparison to the baseline condition (M=15.50,Q1=10.75,Q3=18.75) (Figure 6e-h). This may be related to the fact that, while Papers-to-Posts required understanding a more complex interface, the baseline tool placed greater demands on participants in terms of writing. P24 summarized, ”In this [baseline] version, you have greater autonomy regarding what needs to be done. While this is beneficial, it also requires more effort and planning.”
As shown in Figure 7e-h, participants indicated higher satisfaction with control (treatment: M=5.00,Q1=3.00,Q3=6.00, baseline: M=3.00,Q1=3.00,Q3=4.75) in the treatment condition like in the lab study, but they showed no difference in median satisfaction with respect to perceived control (treatment: M=5.00,Q1=5.00,Q3=6.00, baseline: M=5.00,Q1=3.00,Q3=5.00), perceived ownership (treatment: M=6.00,Q1=4.50,Q3=6.00, baseline: M=6.00,Q1=6.00,Q3=6.00) or satisfaction with ownership (treatment: M=6.00,Q1=4.00,Q3=6.00, baseline: M=6.00,Q1=5.00,Q3=6.00) in the treatment condition. Participants may have been more satisfied with their control of the tool without perceiving a higher control using Papers-to-Posts in the sense that they were comfortable with letting the tool have a certain amount of control in exchange for less effort to generate text.
Overall, we observed benefits from Papers-to-Posts in fulfilling Papers-to-Posts’s design goals. Regarding the first design goal (DG1), participants indicated only a slight improvement in feeling confident that all essential information was covered with Papers-to-Posts (M=5.00,Q1=4.50,Q3=6.00) as opposed to the baseline (M=5.00,Q1=4.25,Q3=6.00). Given that the automatic method for parsing the paper PDF in the deployment study was likely to make more errors than the partially Wizard-of-Oz method for parsing papers in the lab study, it makes sense that participants would now see a reduced benefit from Papers-to-Posts’s paper outline in terms of ensuring all relevant information is included. Multiple participants under both conditions noted that the parsing of their papers was incomplete or flawed. However, participants still indicated an increased ease in incorporating information from the paper in the blog post with Papers-to-Posts (M=6.00,Q1=5.00,Q3=6.00) compared to the baseline (M=4.50,Q1=3.25,Q3=5.75). For instance, P14 commented, ”I liked being able to check the boxes of the portions of the paper I wanted to include in each section.”
With respect to our second design goal (DG2), there was no difference in the median for how well Papers-to-Posts (M=5.00,Q1=4.00, Q3=6.00) versus the baseline (M=5.00,Q1=5.00,Q3=6.00) facilitated production of appropriate text, unlike in the lab study (Figure 8e-h). Nevertheless, in their qualitative feedback, several participants noted how Papers-to-Posts’s generation and modification features helped them to create more appropriate text. For instance, P24 explained that the ”preset custom instructions was very useful because can need to be very specific using these tools,” and P14 noted that the ”provided buttons for making text more or less dramatic was nice.” Furthermore, in their survey responses, participants indicated that Papers-to-Posts (M=6.00,Q1=4.50,Q3=6.00) facilitated iterating on the blog post more so than the baseline (M=5.00,Q1=3.50,Q3=6.00), which is reflected in the lower number of writing actions participants took for a unit change in their initial drafts that we observed in Figure 9(c).
7. Discussion and Design Implications
Based on the results of the lab and deployment studies, we reflect on design implications for other human-LLM scientific long-document summarization tools with long-form outputs.
7.1. Post-Editing is not Enough
Both Papers-to-Posts and the baseline tool included initial LLM-generated drafts generated with powerful LLMs, GPT-4 and GPT-4-32k respectively. Thus, the existence of an initial LLM draft cannot be the primary reason why participants preferred the blog post generated with Papers-to-Posts as compared to the baseline tool. While the participants had the option to simply post-edit the initial draft and refrain from interacting further with the LLM, only two baseline and two treatment participants across the lab and deployment studies did so. This suggests that additional user interfaces and interaction designs can provide better scaffolding supports to further improve upon LLM-generated text compared to simple post-editing.
7.2. Long-Document Outlines
While blog post writing has additional goals and constraints compared to general document summarization, our work also has implications that may be useful for the task of general human-AI long document summarization. Specifically, a main challenge here is the high costs of human summarizer reviewing, locating, and utilizing different essential information in a long document. Prior work in human-LLM summarization has attempted to address this challenge in various manners, including presenting visualizations for the document to be summarized (Zhang et al., 2023b, 2011), dividing the document into chunks (Cheng et al., 2023), and providing document highlights for information already included in the summary (Norkute et al., 2021). However, to the best of our knowledge, no prior work has explored the use of generated outlines to help users review the document to be summarized.
In this work, participants found it easier to find and incorporate essential information from the document using Papers-to-Posts and participants consistently noted the selectable bullet points as one of the most useful features. We found that, given this tool, participants were able to make more change to their initial drafts for a given amount of time or writing actions. We therefore note that presenting LLM-generated outlines of scientific long documents may make the planning step of writing their summaries easier. As participants faced confusion about the connection between the outline and the paper text, we also recommend working to make this connection as clear as possible. How detailed the outline should be remains uncertain. While several participants appreciated the bullet points, a couple mentioned that the outline was overwhelming. As one participant proposed, a hierarchy of bullet points may be more useful.
7.3. Transparent Affordances for Long-form Writing Sub-Tasks
In contrast with shorter forms of scientific writing such as tweetorials (Long et al., 2023) and extended metaphors (Kim et al., 2023) explored in prior work on human-LLM scientific writing, longer scientific texts such as research-paper blog posts require completing a number of sub-tasks common to long-form science communication. For example, because research-paper blog posts are often for broader audiences than the paper itself, Papers-to-Posts includes a ”simpler terms” option to make large swaths of text more digestible for such readers. As another example, because different sections of a research-paper blog post may serve different purposes, we pre-populated the custom instructions for each of the initial draft sections (e.g., for the introduction section, we ask the tool to present a hook, a high-level description of the problem and why it is interesting, etc.). These transparent, section-specific instructions could then be verified or adjusted by the user. Several participants commented on the usefulness of Papers-to-Posts’s different affordances for streamlining different generation and modification sub-tasks such as these. Thus, we observe that similar tools may assist users by including affordances for sub-tasks related to generating and modifying long-form output.
8. Limitations and Future Work
The majority of the participants were PhD students in computer science. Although many of the takeaways discussed in this paper may generalize to other groups of researchers, it would be interesting to see how more diverse populations surface unique needs from the tool. For example, researchers in the field of computer science tend to be better versed in artificial intelligence than those in other fields, so they may not need as much guidance in utilizing an LLM writing tool. Different areas of research may also have very different traditional paper and blog post structures compared to computer science, which may benefit from different tool affordances. Furthermore, PhD students are prime candidates for writing blog posts, which can help them raise awareness of their work as they are starting their research careers. However, more senior researchers may desire specific features for writing their blog posts, such as a feature to coordinate discussion of multiple works or a feature to connect an older work to the current research landscape. We also leave it to future work to investigate how different aspects of the input paper such as length, recency, and author position may affect the support that users want from the tool.
9. Ethical Considerations
There are a few ethical considerations that we would like to note regarding this work. For one, it is well-known that LLMs may hallucinate information included in their generated text (Bubeck et al., 2023). We mitigate this risk by focusing on researchers summarizing their own work, which they should know well; for example, multiple participants noted inaccuracies or exaggerations in generated text, but most were able to correct those issues. A couple participants raised concerns that they had not caught all the errors by the end of their session, but as one of them noted, with more time they could be thorough in ensuring that the text is correct. We believe that LLM usage will increase in the future, and it is important to study effective design. Another ethical concern is whether using an LLM tool to write research-paper blog posts could lead to the loss of important details from the paper. P3 in the baseline condition raised this point: ”…it seems like the easier you make it, i.e., the better the initial summary, the less likely I am to myself remember those things that happen to be missing.” That said, participants overall felt more confident that all essential information was covered in their blog posts with Papers-to-Posts, so this work provided one step in the direction of support for reducing missing information. Furthermore, an LLM tool for writing research-paper blog posts may contribute to making science more accessible in society. This may raise concerns of reducing paper readership. However, blog posts do not provide the same amount of information as papers and do not go through peer review, research papers still hold their own value, and blog posts can act as catalysts for paper reading. Finally, researchers may not acknowledge the use of AI in writing their blog posts, which could mislead readers, particularly if the researcher has not verified the blog post contents. For the participants who shared their blog posts, we found that many chose to disclose its use, but to ensure better transparency around its use, future versions of Papers-to-Posts may highlight the provenance of the output writing (e.g., by visualizing the writer’s interaction with the LLM-powered tool (Naimul Hoque et al., 2023)).
10. Conclusion
We introduce Papers-to-Posts, an LLM-powered research-paper blog post writing tool, which effectively helps researchers plan, draft, and revise their blog posts, Through a formative study (N=6) with an LLM technology probe for this task, we identified the following challenges faced by writers: they 1) appreciated a paper outline for finding and selecting relevant information from the paper but faced difficulties in navigating it and 2) tried to provide recurrent instructions for different kinds of sub-tasks to the LLM without specific affordances to support these sub-tasks. We formulated design goals to address each of these challenges, namely to help writers 1) plan their writing by facilitating review of the long document’s contents and selection of essential information to include in the writing and 2) draft and revise their writing by streamlining commonly desired instructions to the LLM for generating and modifying text. We validated that our LLM-powered tool Papers-to-Posts addresses these goals and helps authors write blog posts for their own papers through two user studies comparing Papers-to-Posts against a baseline providing an LLM blog post draft and free-form LLM prompting. In a within-subjects, time-controlled lab study (N=20 participants), we found that Papers-to-Posts significantly increased writer satisfaction with the output blog post, without increasing task completion time or cognitive load. We also observe that Papers-to-Posts led to significantly greater change to the initial LLM draft blog post, indicating that this increased satisfaction may be due to the tool facilitating iteration on the blog post. A field deployment study (N=37 blog posts, 26 participants) showed that in a more realistic setting Papers-to-Posts again resulted in greater change in the blog post for a given amount of editing time or number of editing actions. We hope that this work may inform the design of future human-LLM tools for scientific long-document summarization into long-form output.
Acknowledgements.
This research was supported by the Allen Institute for Artificial Intelligence (AI2). The authors thank the Semantic Scholar team for helpful feedback on this work and the participants who made this work possible.References
- (1)
- August et al. (2020) Tal August, Lauren Kim, Katharina Reinecke, and Noah A Smith. 2020. Writing strategies for science communication: Data and computational analysis. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP). 5327–5344.
- Avinesh et al. (2018) PVS Avinesh, Carsten Binnig, Benjamin Hättasch, Christian M Meyer, and Orkan Özyurt. 2018. Sherlock: A System for Interactive Summarization of Large Text Collections. Proc. VLDB Endow. 11, 12 (2018), 1902–1905.
- Avinesh and Meyer (2017) PVS Avinesh and Christian M Meyer. 2017. Joint optimization of user-desired content in multi-document summaries by learning from user feedback. In Proceedings of the 55th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). 1353–1363.
- Bayatmakou et al. (2022) Farnoush Bayatmakou, Azadeh Mohebi, and Abbas Ahmadi. 2022. An interactive query-based approach for summarizing scientific documents. Information Discovery and Delivery 50, 2 (2022), 176–191.
- Blanchard (2011) Antoine Blanchard. 2011. 3.4 Science Blogs in Research and Popularization of Science: Why, how and for whom? Common knowledge: The challenge of transdisciplinarity (2011), 219–232.
- Bohn and Ling (2021) Tanner Bohn and Charles X Ling. 2021. Hone as you read: A practical type of interactive summarization. arXiv preprint arXiv:2105.02923 (2021).
- Braun and Clarke (2006) Virginia Braun and Victoria Clarke. 2006. Using thematic analysis in psychology. Qualitative research in psychology 3, 2 (2006), 77–101.
- Bubeck et al. (2023) Sébastien Bubeck, Varun Chandrasekaran, Ronen Eldan, Johannes Gehrke, Eric Horvitz, Ece Kamar, Peter Lee, Yin Tat Lee, Yuanzhi Li, Scott Lundberg, et al. 2023. Sparks of artificial general intelligence: Early experiments with gpt-4. arXiv preprint arXiv:2303.12712 (2023).
- Burns et al. (2003) Terry W Burns, D John O’Connor, and Susan M Stocklmayer. 2003. Science communication: a contemporary definition. Public understanding of science 12, 2 (2003), 183–202.
- Cai et al. (2022) Pengshan Cai, Fei Liu, Adarsha Bajracharya, Joe Sills, Alok Kapoor, Weisong Liu, Dan Berlowitz, David Levy, Richeek Pradhan, and Hong Yu. 2022. Generation of patient after-visit summaries to support physicians. In Proceedings of the 29th International Conference on Computational Linguistics (COLING).
- Cheng et al. (2023) Chen Cheng, Ziang Li, Zhenhui Peng, and Quan Li. 2023. ALens: An Adaptive Domain-Oriented Abstract Writing Training Tool for Novice Researchers. arXiv preprint arXiv:2308.03989 (2023).
- Cheng et al. (2022) Ruijia Cheng, Alison Smith-Renner, Ke Zhang, Joel R Tetreault, and Alejandro Jaimes. 2022. Mapping the design space of human-ai interaction in text summarization. arXiv preprint arXiv:2206.14863 (2022).
- Chung et al. (2022) John Joon Young Chung, Wooseok Kim, Kang Min Yoo, Hwaran Lee, Eytan Adar, and Minsuk Chang. 2022. TaleBrush: Sketching Stories with Generative Pretrained Language Models. Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (2022). https://api.semanticscholar.org/CorpusID:247625751
- CMU ([n. d.]) CMU. [n. d.]. CMU. https://blog.ml.cmu.edu/submissions/.
- Coenen et al. (2021) Andy Coenen, Luke Davis, Daphne Ippolito, Emily Reif, and Ann Yuan. 2021. Wordcraft: a human-ai collaborative editor for story writing. arXiv preprint arXiv:2107.07430 (2021).
- Dang et al. (2022) Hai Dang, Karim Benharrak, Florian Lehmann, and Daniel Buschek. 2022. Beyond text generation: Supporting writers with continuous automatic text summaries. In Proceedings of the 35th Annual ACM Symposium on User Interface Software and Technology. 1–13.
- Ding et al. (2023) Zijian Ding, Alison Smith-Renner, Wenjuan Zhang, Joel R Tetreault, and Alejandro Jaimes. 2023. Harnessing the power of LLMs: Evaluating human-AI text co-creation through the lens of news headline generation. arXiv preprint arXiv:2310.10706 (2023).
- El-Kassas et al. (2021) Wafaa S El-Kassas, Cherif R Salama, Ahmed A Rafea, and Hoda K Mohamed. 2021. Automatic text summarization: A comprehensive survey. Expert systems with applications 165 (2021), 113679.
- Flower and Hayes (1981) Linda S. Flower and J. R. Hayes. 1981. A Cognitive Process Theory of Writing. College Composition and Communication 32 (1981), 365–387. https://api.semanticscholar.org/CorpusID:18484126
- Gao et al. (2020) Yang Gao, Christian M Meyer, and Iryna Gurevych. 2020. Preference-based interactive multi-document summarisation. Information Retrieval Journal 23, 6 (2020), 555–585.
- Gehrmann et al. (2019) Sebastian Gehrmann, Hendrik Strobelt, Robert Krüger, Hanspeter Pfister, and Alexander M Rush. 2019. Visual interaction with deep learning models through collaborative semantic inference. IEEE transactions on visualization and computer graphics 26, 1 (2019), 884–894.
- Gero et al. (2021) K. Gero, Vivian Liu, Sarah Huang, Jennifer Lee, and Lydia B. Chilton. 2021. What Makes Tweetorials Tick: How Experts Communicate Complex Topics on Twitter. Proceedings of the ACM on Human-Computer Interaction 5 (2021), 1 – 26. https://api.semanticscholar.org/CorpusID:239020743
- Gero et al. (2022) Katy Ilonka Gero, Vivian Liu, and Lydia Chilton. 2022. Sparks: Inspiration for science writing using language models. In Designing interactive systems conference. 1002–1019.
- Ghazvininejad et al. (2017) Marjan Ghazvininejad, Xing Shi, Jay Priyadarshi, and Kevin Knight. 2017. Hafez: an interactive poetry generation system. In Proceedings of ACL 2017, System Demonstrations. 43–48.
- Goldfarb-Tarrant et al. (2019) Seraphina Goldfarb-Tarrant, Haining Feng, and Nanyun Peng. 2019. Plan, write, and revise: an interactive system for open-domain story generation. arXiv preprint arXiv:1904.02357 (2019).
- Google ([n. d.]) Google. [n. d.]. BARD. https://bard.google.com/chat.
- Hart and Staveland (1988) Sandra G Hart and Lowell E Staveland. 1988. Development of NASA-TLX (Task Load Index): Results of empirical and theoretical research. In Advances in psychology. Vol. 52. Elsevier, 139–183.
- Holmberg and Thelwall (2014) Kim Holmberg and Mike Thelwall. 2014. Disciplinary differences in Twitter scholarly communication. Scientometrics 101 (2014), 1027–1042.
- Jarreau (2015) Paige Brown Jarreau. 2015. All the science that is fit to blog: An analysis of science blogging practices. Louisiana State University and Agricultural & Mechanical College.
- Jucan and Jucan (2014) Mihaela Sabina Jucan and Cornel Nicolae Jucan. 2014. The power of science communication. Procedia-Social and Behavioral Sciences 149 (2014), 461–466.
- Kim et al. (2023) Jeong Chul Kim, Sang Guk Suh, Lydia B. Chilton, and Haijun Xia. 2023. Metaphorian: Leveraging Large Language Models to Support Extended Metaphor Creation for Science Writing. Proceedings of the 2023 ACM Designing Interactive Systems Conference (2023). https://api.semanticscholar.org/CorpusID:259376400
- Kocielnik et al. (2019) Rafal Kocielnik, Saleema Amershi, and Paul N Bennett. 2019. Will you accept an imperfect ai? exploring designs for adjusting end-user expectations of ai systems. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems. 1–14.
- Koh et al. (2022) Huan Yee Koh, Jiaxin Ju, Ming Liu, and Shirui Pan. 2022. An empirical survey on long document summarization: Datasets, models, and metrics. ACM computing surveys 55, 8 (2022), 1–35.
- Kudos ([n. d.]) Kudos. [n. d.]. Kudos. https://info.growkudos.com/landing/researchers-2023.
- Laban et al. (2023) Philippe Laban, Jesse Vig, Marti A Hearst, Caiming Xiong, and Chien-Sheng Wu. 2023. Beyond the Chat: Executable and Verifiable Text-Editing with LLMs. arXiv preprint arXiv:2309.15337 (2023).
- Lai et al. (2022) Vivian Lai, Alison Smith-Renner, Ke Zhang, Ruijia Cheng, Wenjuan Zhang, Joel Tetreault, and Alejandro Jaimes. 2022. An exploration of post-editing effectiveness in text summarization. arXiv preprint arXiv:2206.06383 (2022).
- Long et al. (2023) Tao Long, Dorothy Zhang, Grace Li, Batool Taraif, Samia Menon, Kynnedy Simone Smith, Sitong Wang, Katy Ilonka Gero, and Lydia B. Chilton. 2023. Tweetorial Hooks: Generative AI Tools to Motivate Science on Social Media. ArXiv abs/2305.12265 (2023). https://api.semanticscholar.org/CorpusID:258833277
- Lu et al. (2023) Guang Lu, Sylvia B Larcher, and Tu Tran. 2023. Hybrid Long Document Summarization using C2F-FAR and ChatGPT: A Practical Study. arXiv preprint arXiv:2306.01169 (2023).
- Luc et al. (2021) Jessica GY Luc, Michael A Archer, Rakesh C Arora, Edward M Bender, Arie Blitz, David T Cooke, Tamara Ni Hlci, Biniam Kidane, Maral Ouzounian, Thomas K Varghese Jr, et al. 2021. Does tweeting improve citations? One-year results from the TSSMN prospective randomized trial. The Annals of thoracic surgery 111, 1 (2021), 296–300.
- Mahrt and Puschmann (2014) Merja Mahrt and Cornelius Puschmann. 2014. Science blogging: An exploratory study of motives, styles, and audience reactions. Journal of science Communication 13, 03 (2014), A05.
- Medium ([n. d.]) Medium. [n. d.]. Medium. https://blog.medium.com/best-practices-for-writing-on-medium-386506ae62b9.
- Mehlenbacher (2019) Ashley Mehlenbacher. 2019. Science communication online: Engaging experts and publics on the internet. The Ohio State University Press.
- Mirowski et al. (2023) Piotr Mirowski, Kory W Mathewson, Jaylen Pittman, and Richard Evans. 2023. Co-Writing Screenplays and Theatre Scripts with Language Models: Evaluation by Industry Professionals. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–34.
- Moramarco et al. (2021) Francesco Moramarco, Alex Papadopoulos Korfiatis, Aleksandar Savkov, and Ehud Reiter. 2021. A preliminary study on evaluating consultation notes with post-editing. arXiv preprint arXiv:2104.04402 (2021).
- Naimul Hoque et al. (2023) Md Naimul Hoque, Tasfia Mashiat, Bhavya Ghai, Cecilia Shelton, Fanny Chevalier, Kari Kraus, and Niklas Elmqvist. 2023. The HaLLMark Effect: Supporting Provenance and Transparent Use of Large Language Models in Writing through Interactive Visualization. arXiv e-prints (2023), arXiv–2311.
- Norkute et al. (2021) Milda Norkute, Nadja Herger, Leszek Michalak, Andrew Mulder, and Sally Gao. 2021. Towards explainable AI: Assessing the usefulness and impact of added explainability features in legal document summarization. In Extended Abstracts of the 2021 CHI Conference on Human Factors in Computing Systems. 1–7.
- NotionAI ([n. d.]) NotionAI. [n. d.]. NotionAI. https://www.notion.so/product.
- OpenAI ([n. d.]) OpenAI. [n. d.]. chatGPT. https://chat.openai.com/.
- Patrick Dunleavy ([n. d.]) Patrick Dunleavy. [n. d.]. LSE. https://blogs.lse.ac.uk/impactofsocialsciences/2016/01/25/how-to-write-a-blogpost-from-your-journal-article/.
- Singh et al. (2022) Nikhil Singh, Guillermo Bernal, Daria Savchenko, and Elena L. Glassman. 2022. A Selective Summary of Where to Hide a Stolen Elephant: Leaps in Creative Writing with Multimodal Machine Intelligence. ACM Transactions on Computer-Human Interaction (2022). https://api.semanticscholar.org/CorpusID:247085270
- UWaterloo ([n. d.]) UWaterloo. [n. d.]. UWaterloo. https://uwaterloo.ca/writing-and-communication-centre/writing-academic-blog-posts.
- Van Someren et al. (1994) Maarten Van Someren, Yvonne F Barnard, and J Sandberg. 1994. The think aloud method: a practical approach to modelling cognitive. London: AcademicPress 11, 6 (1994).
- Wang et al. (2023a) Sitong Wang, Samia Menon, Tao Long, Keren Henderson, Dingzeyu Li, Kevin Crowston, Mark Hansen, Jeffrey V Nickerson, and Lydia B Chilton. 2023a. Reelframer: Co-creating news reels on social media with generative ai. arXiv preprint arXiv:2304.09653 (2023).
- Wang et al. (2023b) Sitong Wang, Zheng Ning, Anh Truong, Mira Dontcheva, Dingzeyu Li, and Lydia B Chilton. 2023b. PodReels: Human-AI Co-Creation of Video Podcast Teasers. arXiv preprint arXiv:2311.05867 (2023).
- Williams et al. (2022) Spencer Williams, Ridley Jones, Katharina Reinecke, and Gary Hsieh. 2022. An HCI Research Agenda for Online Science Communication. Proceedings of the ACM on Human-Computer Interaction 6 (2022), 1 – 22. https://api.semanticscholar.org/CorpusID:253459983
- Xie et al. (2023) Yujia Xie, Xun Wang, Si-Qing Chen, Wayne Xiong, and Pengcheng He. 2023. Interactive Editing for Text Summarization. arXiv preprint arXiv:2306.03067 (2023).
- Yan et al. (2011) Rui Yan, Jian-Yun Nie, and Xiaoming Li. 2011. Summarize what you are interested in: An optimization framework for interactive personalized summarization. In Proceedings of the 2011 conference on empirical methods in natural language processing. 1342–1351.
- Zarinbal et al. (2019) Marzieh Zarinbal, Azadeh Mohebi, Hesamoddin Mosalli, Razieh Haratinik, Zahra Jabalameli, and Farnoush Bayatmakou. 2019. A new social robot for interactive query-based summarization: Scientific document summarization. In Interactive Collaborative Robotics: 4th International Conference, ICR 2019, Istanbul, Turkey, August 20–25, 2019, Proceedings 4. Springer, 330–340.
- Zhang et al. (2023b) Xiaoyu Zhang, Jianping Li, Po-Wei Chi, Senthil Chandrasegaran, and Kwan-Liu Ma. 2023b. ConceptEVA: Concept-Based Interactive Exploration and Customization of Document Summaries. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–16.
- Zhang et al. (2011) Yi Zhang, Dingding Wang, and Tao Li. 2011. iDVS: an interactive multi-document visual summarization system. In Machine Learning and Knowledge Discovery in Databases: European Conference, ECML PKDD 2011, Athens, Greece, September 5-9, 2011, Proceedings, Part III 22. Springer, 569–584.
- Zhang et al. (2023a) Zheng Zhang, Jie Gao, Ranjodh Singh Dhaliwal, and Toby Jia-Jun Li. 2023a. Visar: A human-ai argumentative writing assistant with visual programming and rapid draft prototyping. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology. 1–30.
- Zou and Hyland (2019) Hang Zou and Ken Hyland. 2019. Reworking research: Interactions in academic articles and blogs. Discourse studies 21, 6 (2019), 713–733.
Appendix A Prompts
A.1. Prompt for Initial LLM Draft for Baseline Tool
Figure 10 shows the prompt for the initial LLM draft for the baseline tool.
A.2. Prompts for Initial LLM Draft for Papers-to-Posts
A.3. Prompts for Working with Papers-to-Posts
Appendix B Additional Log Analysis Plots
Figures 25 and 26 show additional log analysis plots. Figure 27 shows the distribution of behavioral actions taken.
Appendix C Formative Study Technology Probe
Figure 28 shows the formative study technology probe.
Appendix D Updated Tool for Deployment Study
Figure 29 shows the updated tool for the deployment study.

### v2

@@
 Abstract.
-Research-paper blog posts help scientists disseminate their work to a larger audience, but translating papers into this format requires substantial additional effort.
-Blog post creation is not simply transforming a long-form article into a short output, as studied in most prior work on human-AI summarization.
-In contrast, blog posts are typically full-length articles that require a combination of strategic planning grounded in the source document, well-organized drafting, and thoughtful revisions.
-Can tools powered by large language models (LLMs) assist scientists in writing research-paper blog posts?
-To investigate this question, we conducted a formative study (N=6) to understand the main challenges of writing such blog posts with an LLM: high interaction costs for 1) reviewing and utilizing the paper content and 2) recurrent sub-tasks of generating and modifying the long-form output.
-To address these challenges, we developed Papers-to-Posts, an LLM-powered tool that implements a new Plan-Draft-Revise workflow, which 1) leverages an LLM to generate bullet points from the full paper to help users find and select content to include (Plan) and 2) provides default yet customizable LLM instructions for generating and modifying text (Draft, Revise).
-Through a within-subjects lab study (N=20) and between-subjects deployment study (N=37 blog posts, 26 participants) in which participants wrote blog posts about their papers, we compared Papers-to-Posts to a strong baseline tool that provides an LLM-generated draft and access to free-form LLM prompting.
-Results show that Papers-to-Posts helped researchers to 1) write significantly more satisfying blog posts and make significantly more changes to their blog posts in a fixed amount of time without a significant change in cognitive load (lab) and 2) make more changes to their blog posts for a fixed number of writing actions (deployment).
+Research-paper blog posts help scientists to disseminate their work to a larger audience, but translating scientific long documents into long-form summaries like blog posts raises unique challenges: 1) planning what paper content to include in the blog post, 2) drafting the selected content in sections amenable to a paper blog post, and 3) revising the blog post to be scientifically accurate but also concise, easy to understand, and engaging.
+Can we harness the power of large language models (LLMs) to assist researchers with these challenges?
+To investigate this question, we developed Papers-to-Posts, an LLM-powered tool that implements a new Plan-Draft-Revise workflow for mixed-initiative long-form paper summarization.
+An LLM-generated paper outline with pre-selected yet adjustable bullet points helps users to plan what information to include.
+Meanwhile, customizable LLM instructions support drafting the text with a suitable structure and revising the text to have an appropriate tone.
+Through two studies, we compared Papers-to-Posts to a strong baseline tool that provides an LLM-generated draft and access to free-form LLM prompting, and we found that Papers-to-Posts improved researchers’ editing power.
+In a within-subjects lab study (N=20 participants), Papers-to-Posts led participants to make significantly more change to initial LLM drafts within a fixed amount of time and to be significantly more satisfied with their final blog post, without increasing cognitive load.
+Furthermore, in a between-subjects deployment study (N=37 blog posts, 26 participants), Papers-to-Posts led participants to make more change to initial LLM drafts within a given amount of time as well as writing actions, without decreasing satisfaction with the final blog posts or increasing cognitive load.
 1. Introduction
@@
 While some tools have been created to help author short-form derivatives like tweets (Long et al., 2023), significant challenges remain for scientists seeking to create longer-form derivatives like blog posts, which require more extensive planning, drafting, and revising writing processes (Flower and Hayes, 1981).
-In the planning step, the blog post author must review the paper, itself a long document, to find and select several pieces of information to include in the blog post.
-In the drafting and revising steps, there are many sub-tasks for completing the full blog post, such as creating an engaging hook or reporting results with sufficient detail but in simpler terms than the paper.
-Integrating large language models (LLMs) in writing tools offers a promising avenue for improving writing outcomes; however, existing approaches are not designed to help create long-form derivatives like blog posts from long documents like academic papers.
-Prior work in human-LLM text summarization has focused on supporting short-form summaries, a sentence or paragraph in length (Zhang et al., 2023b; Lai et al., 2022).
-In the broader area of human-LLM writing, prior work has looked into providing assistance for writing long-form creative texts such as stories (Chung et al., 2022; Goldfarb-Tarrant et al., 2019; Singh et al., 2022) and screenplays (Mirowski et al., 2023), but these tasks do not take a long-form text input.
-Prior work in human-LLM scientific writing has investigated how LLM tools can assist scientists in writing short-form texts such as tweetorials (Long et al., 2023) and extended metaphors (Kim et al., 2023), but to the best of our knowledge, no work has sought to develop an LLM tool for creating long-form scientific texts, such as research-paper blog posts.
-Still, prior work has shown that LLM tools can support planning, drafting, and revising for different forms of writing, such as creative stories (Goldfarb-Tarrant et al., 2019), short summaries (Zhang et al., 2023b), and short scientific texts (Long et al., 2023).
-Thus, in this paper we ask: how best can an LLM tool provide assistance in dealing with the challenges of writing long-form summaries of scientific long documents?
-We designed a technology probe of a simple writing interface with preliminary LLM-based features to better understand the costs and benefits of interacting with LLMs for the purpose of planning, drafting, and revising a research-paper blog post.
-The technology probe provides the paper text, an LLM-generated bullet point outline for the paper, a drafting space to write the blog post in sections, and access to an LLM prompting interface for writing tasks.
-Through a formative study (N=6), we gathered feedback on how users wish to leverage LLMs in this context and common interaction challenges.
-For one, the participants valued having a paper outline for finding and selecting relevant information from the paper to include in the blog post, but they struggled with understanding and utilizing it (e.g., understanding how the bullet points were connected to the paper text).
-Also, participants desired to provide the LLM recurring instructions for generating and modifying text that were not well-supported by the tool (e.g., rewrite a paragraph to better capture readers’ attention, continue writing from a particular sentence).
-Based on the identified challenges, we iterated on the technology probe to develop the prototype tool Papers-to-Posts for LLM-assisted paper blog post writing.
+Integrating large language models (LLMs) in writing tools offers a promising avenue for improving editing power (i.e., change in blog post per minute or writing action).
+However, existing approaches are not designed to help create long-form derivatives like blog posts from scientific long documents like academic papers.
+Prior work in human-LLM text summarization (Zhang et al., 2023b; Lai et al., 2022) and human-LLM scientific writing (Long et al., 2023; Kim et al., 2023) has focused on supporting short-form summaries, a sentence or paragraph in length.
+In the broader area of human-LLM writing, prior work has looked into providing assistance for writing long-form creative texts (Chung et al., 2022; Goldfarb-Tarrant et al., 2019; Singh et al., 2022; Mirowski et al., 2023), but these tasks do not take a long-form text input.
+To the best of our knowledge, no work has developed and evaluated a mixed-initiative, LLM-powered tool for generating long-form summaries of scientific long documents, such as research-paper blog posts.
+Shamsabadi and D’Souza presented a prompt-based tool to support different research tasks including writing blog posts; however, the tool’s main function is to generate a personalized prompt for the user to then provide to an LLM, and the tool was not evaluated (Shamsabadi and D’Souza, 2024).
+Still, prior work has shown that LLM-powered tools can support planning, drafting, and revising different writing forms, such as creative stories (Goldfarb-Tarrant et al., 2019), short summaries (Zhang et al., 2023b), and short scientific texts (Long et al., 2023).
+Thus, in this paper we ask: can an LLM-powered tool provide assistance in dealing with the challenges of writing long-form summaries of scientific long documents?
+As indicated by the many guidelines that discuss the content, formatting, and language of academic blog posts (Kudos, [n.
+d.]; UWaterloo, [n.
+d.]; Patrick Dunleavy, [n.
+d.]; CMU, [n.
+d.]), there are multiple challenges specific to summarizing a research paper in a blog post.
+The first distinct challenge is related to the input being a long document.
+The writer must sift through the paper to determine what information to carry over to the blog post, without missing important details.
+The two other main challenges are related to the output being long-form and for science communication.
+The writer must organize key points in sections amenable to a scientific blog post.
+Also, the writer must ensure that their writing is scientifically accurate like the paper but more concise, easy to understand, and eye-catching than the paper.
+To help researchers write research-paper blog posts, we developed Papers-to-Posts, a mixed-initiative tool designed based on guidelines for academic blog post writing (Kudos, [n.
+d.]; UWaterloo, [n.
+d.]; Patrick Dunleavy, [n.
+d.]; CMU, [n.
+d.]; Mehlenbacher, 2019) as well as pilot study feedback.
 The tool follows our Plan-Draft-Revise workflow (Figure 1).
-In the Plan phase, the user may control text generation by selecting a subset of LLM-generated bullet points corresponding to key points in the source; Papers-to-Posts is the first LLM-powered long-document summarization tool to support direct control of content selection through a source outline, maintaining the source’s narrative structure.
-In the Draft phase, the user may supply additional instructions and preview LLM generations that use the selected bullet points, instructions, and blog post text thus far.
-In the Revise phase, they may use an LLM to further refine text with macros that encapsulate frequent editing transformations (e.g., to have less jargon, to be more dramatic).
-We evaluated Papers-to-Posts through two studies in which participants wrote blog posts for their own papers, and we compared our tool to a baseline tool that provided an LLM-generated draft and integrated prompting.
-In a within-subjects lab study (N=20) in which researchers wrote blog posts for two of their own papers, one per tool, we found that, by helping authors to plan, draft, and revise their blog post, Papers-to-Posts led to significantly higher writer satisfaction with the blog post output alongside significantly greater change to the provided draft, without increasing task completion time or cognitive load.
-To study Papers-to-Posts in a more realistic setting, we also conducted an unmonitored, between-subjects deployment study (N=37), in which participants using Papers-to-Posts again made greater change to the provided drafts for a given amount of time spent or number of writing actions.
-Furthermore, through quantitative survey ratings and behavioral analysis as well as qualitative short answers and semi-structured interviews across the two studies, we observe that an LLM-generated paper outline and affordances for streamlining common long-form text generation and modification sub-tasks contributed to this increased satisfaction.
-In discussing our results, we also reflect on design implications for other human-LLM scientific long-document summarization tools with long-form output.
-In summary, we make the following main contributions:
-A formative study (N=6) producing two primary design goals for an LLM-powered tool for writing research-paper blog posts: (1) help users review and select content from a long source document during the planning stage, and (2) transparently streamline the process of instructing the LLM to generate and modify text during the drafting and revising stages.
-Papers-to-Posts, a prototype tool for writing research-paper blog posts, which addresses the design goals identified in the formative study and implements the Plan-Draft-Revise workflow (Figure 1).
-A within-subjects lab study (N=20 participants) providing evidence that Papers-to-Posts improves upon a powerful baseline (LLM draft and free-form prompting).
-Papers-to-Posts led to significantly higher writer satisfaction with the final blog post (with comparable task completion time and cognitive load) alongside significantly greater changes to the initial draft.
-A field deployment study (N=37 blog posts, 26 participants) comparing Papers-to-Posts to the same strong baseline show that the same benefits hold in realistic scenarios where participants were writing blog posts on their own and sharing them publicly at the end of the study.
-Design implications for other human-LLM scientific long-document summarization tools with long-form output: avoid solely post-editing interactions, present long-document outlines to support content selection, and include transparent affordances to streamline sub-tasks in generating and modifying long-form output text.
+In the Plan phase, the user may adjust the tool’s selection of key bullet points within an LLM-generated paper outline.
+Inspired by prior work on LLM-supported reverse outlining (Dang et al., 2022), Papers-to-Posts is, to the best of our knowledge, the first LLM-powered long-document summarization tool to support direct control of content selection through an LLM-generated document outline, maintaining the source’s narrative structure.
+In the Draft phase, the user may customize preset instructions (e.g., include a hook, make one paragraph) to the LLM for each blog post section.
+The LLM then uses the selected bullet points and adjusted instructions to generate text for the given section.
+In the Revise phase, the user may further refine text with LLM-powered macros that encapsulate frequent editing transformations for scientific blog post writing (e.g., to reduce jargon, to better capture attention).
+We evaluated Papers-to-Posts through two studies in which participants wrote blog posts for their own papers.
+In these studies, we compared our tool to a strong baseline tool–an LLM-generated draft and access to free-form LLM prompting.
+In a within-subjects lab study (N=20 participants) in which researchers wrote blog posts for two papers (one per tool), we found that Papers-to-Posts led to significantly greater change to the provided draft within a fixed amount of time alongside significantly higher writer satisfaction with the final blog post, without significantly increasing cognitive load.
+To study Papers-to-Posts in a more realistic setting, we also conducted an unmonitored, between-subjects deployment study (N=37 blog posts, 26 participants).
+Again, participants using Papers-to-Posts made greater change to the provided drafts for a given amount of time and for a given number of writing actions taken.
+Through surveys, interviews, and logs from across the two studies, we also observed evidence that an interactive LLM-generated paper outline for planning writing as well as preset but customizable LLM instructions for drafting and revising writing contributed to this increased editing power.
+In discussing our results, we reflect on implications for other human-LLM tools for research-paper summarization and long-form scientific communication: avoid solely post-editing interactions, provide an interactive outline of the long document for planning the summary, and include preset but adjustable LLM instructions for generating and modifying the writing.
+In summary, we make the following contributions:
+Papers-to-Posts, a tool for writing research-paper blog posts that implements the Plan-Draft-Revise workflow (Figure 1 and Section 3), including providing an interactive LLM-generated paper outline for planning the blog post, section-specific and customizable LLM instructions for drafting the blog post, and macros for revising text style to suit the blog post.
+Findings from a within-subjects lab study (N=20 participants) showing that Papers-to-Posts significantly increases researchers’ editing power (i.e., change in writing within a fixed amount of time) (Figure 5a and Section 5.1.1), while significantly increasing satisfaction with the output blog post (Figure 6a and Section 5.2.1) and without significantly impacting cognitive load.
+Findings from a between-subjects deployment study (N=37 blog posts, 26 participants) demonstrating that Papers-to-Posts increases researchers’ editing power (i.e., change in writing within an amount of time or writing actions) (Figure 5b-c and Section 5.1.2), without decreasing satisfaction with the output blog post or increasing cognitive load.
 2. Related Work
-2.1. Human-LLM Summarization
+2.1. Human-LLM Text Summarization
 Although there is much work on automatic text summarization (Koh et al., 2022; El-Kassas et al., 2021), issues remain with the output summaries, such as inaccuracies and style misalignment (Lu et al., 2023; Koh et al., 2022).
 To address these issues, other works have looked into having a human and LLM work together to generate a summary (Lai et al., 2022; Moramarco et al., 2021).
-According to Cheng et al., there are five common methods of human-AI summarization: guiding AI output, selecting or rating AI output, post-editing of AI-generated summaries, interactive editing, and writing with AI assistance (Cheng et al., 2022).
+There are several works on human-LLM summarization of long documents (Zhang et al., 2011; Gehrmann et al., 2019) and some for scientific long documents in particular (Zarinbal et al., 2019; Bayatmakou et al., 2022; Zhang et al., 2023b; Cheng et al., 2023).
+Long documents provide a unique challenge for human-LLM summarization, as there is a great deal of information for the human to digest.
+As writing research-paper blog posts entails summarizing portions of a research paper, this work seeks to support human-LLM long-document summarization.
+According to Cheng et al., there are five common methods of human-AI summarization: guiding AI output, selecting or rating AI output, post-editing AI-generated summaries, interactive editing, and writing with AI assistance (Cheng et al., 2022).
 Our tool Papers-to-Posts provides affordances related to guiding and post-editing AI output.
-Prior work in human-AI summarization has mainly focused on direct model guidance (e.g., utilizing concepts/queries (Zhang et al., 2023b; Bayatmakou et al., 2022; Avinesh and Meyer, 2017; Avinesh et al., 2018; Ding et al., 2023; Wang et al., 2023b), typing (Xie et al., 2023), ratings (Gao et al., 2020; Bohn and Ling, 2021), gaze (Bohn and Ling, 2021)) or simply post-editing (e.g., (Lai et al., 2022; Moramarco et al., 2021)).
-A few works have supported summarization based on user-selected points from the source (Bayatmakou et al., 2022; Yan et al., 2011; Zhang et al., 2011; Wang et al., 2023a, b), but the points are ranked or clustered by the system without the same narrative structure as the source document.
+Prior work in human-AI summarization has mainly focused on model guidance through various means (e.g., concepts/queries (Zhang et al., 2023b; Bayatmakou et al., 2022; Avinesh and Meyer, 2017; Avinesh et al., 2018; Ding et al., 2023; Wang et al., 2023b), typing (Xie et al., 2023), ratings (Gao et al., 2020; Bohn and Ling, 2021), gaze (Bohn and Ling, 2021)) or simply post-editing (e.g., (Lai et al., 2022; Moramarco et al., 2021)).
+A few works have supported summarization based on user-selected points from the source artifact (Bayatmakou et al., 2022; Yan et al., 2011; Zhang et al., 2011; Wang et al., 2023a, b), but the points available for selection are ranked or clustered by the system without the same narrative structure as the source.
 Another work proposed but did not implement or evaluate model guidance through in-context content selection (Gehrmann et al., 2019).
 Our approach allows model guidance directly grounded in the input source while maintaining the source narrative structure.
-Providing more direct control of input content is useful when the writer has significant knowledge of the source and can use reminders of its contents to select what to include in the summary.
-We focus on summarizing one’s own content, unlike most prior work, with the exception of work on summarizing medical notes that used the post-editing technique (Moramarco et al., 2021; Cai et al., 2022).
-Thus, we take advantage of the writer’s knowledge of the source document by having them select content in context of the source’s layout.
-There are several works on human-LLM summarization of long documents (Zhang et al., 2011; Gehrmann et al., 2019) and some for scientific long documents in particular (Zarinbal et al., 2019; Bayatmakou et al., 2022; Zhang et al., 2023b).
-Long documents provide a unique challenge for human-LLM summarization, as there is a great deal of information for the human to digest.
-As writing research-paper blog posts entails summarizing portions of a research paper, this work seeks to support human-LLM long-document summarization.
-2.2. Human-LLM Writing
-While LLMs have impressive capabilities with respect to quickly generating large amounts of text, interacting with swaths of text can be difficult to manage without additional affordances.
-Popular chatbot LLMs such as chatGPT (OpenAI, [n.
-d.]) and BARD (Google, [n.
-d.]) offer some options for the user to guide their output beyond general prompting.
-Both provide example inputs for the user to select.
-chatGPT includes the option to provide custom instructions for the questions ”What would you like chatGPT to know about you to provide better responses?” and ”How would you like chatGPT to respond?”.
-BARD allows the user to modify its generated text to be shorter, longer, simpler, more casual, more professional, or modified based on a custom prompt.
-Tools from prior research also provide users with means to guide LLM writing beyond general prompting, such as keyword specifications (Zhang et al., 2023b), style specifications (Ghazvininejad et al., 2017; Coenen et al., 2021), and input modification (Zhang et al., 2011).
-With respect to planning text, a unique aspect of our work is providing an LLM-generated outline for source content selection.
-There are a couple works in human-LLM writing that have utilized outlines, but they focused on outlining output content rather than input content.
-Zhang et al.
-introduce the tool VISAR, a human-AI argumentative writing assistant that allows users to iterate on their draft with adjustments to a visual outline of the draft (Zhang et al., 2023a).
-Dang et al.
-introduce a text editor that generates summaries of each paragraph written in the editor to comprise a reverse outline of the writing (Dang et al., 2022).
+Providing more direct control of input content is useful when the writer has significant knowledge of the source and can use reminders of its contents when selecting what to include in the summary.
+We focus on summarizing one’s own content, unlike most prior work; the exception is work on summarizing medical notes, which used the post-editing technique (Moramarco et al., 2021; Cai et al., 2022).
+We take advantage of the writer’s knowledge of the source document by having them select content within the context of the source’s layout.
+2.2. Outlines in Human-LLM Writing
+With respect to planning blog post text, a unique aspect of our work is providing an interactive LLM-generated outline for source content selection.
+There are some works in human-LLM writing that have utilized outlines for writing, but they focused on outlining output content rather than input content.
+For example, Zhang et al.
+introduced the tool VISAR, a human-AI argumentative writing assistant that allows users to iterate on their draft with adjustments to a visual outline of the draft (Zhang et al., 2023a), and Lu et al.
+created a tool to support writing workplace social media posts that includes a feature to help generate an outline for the post (Lu et al., 2024).
+Most related to our work, Dang et al.
+introduced a text editor that generates summaries of each paragraph written in the editor to comprise a reverse outline of the writing (Dang et al., 2022).
 The summaries then become affordances to help writers revise their writing (e.g., restructuring the writing).
-Our work also generates a reverse outline but as an affordance to guide summary generation based on the input content.
+Our work also generates a reverse outline, but it is an outline of the input paper for users to interactively guide blog post generation.
 For summarizing a long source document, an outline is particularly useful, as reviewing the entire source may require substantial time and effort.
-With respect to revising text, like BARD (Google, [n.
-d.]) and other LLM-powered writing tools (NotionAI, [n.
-d.]), our final tool allows writers to utilize preset modifications, but some of these are set to be more specific to blog post writing.
-Closely related to our work, Laban et al.
-present an editing interface to help users revise their writing with LLM edit suggestions that can be integrated into their writing (Laban et al., 2023).
-This editing interface focuses solely on the revising step and on smaller edits than our tool.
+2.3. Preset, Customizable LLM Instructions in Human-LLM Writing
+For drafting blog post text, our tool presents LLM instructions preset to generate four unique research-paper blog post sections; the user may then adjust these section-specific instructions in multiple manners to draft text that better aligns with their needs.
+Prior works also break text into chunks to support drafting and revising writing.
+For example, Laban et al.
+present an editing interface that provides edit suggestions for relatively small chunks of writing (Laban et al., 2023), while Dhillon et al.
+provide next-sentence and next-paragraph suggestions to continue writing (Dhillon et al., 2024).
+There are also prior works on human-LLM writing that, like ours, allow users to provide customized instructions to LLMs for drafting text (e.g., (Lim et al., 2024; Coenen et al., 2021)).
+Our customizable LLM instructions for drafting text are specifically designed to support writing common sections of research-paper blog posts.
+For revising blog post text, our tool provides preset macros as well as the ability to perform custom text modifications.
+Several existing tools (e.g., (Google, [n.
+d.]; NotionAI, [n.
+d.]; Lu et al., 2024; Coenen et al., 2021; Reza et al., 2024)) also allow users to modify their text with macros for actions such as adjusting the text’s length, simplicity, or professionalism, and some allow custom modifications.
+However, our tool’s macros are designed specifically for blog post writing.
 2.4. Human-LLM Scientific Writing
@@
 Gero et al.
-present a system for generating scientific ”sparks”, or inspiring sentences (Gero et al., 2022).
+present a system for generating scientific “sparks”, or inspiring sentences (Gero et al., 2022).
 Long et al.
@@
 investigate how science writers can generate extended metaphors for scientific ideas with the help of an LLM (Kim et al., 2023).
-The couple works that have explored human-AI scientific summarization do not allow for in-context selection of source content but rely on selection of concepts (Zhang et al., 2023b) or out-of-context sentences extracted based on queries (Bayatmakou et al., 2022).
-These previously studied forms of human-LLM scientific writing are the length of a sentence or paragraph.
+The few works that have explored human-AI scientific summarization do not allow for in-context selection of source content but rather allow users to select concepts (Zhang et al., 2023b), out-of-context sentences extracted based on queries (Bayatmakou et al., 2022), suggested summary content (Lim et al., 2024).
+Lim et al.’s Q-and-A research-paper summarization tool AQUA also allows users to prompt the LLM with relevant questions (Lim et al., 2024).
+Meanwhile, Google’s tool NotebookLM for general human-AI document summarization does not permit any content selection (Google, [n.
+d.]).
 Based on prior work (Mehlenbacher, 2019), popular resources (Kudos, [n.
@@
 d.]; CMU, [n.
-d.]; Medium, [n.
-d.]), and feedback from three communications experts, we determined a set of guidelines for writing research-paper blog posts, which indicate that the shortest blog posts are recommended to be around 500 words in length (Medium, [n.
-d.]; Kudos, [n.
-d.]; Patrick Dunleavy, [n.
-d.]), making blog posts a long-form output.
-Because research-paper blog posts are longer than a sentence or a paragraph, they require coordinating more writing sub-tasks, such as rewriting a sentence about a method to be easier for a broader audience to understand or summarizing a paper’s results section.
+d.]), and feedback from three communications experts, we determined a set of guidelines for writing research-paper blog posts, which indicate that the shortest blog posts are recommended to be around 500 words in length, making blog posts a long-form output.
+Because research-paper blog posts provide long-form scientific communication, drafting them requires LLM instructions unique to common blog post sections, and revising them requires balancing the writing style for accuracy, engagement, ease of understanding, and concision.
+Most of the previously cited forms of human-LLM scientific writing above are the length of a sentence or paragraph.
+The exceptions are AQUA’s Q-and-A paper summary (Lim et al., 2024) and NotebookLM’s many summary forms (e.g., FAQ, study guide, briefing doc, podcast).
+However, AQUA’s scaffolding is limited to suggested questions and answers as well as a space for user-provided additional prompting, and NotebookLM provides no scaffolding to support customization of their document summaries.
+Furthermore, our work is unique in its dedication to supporting mixed-initiative writing of research-paper blog posts.
 2.5. Scientific Blog Posts
 Scientists can engage with broad community through various channels, including press releases, magazines, journals, and tweets (August et al., 2020; Gero et al., 2022).
-This work focuses on scientific blog posts, which offers increased recognition and transparency of one’s work, writing practice, editorial freedom, and public engagement (Jarreau, 2015; Blanchard, 2011; Zou and Hyland, 2019).
-There are several forms of scientific blog posts, including academic commentary, free commentary, political commentary, and mediation of research to laypersons (Mahrt and Puschmann, 2014).
-Here, we focus on blog posts with the primary goal of communicating information from a research paper to scientists in the broad area of the paper (e.g., computer science, psychology).
-3. Formative Study and System Iteration
-We conducted a formative study to identify challenges that people encounter in working with an LLM-powered technology probe for writing blog posts about their research papers, and corresponding design goals for our final LLM tool.
-3.1. Participants
-Via academic social networks, we recruited 6 participants for the formative study and compensated them with $25 for the 45-minute study.
-Participants were PhD students, a population frequently tasked with writing blog posts for papers they have authored.
-Five worked in computer science and one in information science.
-To circumvent initial time required for getting used to LLMs themselves, we recruited people who already had some experience of working with LLMs in general (6-10 sittings: 1; ¿10 sittings: 5).
-We also required participants to have previously written research papers (3-10 papers: 3; ¿10 papers: 3), since the purpose of such a tool is to write blog posts for one’s own research papers.
-Participants had a mixture of experience with writing research-paper blog posts (0 posts: 4; 1-2 posts: 2).
-Also, for sufficient participant motivation, we made sure that all participants were more interested than not (at least 5 on a 7-point Likert scale where 7 indicates the highest interest) in writing blog posts for their research papers.
-3.2. Study Procedure
-The study consisted of 3 iterations, each of which involved sessions with 2 participants.
-Based on participants’ interactions and feedback, we updated the probe after each iteration for better insight into remaining challenges.
-Prior to each study session, the participant completed a participation agreement.
-The session coordinator recorded and transcribed each session in Google Hangouts.
-The coordinator asked the participant to assume their audience was laypeople, as a common advantage of writing scientific blog posts is to communicate with a broader audience than the paper does (Jarreau, 2015; Blanchard, 2011).
-The participant then used a link to access the technology probe initialized for one of their research papers that they provided.
-In order to see how participants would intuitively use the probe, the coordinator purposefully did not initially explain how to use it.
-For 30 minutes, the participant used the probe to write as much as possible for a blog post about their research paper.
-We asked participants to think aloud as they worked (Van Someren et al., 1994).
-The coordinator occasionally asked follow-up questions about what the participant said aloud.
-Participants could ask for help when they did not understand an aspect of the tool.
-Once 30 minutes had passed, the coordinator engaged the participant in a semi-structured interview.
-They asked the participant about the difficulties and benefits of using the probe to 1) navigate the paper and outline, 2) generate and edit text for a blog post section, and 3) edit the blog post layout.
-Finally, they asked for the participant’s thoughts on receiving different types of feedback on their blog post, but this topic was left for future work.
-3.3. Initial Technology Probe
-In the first iteration of the technology probe interface (Appendix 28), the left side consisted of blog post sections pre-populated with an LLM-generated draft for drafting and revising the blog post.
-The right side contained the paper text with selectable paragraphs and an LLM-generated paper outline with bullet points for planning.
-Under the currently selected blog post section, there was an LLM-supported text generator that contained buttons to generate section text from the selected paragraphs and bullet points as well as to expand and condense the text in the blog post section editor.
-3.4. Design Goals and Tool Iteration
-After each iteration of the formative study, the first author conducted inductive thematic analysis (Braun and Clarke, 2006) of the participants’ think-aloud and interview comments on their experience with the probe.
-Participants described difficulties in working with the probe that, aside from general usability issues, may be classified under two overarching challenges.
-3.4.1. Challenge 1: Planning Paper Content to Use
-An academic paper is a long document with dense information, so we expected that identifying information to include in the blog post would be a challenge.
-For this reason, our technology probe parsed the paper text and generated a paper outline with an LLM in order to support navigating and utilizing the paper.
-Still, participants faced difficulties with finding and selecting information from the paper.
-Our associated design goal was to help writers plan their writing by facilitating review of a long document’s contents and selection of essential information to include in the writing (DG1).
-Although participants faced issues with the paper outline, four participants expressed that they liked having the paper outline, and four appreciated the ability to select paper paragraphs or bullet points from which to generate text.
-Because participants overall liked and used the paper outline for identifying and using essential information from the paper, we iterated on this affordance to address DG1.
-We identified the following requested functionality associated with this design goal:
-Transparent Connection Between Paragraphs and Bullet Points.
-In the first iteration, P2 noted confusion around how the bullet points in the outline corresponded to the original paper text, despite the existence of a button to the left of each set of bullet points that revealed the associated paragraph when clicked.
-We therefore added a bright line to the left of each set of bullet points to better visualize its grouping and its connection to the corresponding paragraph (see Figure 2).
-In the second iteration, P3 expressed uncertainty about how the selection of paragraphs and bullet points were related.
-For example, they were surprised that a paragraph was selected but only one of three corresponding bullet points was selected.
-During the formative study, the selection logic was as follows.
-If a bullet point is selected, the associated paragraph needs to be selected, but if a paragraph is selected, no bullet point needs to be selected.
-This was done in an attempt to make it clearer that the associated paragraph was given to the model as context in addition to any bullet points selected.
-However, this logic seemed to confuse the participants.
-Filter for Selected Content.
-In the third iteration, P6 said that they were not sure if they had selected everything that they wanted to select.
-Throughout all the iterations, the probe only communicated the number of selected paragraphs and bullet points.
-P6 desired a way to view only the selected paragraphs and bullet points.
-3.4.2. Challenge 2: Drafting and Revising with Recurring LLM Instructions
-The second challenge was attempting to provide the LLM instructions for generating and modifying text that were not well-supported by the probe.
-Some of these instructions were frequently desired, but the participant had no reasonable way to input them or, once directly prompting the LLM was available, had to determine a prompt to fit their needs.
-Participants also faced confusion regarding the prompts being used in the back-end.
-Our corresponding design goal was therefore to help writers draft and revise their writing by transparently streamlining commonly desired instructions to the LLM for generating and modifying text (DG2).
-We identified the following requested functionalities under this design goal:
-Transparent Prompt Input.
-In the second iteration, P3 liked a default blog post section and was confused that it appeared significantly different from the new version they generated without adjusting the pre-selected bullet points.
-For one, there was a difference in the requested length in the prompt for the initial-draft section versus the live-generation section.
-Also, the participant was not aware of the section-specific instructions that were provided to the model for the initial-draft section.
-To provide access to default settings, which may often be desirable, we updated the probe to show these section-specific instructions as preset (but adjustable) custom instructions.
-While the model was always prompted to add 1 to 2 sentences in the third iteration, P5 was confused regarding if and how much the text would expand after clicking the ”expand” button multiple times.
-Meanwhile, P6 commented that they were unsure if the model had context for the blog post sections above the one for which they were generating text.
-Custom Instructions.
-In the first iteration, our design probe did not support direct custom instructions.
-Users could only generate or modify the text using the pre-defined operations.
-P1 tried to work around this limitation by giving specific instructions to the LLM in the blog post section editor, custom bullet points, and custom instructions.
-The participant seemed unaware that the model was not taking these instructions into account.
-To provide users with more control over the generated text, we added a field for custom instructions to the generator box.
-New Section Generation. In the second iteration, P4 noted wanting to generate entire new sections with the model, perhaps based on the section header.
-Flexible Input for Text Modification.
-In the first iteration, P2 expected and wanted the ”condense” button to iterate on the generated text rather than the editor text, so we updated the ”expand” and ”condense” buttons to do that.
-However, in later iterations, P4 expressed interest in the model expanding on something they wrote without reference to the paper, and P6 expressed interest in rewriting portions of the original paper in different manners.
-Preset Modification Actions.
-We observed participants attempting to rewrite text in manners that seemed applicable to many research-paper blog posts.
-For instance, P1 wanted a button to make text more approachable for their audience of laypeople, while P3 rewrote text that seemed ”too lofty” in its claims.
-For the third iteration, we therefore added buttons for such modifying actions, which included ”simpler terms”, ”less dramatic”, and ”more dramatic.” When the participant clicked the button, the modification would be applied to the last generation.
+This work focuses on scientific blog posts, which offer increased recognition and transparency of one’s work, writing practice, editorial freedom, and public engagement (Jarreau, 2015; Blanchard, 2011; Zou and Hyland, 2019).
+There are several forms of scientific blog posts, such as academic commentary, free commentary, and mediation of research to laypersons (Mahrt and Puschmann, 2014).
+Here, we focus on blog posts with the primary goal of communicating information from a research paper to people unfamiliar with the paper’s specific research topic.
 3. The Papers-to-Posts System
-4.1. Paper Outline and Initial Draft Generation
-The paper text was extracted from the paper PDF using a public tool for converting scientific PDFs to HTML.111https://papertohtml.org We manually adjusted the parsed HTML so that the section headers were correct, no section or subsection was completely missing, and only the main paper text was included (i.e., no appendix, references, or footnotes).
-Within the paper text box, participants could click a button to see a paper outline generated for the paper text.
-To obtain this paper outline, we used an LLM to generate 1 to 3 bullet points to summarize each paragraph in the paper, depending on the length of the paragraph.
-To generate the initial blog post draft for the tool, all of the paper’s bullet points were given to the model along with the request to select 10 bullet points relevant to the initial four blog post sections: introduction, methods, results, and conclusion.
-To generate each draft section, the model was provided the selected bullet points, their associated paragraphs, the portion of the blog post that had already been generated, and section-specific guidelines based on prior work (Mehlenbacher, 2019), popular resources (Kudos, [n.
+We developed Papers-to-Posts based on key design goals and pilot study feedback, which helped us to identify system features to add for increased transparency, flexibility, and ease of use.
+This section describes the design goals for Papers-to-Posts and provides a walkthrough of the resulting system’s four main areas.111A walkthrough video of the tool may be found in the supplementary materials.
+3.1. Design Goals
+We determined the design goals for Papers-to-Posts based on guidelines for academic blog posts from popular resources (Kudos, [n.
 d.]; UWaterloo, [n.
@@
 d.]; CMU, [n.
-d.]; Medium, [n.
-d.]), and feedback from three communications experts.
-We found that this process of generating one draft section at a time with the context of prior sections helped make the draft more coherent while avoiding overwhelming the model with instructions for many sections.
-For both the initial draft and user generations, the model was provided the associated paragraphs for any bullet points in order to ground the model in the actual paper.
-We used GPT-4 for all LLM tasks except for extracting bullet points from a paragraph, which we produced using GPT-3.5-Turbo after observing that it produced shorter bullet points that were quicker to read.
-It was also many times cheaper for this token-intensive generation task.
-4.2. User Interface
-Developed over the course of the formative study, the Papers-to-Posts system consists of two main panes split vertically.
-On the left is the pane for writing the blog post (Figure 2a), and on the right is the area for generating and modifying text.
-The blog post is divided into sections, which can be named, reorganized, and deleted.
-Each section is connected to a unique LLM state, so the history of all inputs and outputs for text generation and modification are stored separately for each section.
-The initial blog post draft is generated as described in Section 3.3.
-The blog post section currently being edited is highlighted in purple to indicate that it is the section for which text is generated.
-Using the controls on the top of the left pane, users can view the blog post’s word count, a preview of the full blog post, and tips for writing blog posts based on prior work (Mehlenbacher, 2019) and popular resources (Kudos, [n.
+d.]), prior work (Mehlenbacher, 2019), and feedback from three communications experts, recruited from our institutions.
+There are multiple guidelines on the content and format for different sections of the academic blog post (e.g., introduction, results) (Kudos, [n.
 d.]; UWaterloo, [n.
@@
 d.]; CMU, [n.
-d.]; Medium, [n.
-d.]), as well as feedback from 3 communications experts.
-There is a button to submit the blog post at the bottom.
-The writer can also add a new section to the blog post using the plus-sign button next to each blog post section.
-When the writer clicks this button, they can either create a blank section or generate a section based on the section header they provide.
-To generate a new section, we prompt GPT-4 to select 10 bullet points for the section based on the provided header and to generate text for the section focused on those bullet points (Appendix A).
-The pane for generating and modifying text on the right has three areas.
-The first is the planning-support area (Figure 2b), which contains the Paper Outline to address DG1.
-The Paper Outline supports writers in planning their blog post, through review of the paper’s contents and selection of essential information to include in the post.
-The paper view at the top of the pane offers the writer two views of the paper in two tabs: Paper Text and Paper Outline.
-Under the Paper Text tab, the writer may view the original paper text, divided into paragraphs.
-The paper is parsed into collapsible sections.
-The figures and tables are presented in a separate final section of the paper.
-Their captions can be copied for pasting in the editor so that the writer knows where they should go in the final blog post.
-Under the Paper Outline tab, the writer may view the paper outline, divided into sets of bullet points associated with the original text’s paragraphs.
-The writer may view a bullet point set’s corresponding paragraph by clicking the button to the left, and the reverse is true under the Paper Text tab.
-The outline is generated, and the pre-selected bullet points for each blog post section are displayed.
-The writer may select or deselect any bullet points or paragraphs that they want the LLM to use in generating text for the blog post section on which they are working.
+d.]; Mehlenbacher, 2019).
+With respect to the blog post’s tone, guidelines suggest avoiding jargon (Kudos, [n.
+d.]; UWaterloo, [n.
+d.]; Patrick Dunleavy, [n.
+d.]; CMU, [n.
+d.]), sensationalism (noted by a communications expert), and wordiness (Kudos, [n.
+d.]; UWaterloo, [n.
+d.]; Patrick Dunleavy, [n.
+d.]).
+The resulting design goals are:
+DG1: Help researchers to plan what paper content to include in their blog post.
+DG2: Help researchers to draft their content in sections amenable to a research-paper blog post.
+DG3: Help researchers to revise the blog post’s tone to be scientifically accurate but also concise, easy to understand, and engaging.
+3.2. Walkthrough of Start-Up Step
+3.2.1. Frontend
+In the Blog Post Area (Figure 2a), the user starts by reviewing the initial blog post draft.
+The blog post draft is divided into sections.
+The user clicks on whichever section they want to work on first, and the tool highlights that section for reference.
+At the top, the user views the blog post’s word count, a preview of the full blog post, and tips for writing blog posts, which are based on popular resources (Kudos, [n.
+d.]; UWaterloo, [n.
+d.]; Patrick Dunleavy, [n.
+d.]; CMU, [n.
+d.]), prior work (Mehlenbacher, 2019), and feedback from three communications experts.
+3.2.2. Backend
+In order to obtain the initial blog post draft for a given research paper, we completed a few steps (Figure 1a).
+In the first step, we used a public tool for converting scientific PDFs to HTML222https://papertohtml.org to extract the paper text from the paper PDF.
+For the lab study, we then manually adjusted the parsed HTML so that the section headers were correct, no section or subsection was completely missing, and only the main paper text was included (i.e., no appendix, references, or footnotes).
+We also manually retrieved the tables and figures from the parsing.
+For the real-world deployment study, the parsing was automatic to better align with a real-world setting.
+For the second step, we used GPT-3.5 Turbo to generate between one and three bullet points to summarize each paragraph in the parsed paper, depending on the length of the paragraph.
+While we used GPT-4 (GPT-4-0125-preview in the deployment study) for all of Papers-to-Posts’s other tasks, we decided to use GPT-3.5-Turbo for this task for two reasons: 1) it produced shorter bullet points that were quicker to read, and 2) it was many times cheaper for this token-intensive generation task.
+All together, the generated bullet points formed a paper outline.
+For the third step, we provided the LLM with all of the paper’s bullet points along with the request to select 10 bullet points relevant to the initial four blog post sections: introduction, methods, results, and conclusion.
+We requested 10 bullet points per blog post section in order to provide the model sufficient content without overwhelming it.
+For the fourth step, we prompted the LLM to generate each section in succession using the selected bullet points for that section, the paragraphs associated with the selected bullet points (to better ground the model), the portion of the blog post that had already been generated, and section-specific guidelines based on popular resources (Kudos, [n.
+d.]; UWaterloo, [n.
+d.]; Patrick Dunleavy, [n.
+d.]; CMU, [n.
+d.]), prior work (Mehlenbacher, 2019), and feedback from three communications experts.
+We observed that this process of generating one draft section at a time with the context of prior sections helped to make the draft more coherent while avoiding overwhelming the model with instructions for multiple sections.
+The model was prompted to make each section around 125 to 250 words and around one to three paragraphs.
+The prompts for generating the initial draft may be found in Appendix B.2.
+An example initial draft and associated selected bullet points is in Appendix F.1 and F.2.
+3.3. Walkthrough of Planning Step
+3.3.1. Frontend
+In the Planning-Support Area (Figure 2b), the user works with the tool to plan the blog post section on which they are currently working (DG1).
+This area has two tabs labeled Paper Outline and Paper Text.
+Under the Paper Outline tab, the user views the paper’s outline, divided into sets of one to three bullet points; each set is associated with a paragraph from the original text.
+The user views a bullet point set’s corresponding paragraph by clicking the button to the left, and the reverse is true under the Paper Text tab, where the user views the original paper text divided into paragraphs.
+Under both tabs, the information is divided into collapsible sections based on the paper’s highest-level sections.333In the lab study, the tool presented the figures and tables in a final collapsible section.
+Their captions could be copied for pasting in the editor to help the user plan where to put figures in the final blog post.
+We did not include this feature in the deployment study, as the automatic paper parsing did not extract this information.
+Under the Paper Outline tab, the user reviews the tool’s pre-selected bullet points for the current blog post section by scrolling through the full outline or clicking the “Show Selected Only” button below the paper area.
+When this filter button is activated and the user is under the Paper Text tab, if a paragraph is not selected but a corresponding bullet point is selected, the tool notes that “A corresponding bulletpoint(s) is still selected”; the opposite is true for the Paper Outline tab.
+The user clicks the “Close All Paper Sections” button below the paper area if they want to get a higher-level view of the paper.
+The user then adjusts the selection of bullet points or paragraphs for the LLM to use in generating text for the current blog post section.
 Selecting a paragraph does not affect the selection of associated bullet points, and vice versa.
-On the back-end, the model still receives the context of the associated paragraph for each bullet point selected.
-There are buttons below the paper view to clear all selected bullet points and paragraphs, filter for only selected bullet points and paragraphs, and close or open all the paper sections.
-When the filter button is activated, if the writer is in the Paper Text tab, if a paragraph was deselected but a corresponding bullet point is still selected, the tool notes that ”A corresponding bulletpoint(s) is still selected,” and vice versa.
-Below the planning-support area is the drafting-support area (Figure 2c).
-There are several inputs to address DG2 by transparently streamlining commonly desired instructions to the LLM for generating text.
-First, there is a menu to select the desired length for the generated text.
-The default length is labeled as ”auto,” and the other options are 1 sentence, 1 paragraph, or a few paragraphs.
-The ”auto” length is the same as for the initial LLM draft sections (around 125 to 250 words) There is next a toggle turned on by default for the LLM to use the context of blog post sections above the one currently being edited.
-The writer can optionally utilize text fields to provide custom bullet points, custom instructions, and starting text to the LLM.
-For the initial introduction, methods, results, and conclusion sections of the blog post, the custom instructions field is pre-filled with the section-specific instructions used to generate those sections.
-The text generator’s ”generate” button utilizes the inputs to generate text for the current blog post section using GPT-4.
-The tool also provides a revising-support area (Figure 3d, located below the drafting-support area in the tool), which addresses DG2 by transparently streamlining commonly desired instructions to the LLM for modifying text.
-First, there is a text field for inputting the text to be modified.
-This can be anything, including text from the blog post draft, the paper, or a generation.
-There is then a menu to select the desired length for the modified text.
-The options are the same as for generated text, with ”auto” being unique to each modifying action (see Appendix A).
-Next, there is a toggle, on by default, for the LLM to use the context of blog post sections above the one currently being edited.
-There is also a toggle turned on by default for the LLM to base modified writing on selected paragraphs and bullet points.
-Underneath the toggles, there is a text field for providing any custom instructions for the modification action.
-At the bottom, there are 5 preset modification buttons: ”expand”, ”condense”, ”simpler terms”, ”less dramatic”, and ”more dramatic”.
-There is also a ”custom modification” button.
-Aside from the ”expand” and ”condense” buttons, all of the affordances for generating and modifying text were additions made to support recurring types of LLM instructions observed in the formative study.
-In the lab study, participants utilized each affordance to some degree.
-5. Lab Study
-We conducted a lab study to determine if and how the LLM tool resulting from the formative study may support researchers in writing blog posts about their research papers.
-5.1. Hypotheses
-We hypothesized that, compared to a baseline consisting of an LLM draft and the ability to prompt the LLM, Papers-to-Posts scaffolding and affordances would provide participants a better experience.
+If the user wants to reset the content selection, they click the “Clear All Selected” button below the paper area.
+3.3.2. Backend
+As shown in Figure 1b, all planning actions (as well as drafting and revising actions) are connected to the user’s selected blog post section.
+In other words, the state and history of all inputs and outputs for text generation are stored separately for each section.
+This is somewhat similar to the fragmented document history proposed by Buschek (Buschek, 2024).
+3.4. Walkthrough of Drafting Step
+3.4.1. Frontend
+In the Drafting-Support Area (Figure 2c), the user works with the tool to synthesize text for the blog post section on which they are currently working (DG2).
+The user reviews and adjusts the LLM’s section-specific instructions for generating text.
+At the top of the Drafting Support Area, the user changes the desired length of the generated text to their liking.
+The length defaults to “auto” and may be updated to one sentence, one paragraph, or a few paragraphs.
+Below, the user notes the toggle that is turned on by default.
+This toggle lets the user know that the context of blog post sections above the highlighted one is provided to the LLM in order to generate text for this section, and they may turn this toggle off if they do not want this context to influence the generated text for this section.
+In the “Custom Bulletpoints” text field, the user enters any additional bullet points that they want their generated text for the highlighted section to cover.
+In the “Custom Instructions” text field, the user adjusts the instructions for the highlighted section’s generated text.
+For each of the initial blog post sections, this box is pre-filled with section-specific instructions, which are based on the blog post guidelines we collected.
+In the text field labelled “Starting Text to Continue from”, the user optionally provides text that should be at the start of the generated text.
+Finally, the user clicks the “generate” button to generate text for the highlighted section, which appears below.
+The user then copies and pastes the generated text into the highlighted blog post section and edits it as they see fit.
+The user may peruse a history of their content selection, LLM instructions, and text generations for the highlighted section.
+The user also drafts text for the blog post by adding a new section.
+In the Blog Post Area, the user clicks the plus-sign button next to an existing blog post section, and the tool places the new section below the existing one.
+The tool gives the user the option to either create a blank section or generate a section based on the section header that they provide.
+If the user decides the new section should be in another location within the blog post, they use the section’s associated up and down arrow buttons to move the section to the desired point.
+For this new section, the user may return to the planning step or adjust the LLM’s instructions for generating text.
+If the user decides not to include the section in the blog post, they delete it using the trash button.
+3.4.2. Backend
+As shown in Figure 1c, the LLM (GPT-4 in lab study; GPT-4-0125-preview in deployment) generates text for the highlighted blog post section with the user’s adjusted selection of bullet points and paragraphs as well as their adjusted section-specific instructions (described in Section 3.4.1).
+With respect to the selected bullet points, the model receives the context of the associated paragraph for each bullet point selected, even if the associated paragraph was not selected by the user.
+This is to ensure the model’s generation is grounded in the actual paper text.
+When the user desires the “auto” length for generated text, the model is instructed to generate text that is 125 to 250 words and between one and three paragraphs.
+To generate a new section, we prompt the LLM to select 10 bullet points for the section based on the provided header and to generate text focused on those bullet points.
+Again, the model is provided the context of the bullet points’ associated paragraphs.
+The prompts for drafting text may be found in Appendix B.4, and example outputs are in Appendix F.1 and F.2.
+3.5. Walkthrough of Revising Step
+3.5.1. Frontend
+In the Revising-Support Area (Figure 2d), the user works with the tool to revise the blog post section on which they are currently working (DG3).
+At the top of this area, the user pastes the text that they want to modify in the “Writing to Modify” text field.
+This text may be anything, including text from the blog post draft, the paper, or a previous model generation.
+The user then adjusts the desired length of the modified text if need be; the options are the same as in the Drafting-Support Area (Section 3.4).
+Also as in the Drafting-Support Area, the user may adjust the first toggle, which is on by default to indicate that the blog post sections above the one being edited are being provided as context to the LLM.
+The user then notes the second toggle, which is on by default to indicate that the LLM will utilize the currently selected paragraphs and bullet points to generate the modified text.
+At the bottom of the Revising-Support Area, there are five preset modification buttons to 1) expand the text, 2) condense the text, 3) rewrite the text in simpler terms to be more understandable to a layperson, 4) rewrite the text in a less dramatic tone to align better with the unadorned language of scientific writing, and 5) rewrite the text in a more dramatic tone to better capture readers’ attention.
+There is also a “custom modification” button.
+If the user wants to use the “custom modification” button, they enter their custom instructions for the text modification in the text field above.
+If the user wants to use a different modification button, they may still provide custom instructions for the LLM to take into consideration.
+The user copies and pastes the modified text into the highlighted blog post section and edits it as they see fit.
+The user may peruse a history of their LLM instructions and text modifications for the highlighted section.
+3.5.2. Backend
+As shown in Figure 1d, the LLM (GPT-4 in lab study; GPT-4-0125-preview in deployment) modifies text for the highlighted blog post section with the user’s selected text snippet and their adjusted instructions for the modification (described in Section 3.5.1), including the selected text modification.
+The desired length “auto” is unique to each modifying action.
+The prompts for revising text may be found in Appendix B.4, and example outputs are in Appendix F.3.
+4. Evaluation Studies
+4.1. Lab Study
+We conducted a lab study to determine if and how Papers-to-Posts may support researchers in writing blog posts about their research papers.
+4.1.1. Hypotheses
+We hypothesized that, compared to a baseline consisting of an LLM draft and the ability to prompt the LLM, Papers-to-Posts would provide participants a better experience and outcome.
 Our individual hypotheses were as follows:
@@
 H3: Papers-to-Posts leads to lower cognitive load than the baseline.
-H4: Papers-to-Posts leads to lower task completion time than the baseline in generating a blog post that the writer would be comfortable posting publicly.
-5.2. Participants
+H4: Papers-to-Posts leads to less time spent generating a blog post that the writer would be comfortable posting publicly, in comparison to the baseline.
+4.1.2. Participants
 We recruited 20 participants (M: 14, W: 6) via academic social networks and institutional mailing lists and compensated them with $83.33 USD for 2.5 hours of their time.
@@
 In order to simulate a naturalistic setting, participants were required to have authored at least 2 research papers for use in the study (2 papers: 1, 3-10 papers: 14, ¿10 papers: 5) and to be more interested than not (at least 5 on a 7-point Likert scale where 7 indicates the highest interest) in writing blog posts for their research papers.
-5.3. Study Conditions
-The treatment condition involved interacting with Papers-to-Posts (Figure 2), which is described in Section 4.
+4.1.3. Study Conditions
+The treatment condition involved interacting with Papers-to-Posts which is described in Section 3.
 The tool in the baseline condition was designed to be a simplified version of Papers-to-Posts without affordances developed to address the design goals (Figure 4).
 The baseline provided input paper text, sourced from the PDF and divided into paragraphs and collapsible sections as in Papers-to-Posts.
-The baseline also provided the ability to prompt GPT-4 with any instructions.
+The baseline also provided the ability to prompt GPT-4 with any instructions and copy its output.
 A history of generations and their associated instructions was retained for the user’s perusal.
-Lastly, the baseline had an editor for writing the blog post, which was pre-populated with an LLM draft generated by GPT-4-32k using the entirety of the input paper.
+Lastly, the baseline had an editor for writing the blog post, which was pre-populated with an LLM draft generated by GPT-4-32k using the entirety of the input paper.444An example of a baseline initial draft is in Appendix E.
 GPT-4-32k was used rather than GPT-4 because the context window for GPT-4 was insufficient for longer research papers.
-The prompt was designed to mirror the prompt for Papers-to-Posts’s initial draft sections, so the resulting draft also had introduction, methods, results, and conclusion sections.
+The prompt was designed to mirror the prompt for Papers-to-Posts’s initial draft sections (Appendix B.1), so the resulting draft also had introduction, methods, results, and conclusion sections.
 However, the prompt did not contain section-specific instructions.
-The full prompt can be found in Appendix A.
 In each condition, the tool was set up for one of the participant’s research papers submitted in the recruitment survey.
@@
 There was no significant difference in publication dates between conditions (Paired-Samples t-Test, t(19)=0.90, p=n.s.).
-The papers were parsed as described in Section 3.3.
 The average paper length across conditions differed by less than 10% (baseline: 8452 words, treatment: 7742 words).
-5.4. Study Procedure
-The study was within-subjects, so each participant completed two study sessions, one in the treatment condition and one in the baseline condition, in randomized order and following a counter-balanced design.
-The participants’ selected papers were assigned randomly to conditions.
+4.1.4. Study Procedure
+The study was within-subjects, so each participant completed two study sessions, one in the treatment condition and one in the baseline, in randomized order and following a counter-balanced design.
+The participants’ selected papers were also assigned to conditions in randomized order.
 Each study session lasted up to about 75 minutes.
-To provide sufficient flexibility for recruited participants, they can select two study sessions within a week of each other, but the majority of sessions were a day apart (1 day: 13, 2 days: 3, 3 days: 3, 4 days: 1).
-The condition order was counter-balanced within the group of participants with sessions 1 day apart and within the group of participants with sessions over 1 day apart.
+For sufficient flexibility, participants selected two study sessions within a week of each other, but the majority of sessions were a day apart (1 day: 13, 2 days: 3, 3 days: 3, 4 days: 1).555The condition order was counter-balanced within the group of participants with sessions one day apart and within the group of participants with sessions over one day apart.
 Participants completed a participation agreement before the study.
-We recorded and transcribed each session in Google Hangouts.
-To start, the session coordinator introduced participants to the task of writing a blog post with an AI tool.
-The coordinator then presented a tutorial on the tool.
-After, participants had 2 minutes to interact with the tool in the context of a sample paper that was not their own, and they could ask questions about the tool.
-Participants next opened a tool link initialized with one of their research papers.
-The coordinator asked them to assume that their audience was people in their broad area, which was computer science for all participants.
-We made this assumption because it was the most common audience for whom potential recruits wanted to write blog posts (52 of 60), according to the recruitment survey.
+We recorded and transcribed each session in Google Hangouts.666The lab study session script may be found in the supplementary materials.
+To start, the session coordinator introduced the task and presented a video tutorial on the assigned tool.777See the tutorial scripts in the supplementary materials.
+Participants then had two minutes to interact with and ask questions about the tool in the context of a sample paper that was not their own.
+Next, participants opened a tool link initialized with one of their research papers.
+The coordinator asked them to assume that their audience was people in their broad area (i.e., computer science), which was the most common audience for whom potential study recruits wanted to write blog posts (52/60).
 The coordinator instructed participants to work as if they were not there, but to let them know if any questions about the tool arose.
-They also instructed participants not to use any tools or information outside of the tool provided.
-We made an exception when participants wanted to view the PDF of their paper due to issues working with its parsing in the tool.
-This only occurred twice.
+They also instructed participants not to use any external tools or information.
+We made an exception twice when participants wanted to view their paper’s PDF due to issues with its parsing in the tool.
 Lastly, the coordinator asked participants to let them know if they reached a point at which they would feel comfortable publicly publishing the blog post in its current state.
 Each participant had up to one hour to work on the blog post.
-At the 20-minute and 40-minute marks, the coordinator reminded them to let them know if they reached the point of comfort with publication.
+At the 20-minute and 40-minute marks, they were reminded to let the coordinator know if they reached the point of comfort with publication.
 If and when the participant reached that point, the coordinator asked them to save their draft and continue revising the blog post.
 After the first four study sessions, we noticed that participants were sometimes done early and did not desire to revise further.
-Thus, we added a rule so that the participant would let the coordinator know if they reached a point at which they were proud of the blog post and saw no reason to revise further.
-If and when the participant reached that point, or if an hour had passed, they were asked to submit and moved on to the follow-up survey regarding their experience with the tool.
-The survey consisted of Likert-type questions regarding perceived efficiency, cognitive load, ease of using the tool in relation to the design goals, satisfaction with the tool, satisfaction with the final blog post, perceived control, and perceived ownership.
-There were also short answer questions regarding the difficulties and benefits of using the tool in relation to the design goals.
-If it was the participant’s second session, the survey ended with a few questions directly comparing the two conditions.
+Thus, we added an additional rule that the participant would let the coordinator know if they reached a point at which they were proud of the blog post and saw no reason to revise further.
+If and when the participant reached that point, or if an hour had passed, they submitted their final blog post and moved on to a survey regarding their experience with the tool.
+The survey888The survey may be found in the supplementary materials.
+consisted of Likert-type questions regarding perceived efficiency, cognitive load, ease of using the tool in relation to the design goals, satisfaction with the tool, satisfaction with the final blog post, perceived control, and perceived ownership.999See Appendix A for a discussion of the exploratory results regarding participants’ perceived control and ownership.
+There were also short answer questions regarding the difficulties and benefits of using the tool in relation to the design goals.101010If it was the participant’s second session, the survey ended with a few questions directly comparing the two conditions.
 These were not for analysis but only to help us get a quick sense of the results.
-However, they were not allowed to change their first-session responses.
-At the end of the session, the coordinator told participants to close out of the tool and, if it was their first session, not to use the tool before the next session.
-Each time the participant prompted the LLM, saved, or submitted, a log of the user action along with the full state of the blog post draft was saved for analysis.
-5.5. Results
-We evaluated our 4 hypotheses (Section 5.1) for statistical significance and found that blog post satisfaction is significantly different between the two conditions, while tool satisfaction, cognitive load, and difference in task completion time are not.
-We also investigated how well the design goals were fulfilled, and the extent of participants’ editing.
-We saw evidence that our tool addressed the design goals and increased participants’ satisfaction with control over the tool.
-In addition, we observed that participants did significantly more editing of the initial LLM draft using Papers-to-Posts (we note that this significance test was not part of our initial hypotheses and is therefore exploratory in nature).
-5.5.1. Blog Post Satisfaction
-We utilized the Likert responses (see Figure 5) as a measure of participants’ satisfaction with their final blog post under both conditions.
-Given the Likert nature of the data, we employed a non-parametric test.
-Furthermore, we used a sign test rather than a Wilcoxon signed rank test, which is normally used to analyze non-parametric within-subjects data for two conditions of one factor, due to the violation of that test’s assumption of a symmetric distribution of the paired differences about the median.
-We observed that participants were significantly more satisfied with the final blog post generated with Papers-to-Posts (M=6.00,Q1=5.00,Q3=7.00) as compared to the baseline tool (M=5.00,Q1=5.00,Q3=6.00) (Dependent-Samples Sign Test, S=11, p¡.05).
-The majority of participants (11) were more satisfied in the treatment condition, 7 were equally satisfied in both conditions, and only 2 were more satisfied in the baseline condition.
-Inspecting participants’ Likert responses regarding aspects of their blog post satisfaction, we see that participants particularly demonstrated more satisfaction with the content covered in the Papers-to-Posts blog post than the baseline blog post (Figure 5b).
-The majority of participants (14) were more satisfied with the content under the treatment condition (equally: 2, less: 4).
-Overall, participants were more satisfied with the blog post structure using Papers-to-Posts than the baseline (Figure 5d).
-The plurality of participants (9) were more satisfied with the structure under the treatment (equally: 6, less: 5).
-Participants did not show increased satisfaction with the blog post’s tone under the treatment compared to the baseline (Figure 5c).
-Looking at participants’ perceived task performance (Figure 6c), we observe that participants had higher perceived performance with Papers-to-Posts than with the baseline.
-The plurality of participants (9) thought they were more successful in writing a research-paper blog post with Papers-to-Posts compared to the baseline (equally: 7, less: 4).
-This further corroborates participants’ blog post satisfaction under the treatment condition.
-5.5.2. Tool Satisfaction
-For measuring tool satisfaction, we used the 5-question Likert scale on tool acceptance from Kocielnik et al.
-(Kocielnik et al., 2019) (Figure 5).
-The 5 responses were averaged to obtain one score.
-We used the non-parametric sign test to evaluate the data, as it was Likert data and violated the Wilcoxon signed rank test’s assumption of a symmetric distribution of the paired differences about the median.
-While participants had a higher median satisfaction with Papers-to-Posts (M=5.9,Q1=5.2,Q3=6.5) than the baseline tool (M=5.6,Q1=4.8,Q3=6.0), this difference was not significant (Dependent-Samples Sign Test, S=13, p=n.s.).
-In Figure 5e-h, we see that, under the treatment condition, participants were more likely to find the tool helpful, use it if it were available, recommend it, and perceive it as useful for productivity.
-For each of these questions, at least 4 more participants increased rather than decreased their Likert rating from the baseline to the treatment condition.
-However, for the question about the tool inducing annoyance, only 2 more participants indicated decreased rather than increased annoyance with Papers-to-Posts.
-5.5.3. Cognitive Load
-To measure cognitive load, we used the NASA TLX Index questionnaire (Hart and Staveland, 1988) (Figure 6a-d) and summed the responses for the mental demand, temporal demand, effort, and frustration questions (see 5.5.1 for discussion of the responses for performance).
-We utilized the non-parametric sign test to evaluate the data, as it was Likert data and violated the Wilcoxon signed rank test’s assumption of a symmetric distribution of the paired differences about the median.
-We saw no significant difference between the baseline and treatment conditions, although participants reported slightly lower cognitive load with Papers-to-Posts (M=13.0,Q1=11.5,Q3=16.3) than the baseline tool (M=14.5,Q1=12.0, Q3=18.3) (Dependent-Samples Sign Test, S=8, p=n.s.).
-Figure 6a-d presents little difference in mental demand, temporal demand, and frustration between the treatment and baseline conditions.
-However, with respect to effort, we note a divergence.
-The majority of participants (11) reported using less effort with Papers-to-Posts than with the baseline tool, whereas only 4 participants reported using more effort.
-5.5.4. Task Completion Time
-We recorded two metrics for task completion time.
-One was the amount of time participants needed during the session before they would be comfortable publicly publishing the blog post.
-The other was the perceived amount of time participants would need after the session before they would be comfortable publicly publishing the blog post.
-As participants reached that point of comfort during most sessions, we focused on the former metric.
-Because this data was numeric within-subjects data for two conditions of one factor, we used a paired-samples t-test.
-Participants worked slightly faster with the baseline tool (M=44.3 minutes,SD=12.5) in comparison to Papers-to-Posts (M=46.0 minutes,SD=13.9), but this difference was not significant (Paired-Samples t-Test, t(19)=0.55, p=n.s.).
-5.5.5. Control and Ownership
-Participants indicated higher perceived control (treatment: M=5.00,Q1=4.50,Q3=6.00, baseline: M=5.00, Q1=3.00,Q3=6.00) and satisfaction with their control (treatment: M=5.00,Q1=2.75,Q3=6.00, baseline: M=5.00,Q1=3.00,Q3=5.25) under the treatment condition only if we look within-subjects (Figure 7a-b).
-Eleven participants had higher perceived control with Papers-to-Posts than the baseline tool (equal: 2, lower: 7), and 10 participants had higher satisfaction with Papers-to-Posts (equal: 4, lower: 6).
-On the other hand, participants did not demonstrate increased perceived ownership (treatment: M=6.00,Q1=5.75,Q3=6.00, baseline: M=6.00,Q1=5.00,Q3=6.00) or satisfaction with ownership (treatment: M=6,Q1=4.00,Q3=6.25, baseline: M=6.00,Q1=5.00,Q3=6.00) under the treatment condition (Figure 7c-d).
-Although they felt that they had more control over Papers-to-Posts, participants may not necessarily have felt that the LLM tool contributed more or less to the creation of the blog post.
-5.5.6. Fulfillment of Design Goal 1
-Looking at Figure 8a-b, we see that participants were more confident that all essential information was in the blog post (treatment: M=6.00,Q1=3.00,Q3=6.25, baseline: M=5.00,Q1=3.00,Q3=6.00) and found it easier to include essential information from the paper (treatment: M=6.00,Q1=5.00,Q3=6.00, baseline: M=5.00,Q1=4.75,Q3=6.00) using Papers-to-Posts.
-In addition, the interaction logs show that participants utilized Papers-to-Posts’s DG1-inspired affordance to select relevant paper bullet points and paragraphs to incorporate in the blog post.
-All participants adjusted the LLM’s pre-selected bullet points to some degree.
-On average, participants added or removed 28 bullet points and 13 paragraphs.
-We note that some added bullet points come from automatic selection of bullet points used to generate a new section rather than manual selection.
-Participants also provided an average of 1.70 custom bullet points.
-In their short answers regarding aspects of Papers-to-Posts that they found difficult and helpful, 10 participants noted the paper outline being useful in some manner.
-Four participants described the outline as helpful for generating text.
-For instance, P15 noted, ”I could rewrite entire sections or direct it with just a few bullets.” Three participants mentioned that the outline was useful for reviewing the information in or structure of the paper.
-P4 commented, ”The outline was really helpful in providing a quick overview of what content I had written.” For two participants, the bullet points provided transparency around the origins of the initial LLM draft and generations.
-P6 explained, ”The checked-marked bullet points helped me understand where the pre-filled content on the left was from,” and P12 shared, ”I liked the bullet point features to see what was being fed in vs not.” Five participants also specified a difficulty related to the paper outline.
-Two participants were overwhelmed with the number of bullet points.
-P9 described the outline as ”really bloated and hard to navigate.” Three more participants were confused or encumbered by aspects of utilizing the bullet points.
-For example, P3 ”felt stressed by choosing between the entire paragraph at hand and the (presumably) AI-generated summary bullets,” and P20 did not realize that bullet points could come from different sections of the paper.
-5.5.7. Fulfillment of Design Goal 2
-Participants’ Likert responses regarding DG2 indicate that they found it easier to produce appropriate blog post text (treatment: M=6.00,Q1=5.00,Q3=6.00, baseline: M=5.00,Q1=4.00,Q3=6.00) and iterate on the blog post text (treatment: M=6.00,Q1=5.00,Q3=6.00, baseline: M=5.00,Q1=4.00,Q3=6.00) in the treatment condition (Figure 8c-d).
-Furthermore, the interaction logs for the treatment condition indicate that participants took advantage of DG2-based affordances for generating and modifying text for sub-tasks (Appendix 27).
-For an average of 6.85 generations per treatment condition, participants often utilized custom instructions (5.70 times), selected paragraphs or bullet points (6.35 times), and the context of prior blog post sections (6.00 times), though we note that these features were toggled on by default.
-In addition, they modified text 6.85 times on average, with the ”condense” (1.85 times), ”custom” (1.80 times), and ”simpler terms” (1.40 times) modifications being the most common.
-In their short answers, 5 participants noted the usefulness of being able to provide custom instructions for either generating or modifying text using Papers-to-Posts.
-P9 described the custom instructions as providing a greater sense of control: ”…being able to include explicit instructions for the model to generate text from was helpful in being able to control the information in the text that it generated.” With respect to revising writing in particular, 12 participants noted in their short answers that Papers-to-Posts’s affordances for modifying text were useful.
-All of the preset modification buttons (”simpler terms”, ”condense”, ”expand”, ”more dramatic”, ”less dramatic”) were specifically mentioned as helpful affordances at least once.
-The most commonly noted one was the ”condense” button.
-For example, P12 commented, ”I sometimes put the modify text through multiple rounds of expand and contract, focusing on different parts.
-I found that helpful.” Two participants commented that they desired more preset modification buttons.
-P16 wanted a ”more technical” button, while P7 simply shared, ”I wish there are more default buttons for similar functions.” P7 then added, ”Or even let user to add their favorite rewrite button.” Interestingly, P2 found the preset modification buttons useful for understanding the LLM tool’s capabilities.
-They reflected, ”It was nice to know what options the tool already knew of, and would perhaps have success with.
-For example, I would probably not have gone to a blank-box tool and said, ’please make the following more dramatic’.
-I would have stuck to more basic things like ’please summarize the following’.”
-Participants also discussed difficulties that they encountered in drafting and revising writing with Papers-to-Posts.
-Eight participants commented that they appreciated the simplicity of the baseline tool or that the baseline tool gave them more freedom over the instructions sent to the tool.
-For instance, P9 noted, ”the ’generate writing’ panel felt overwhelming in terms of features,” and P19 explained, ”It’s very open end so I felt like I got more freedom what to ask it to do.” Meanwhile, in the treatment condition, 4 participants expressed wanting to take specific actions that were not explicitly supported by Papers-to-Posts.
-As an example, 2 participants wanted assistance with checking for redundancies in the blog post.
-Another 2 wanted feedback on the blog post writing.
-Regarding the affordances for modifying text in particular, 8 participants noted a relative difficulty.
-A common issue, raised by 4 participants, was confusion about how or when to use certain affordances.
-As an example, P3 ”didn’t really understand if the custom instructions only applied for the ’custom modification’ option,” when it in fact would be applied to any modification action.
-Meanwhile, P6 was unsure of ”when to use which one of the six [modification] options.” Another recurrent issue noted by 5 participants was a modification affordance not working as desired.
-For P20, setting the desired length of the modified writing to one paragraph still resulted in three, as the back-end prompts are not guaranteed to work as planned.
-For P12, modifications were sometimes too exaggerated, while for P16, a custom modification to make writing more technical was too subdued.
-Despite these comments about feeling restricted or confused by Papers-to-Posts, participants still reported higher perceived control and satisfaction with control in the treatment condition, which we discuss in Section 5.5.5.
-5.5.8. Extent of Change to Blog Post
-Interestingly, we found that participants made larger changes to the initial LLM draft blog post when using Papers-to-Posts as compared to the baseline tool.
-In Figure 9a, for each condition, we show the Levenshtein distance, a common metric for measuring the amount of editing done to text, between participants’ initial LLM draft blog posts and their blog posts at the point at which the writer would feel comfortable publishing it publicly.
-For exploratory statistical testing, we employed a paired-samples t-test and observed that this Levenshtein distance is significantly higher with Papers-to-Posts (M=3977.05,SD=1146.46) compared to the baseline (M=2764.15,SD=1042.16) (Paired-Samples t-Test, t(19)=4.31, p¡.0005).
-In Figure 9a, we show how the Levenshtein distance progressed over time, with the removal of one outlier.
-The difference in Levenshtein distance does not appear attributable to the time spent on the task or the change in length of the blog post.
-Participants did not spend substantially more time to reach the point at which they were comfortable publicly publishing the blog post using Papers-to-Posts (M=49.37, Q1=35.43, Q3=59.54) as compared to the baseline (M=46.81, Q1=33.92,Q3=56.60) (Appendix 25).
-Furthermore, we saw that participants changed the length of their blog posts, measured in characters, less using Papers-to-Posts (M=-162.00, Q1=-1126.25, Q3=297.00) than using the baseline (M=1971.00, Q1=965.25,Q3=2514.25) (Appendix 25(b)).
-Thus, we suspected that Papers-to-Posts may have required a decreased number of writing actions (inserting, deleting, replacing text in text fields) in order to achieve a given increase in Levenshtein distance.
-We investigated this potential reduction in effort in the follow-up deployment study.
-6. Deployment Study
-Through the controlled lab study, we observed that participants produced final blog posts that were significantly more distant from their initial LLM-generated drafts when using Papers-to-Posts than the baseline tool.
-We wanted to further understand how participants in a more realistic setting perceive Papers-to-Posts versus the baseline tool (RQ2), and whether Papers-to-Posts would lead participants to change their provided drafts more in a given time period (RQ1).
-For this, we conducted a deployment study where participants wrote blog posts on their own time with the explicit goal of posting the article publicly and sharing it with others.
-6.1. Participants
+At the session’s end, the coordinator told the participant to close the tool and not to access it before the next session if they had one.
+Each time the participant prompted the LLM or submitted the blog post, a log of the action along with the state of the blog post was saved for analysis.
+4.2. Deployment Study
+We conducted a deployment study to understand if and how the results of the lab study translate to a less controlled and more realistic setting.
+To make the deployment study more realistic than the lab study, participants wrote blog posts on their own time with the explicit goal of sharing the blog post with others (e.g., on social media, in an email to colleagues).
+4.2.1. Research Questions
+Having observed that the lab study participants had significantly more editing power when using Papers-to-Posts rather than the baseline tool, we investigated if Papers-to-Posts leads participants to have greater editing power in this real-world context (RQ1).
+As in the lab study, we also evaluated satisfaction with the final blog post, satisfaction with the tool, and cognitive load (RQ2).
+We did not evaluate task completion time in the deployment study, as the lab study did not demonstrate any interesting difference between the conditions.
+RQ1: How is potential users’ editing power impacted by Papers-to-Posts versus the baseline tool in a real-world setting?
+RQ2: How do potential users perceive the outcome (blog post satisfaction (a)) and experience (tool satisfaction (b), cognitive load (c)) of interacting with Papers-to-Posts versus the baseline tool in a real-world setting?
+For the deployment study, we posed research questions rather than evaluated hypotheses for statistical significance because power analyses indicated that we would need a larger sample size than we could obtain to see the same significant results from the lab study.111111The significant editing power result would have required 29 participants per condition (power=0.95, alpha=0.05, effect size=0.45), and the significant blog post satisfaction result would have required 83 participants per condition (power=0.95, alpha=0.05, effect size=0.98).
+4.2.2. Participants
 We recruited 26 participants (M: 17, W: 8, undisclosed: 1) through academic social networks and institutional mailing lists.
@@
 Lastly, the papers that the participants want to write about had to be compatible with our tools’ PDF parser.
-6.2. Study Conditions
+4.2.3. Study Conditions
 While the treatment and baseline conditions were largely the same as in the lab study, several minor usability issues identified in the lab study were addressed before proceeding with the deployment study.
 For example, there were better instructions and less intrusive highlight colors (Appendix 28).
-We also replaced GPT-4 and GPT-4-32k with the newer GPT-4-0125-preview model with a larger context window that became available after the lab study.
+We also replaced GPT-4 and GPT-4-32k with the newer GPT-4-0125-preview model, which had a larger context window and became available after the lab study.
 This enabled Papers-to-Posts to support longer inputs.
-In the deployment study, participants could freely close and reopen the tool in their browsers to work on their drafts when convenient.
+In the deployment study, participants could freely close and reopen their assigned tools in their browsers to work on their drafts when convenient.
 Participants could save their progress across sessions with a newly added save button.
@@
 The average parsed paper length across conditions differed by less than 5% (baseline: 37064 characters, treatment: 35845 characters).
-6.3. Study Procedure
-The task for the deployment study was to share a blog post with others, which can require significantly higher effort than the controlled lab study, which only required voicing comfort in publicly posting the blog post.
-To control for the study length and ease of recruitment, we opted for a between-subjects study so that each participant only had to write one blog post.
-At the same time, we allowed participants who wished to participate in the study twice (N=11/26 participants) to use both tools each time in randomized, counter-balanced order.
-When analyzing the results we treated them as non-paired data.
-Eighteen participants completed the baseline condition, and 19 participants completed the treatment condition.
-For participants who wrote two blog posts, their selected papers were assigned randomly to each condition.
-Participants had up to six days to complete a research-paper blog post using the assigned tool on their own time.
-We provided participants with a video tutorial on the assigned tool, a link to the paper about which they would be writing, and a link to the assigned tool.
-We noted that their aim should be to share the final blog post with others, but if they felt that they could not reach the point of being comfortable with sharing it with others, they could let us know and still be compensated.
-Participants were required to spend at least 20 minutes using the tool to write the blog post.
-To better understand how participants utilized the two systems, we recorded their interactions, including prompting, inserting, deleting and replacing text.
-In addition, we also took periodic snapshots of the drafts whenever they (re)opened the systems in their browsers and every 60 seconds during active usage.
-Once participants were satisfied with their draft, participants shared a blog post based on their drafts.
-We provided instructions on how to create blog posts on the Medium platform,222https://medium.com/ but they were allowed to use other publishing platforms if they preferred.
+4.2.4. Study Procedure
+The task for the deployment study was to share a blog post with others, which can require substantially higher effort than voicing comfort in publicly posting the blog post, as in the lab study.
+Considering this difference for participant recruitment, we opted for a between-subjects design, so each participant only had to write one blog post.
+Still, we allowed participants who wished to participate twice (N=11/26 participants) to use both tools in randomized, counter-balanced order.
+Their selected papers were assigned randomly to each condition.
+When analyzing these participants’ results, we treated them as unpaired data.
+Eighteen participants completed the baseline, and 19 completed the treatment.
+Participants had up to six days to complete a paper blog post using the assigned tool on their own time.
+We provided a link to the tool, a video tutorial on the tool, and a link to the paper about which to write.
+We noted that participants should aim to share the final blog post with others, but if they felt that they could not reach that point, they would still be compensated.
+Participants were required to spend at least 20 minutes using the tool.
+We recorded participants’ interactions, including prompting and text edits.
+In addition, we also took snapshots of the blog post whenever they (re)opened the systems in their browsers and every 60 seconds during active usage.
+Once participants were satisfied with their draft, they shared it.
+We provided instructions on how to create blog posts on the Medium platform,121212https://medium.com/ but they were allowed to use other publishing platforms if they preferred.
 They then emailed us the blog post link and a screenshot or link showcasing evidence that they shared the blog post with others (e.g., tweet, email to co-authors, Slack channel message, LinkedIn post).
-Only 2 of the 26 participants (one under both conditions and one under the treatment condition) alerted us to the fact that they could not reach the point at which they felt comfortable sharing the blog post with others in the allotted time.333We note that some participants indicated that the blog post was part of an experiment when sharing with others, which may have increased their willingness to share it; regardless, they were willing to share a post at the end of the study.
-Participants then completed the same survey from the lab study regarding their experience with the tool.
-We note that four participants who we realized had not completed 20 minutes of interaction were asked to do so after completing the survey.
-The only difference was the removal of a few questions irrelevant to this study: a question regarding how much additional time would be needed to feel comfortable posting the blog post publicly and questions regarding a direct comparison between the two tools.
-Lastly, for the 15 participants who were willing, the first author engaged them in a 15-minute semi-structured interview to discuss their experience with the tool.
-6.4. Results
-We did not plan to investigate statistical significance in this between-subjects study, given power analyses indicating that we would need a larger sample size than we could obtain to see a significant result for our metric of Levenshtein distance (power=0.95, alpha=0.05, effect size=0.45; requires 29 participants per condition) or blog post satisfaction (power=0.95, alpha=0.05, effect size=0.98; requires 83 participants per condition).
-Throughout our analysis of the deployment study logs, we exclude the three study runs in which the participant made no change to the initial draft, as there is no interaction in producing the blog post.
-We also exclude six study runs in which the participant claimed to have made edits outside the tool beyond small things like grammar, formatting, and adding images; this is to avoid drawing conclusions based on missing data about effort spent.
-The resulting log data used thus includes 28 study runs (treatment: 14, baseline: 14).
-The first author conducted inductive thematic analysis (Braun and Clarke, 2006) of the participants’ survey and interview comments.
-6.4.1. RQ1: Blog Post Iteration
-When given a time constraint in the lab study, we observed that participants made larger changes to the tool’s initial draft with Papers-to-Posts than with the baseline tool.
-The deployment study allowed us to investigate if this evident decrease in effort could be reflected in a realistic setting and in terms of other metrics of effort (i.e., number of writing actions).
-We found that participants using Papers-to-Posts took fewer writing actions in order to create one unit of change in Levenshtein distance to the blog post (Figure 9(c)).
-We also observed, in alignment with the lab study, that Papers-to-Posts led to greater changes in Levenshtein distance for a given amount of active time, which we define as time taking any action in the tool (Figure 9(b)).
-Therefore, when using Papers-to-Posts as compared to the baseline, participants experienced reduced effort in iterating on the blog post with respect to both writing actions needed and time needed.
-Once more, this is despite the fact that participants did not spend substantially more active time using Papers-to-Posts (M=22.00, Q1=17.25,Q3=34.00) than the baseline (M=22.50,Q1=16.00,Q3=27.00) (Appendix 26(a)).
-Furthermore, we again saw that participants changed the length of their blog posts, measured in characters, less using Papers-to-Posts (M=-5.5, Q1=-525.25, Q3=961.00) than using the baseline (M=710.00, Q1=47.25, Q3=4261.25) (Appendix 26(b)).
-In fact, Figure 9 shows that the maximum active time and writing actions taken by a baseline participant is higher than that of a treatment participant.
-This provides some evidence that, for writers who tend to spend more time than average on the task, Papers-to-Posts may help them to reduce the time they spend on it.
-Participants’ survey and interview comments reflected this discrepancy between the two tools.
-With respect to the baseline, participants most commonly cited the initial draft as a helpful aspect of the tool.
-Nine participants under the baseline condition specifically noted that it was useful as a starting point.
-Focusing on the initial draft, baseline participants often spoke of editing or adding to it rather than replacing it.
-For example, P20 explained, ”It’s easier to critique than it is to generate,” and P26 shared, ”It was very useful to have a straw-person of what the blog post could look like (generated by the LLM).
-I mostly used the draft as a place to edit and steer in the direction I wanted the post to take.”
-On the other hand, 15 participants noted the selectable paper bullet points and paragraphs as a helpful feature of Papers-to-Posts.
-Focusing on the feature designed to generate full blog post sections, participants using Papers-to-Posts often reported doing just that.
-As an example, P12 commented that the ”outline [to] text generation process is very helpful to get the first draft of the post.” Furthermore, because the tool shows users what bullet points the tool used to generate the initial draft, participants may have been inclined to inspect and iterate upon this selection.
-This is illustrated by P20, who observed, ”The automatic selection of bullet points to include did not capture most of the information I wanted it to, and so required a lot of manual selection of bullet points, but once I selected those, the generation made it easy to write.” Participants also frequently mentioned the utility of the Papers-to-Posts’s modification abilities.
-P18 reflected, ”The ”MODIFY WRITING” section is helpful.
-I like that the interface contains this particular section to support iterating on existing text.” By facilitating modifications to the initial draft, Papers-to-Posts may have encouraged participants to iterate more heavily on the initial draft.
-6.4.2. RQ2: Real-World Tool Perceptions
-We evaluated participants’ survey results in the same manner as in the lab study (see Section 5.5).
-With respect to satisfaction with the final blog post quality, we observed no difference in the median satisfaction under Papers-to-Posts(M=6.00,Q1=5.00,Q3=6.00) versus the baseline (M=6.00,Q1=5.25, Q3=6.00) (Figure 5j-m).
-This discrepancy with the lab study may be related to the fact that participants were not writing under a time constraint in the deployment study.
-With sufficient time to write the blog post to their liking (only 3 blog posts were not shared with others), participants may not have noticed a pronounced difference in blog post quality between conditions.
+Only 2 of the 26 participants (one under both conditions and one under the treatment) could not reach the point at which they felt comfortable sharing the blog post with others in the allotted time.131313We note that some participants indicated that the blog post was part of an experiment when sharing with others, which may have increased their willingness to share it; regardless, they were willing to share a post at the end of the study.
+Participants then completed the same survey from the lab study regarding their experience (see Section 4.1.4).141414We note that four participants who we realized had not completed 20 minutes of interaction were asked to do so after completing the survey.
+The only difference was the removal of a few questions irrelevant to this study151515a question regarding how much additional time would be needed to feel comfortable posting the blog post publicly and questions regarding a direct comparison between the two tools.
+Lastly, the first author engaged willing participants (15) in a 15-minute semi-structured interview to discuss their experience and survey responses.
+5. Evaluation Studies’ Results
+In this section, we present the quantitative and qualitative results of the lab and deployment studies in parallel.
+To obtain the qualitative results, the first author conducted inductive thematic analysis (Braun and Clarke, 2006) of the participants’ interview and short-answer responses across the two studies.
+Please note that the hypotheses from the within-subjects lab study were not re-evaluated for significance in the between-subjects deployment study because there were not enough participants to do so (see Section 4.2.1).
+5.1. Editing Power
+5.1.1. Lab Study - Exploratory
+Participants made greater change to the initial LLM draft blog post when using Papers-to-Posts as compared to the baseline tool.
+For each condition, we calculated the Levenshtein distance between participants’ initial LLM draft blog posts and their blog posts at the point at which they would feel comfortable publicly publishing the blog post.
+Levenshtein distance measures how many character insertions, deletions, and replacements would be needed to transform one text into another.
+For statistical testing, we employed a paired-samples t-test and observed that this Levenshtein distance is significantly higher with Papers-to-Posts (M=3977.05, SD=1146.46) compared to the baseline (M=2764.15, SD=1042.16) (Paired-Samples t-Test, t(19)=4.31, p¡.0005).161616We note that this significance test was not part of our initial hypotheses and was therefore exploratory in nature.
+The difference in Levenshtein distance does not appear attributable to the time spent on the task or the change in blog post length.
+Participants did not spend substantially more time to reach the point at which they were comfortable publicly publishing the blog post using Papers-to-Posts as compared to the baseline (see Section 5.5).
+Figure 5a shows Levenshtein distance plotted against the time taken to get to a publishable blog post, with the removal of one outlier.
+Furthermore, participants changed the length of their blog posts, measured in characters, less with Papers-to-Posts (M=-162.00, Q1=-1126.25, Q3=297.00) than with the baseline (M=1971.00, Q1=965.25, Q3=2514.25) (Appendix 26(b)).171717As you can see in Appendix 26(b), the initial length of treatment drafts was significantly higher than the initial length of baseline drafts.
+5.1.2. Deployment Study - RQ1
+We investigated if the increase in editing power observed in the lab study would be reflected in a more realistic setting and if Papers-to-Posts required fewer writing actions (i.e., inserting/deleting a character or text span in the draft or LLM instructions)181818We did not count when participants moved, inserted, or deleted a blog post section in Papers-to-Posts as a writing action because these actions involved clicking a button rather than manipulating text.
+in order to achieve a given amount of change in the blog post.
+To analyze editing power, we excluded the three study runs in which the participant made no change to the initial draft.
+We also excluded six study runs in which the participant claimed to have made edits to the blog post outside the tool, beyond small edits like those related to grammar, formatting, and adding images.
+The resulting analysis thus included 28 study runs (treatment: 14; baseline: 14).
+Participants using Papers-to-Posts took fewer writing actions to create one unit of change in Levenshtein distance to the blog post (Figure 5(c)).
+Papers-to-Posts also led to a greater final Levenshtein distance for a given amount of active time, which we define as time during which participants were taking any actions in the tool (Figure 5(b)).
+Once more, this is despite the fact that participants did not spend substantially more active time using Papers-to-Posts (M=25.71 minutes, SD=14.72) than the baseline (M=35.93 minutes, SD=34.38) (Appendix 27(a)).191919Considering that there were three outliers with a high amount of active time in the baseline condition, we can also look at the median behavior.
+We see little difference between the conditions (treatment: M=22.00, Q1=17.25, Q3=34.00; baseline: M=22.50, Q1=16.00,27.00).
+Furthermore, participants changed the length of their blog posts, measured in characters, less using Papers-to-Posts less using Papers-to-Posts (M=-5.50, Q1=-525.25, Q3=961.00) than using the baseline (M=710.00, Q1=47.25, Q3=4261.25) (Appendix 27(b)).202020As you can see in Appendix 27(b), the initial length of treatment drafts was significantly higher than the initial length of baseline drafts.
+5.1.3. Summary
+In the lab study, our exploratory analysis revealed that Papers-to-Posts led to increased editing power, in the sense that participants made more change to their blog posts in a given amount of time.
+In the deployment study, participants experienced the same outcome in a more realistic setting.
+We further found that participants made more change to their blog posts for a given number of writing actions.
+5.2. Blog Post Satisfaction
+To measure participants’ satisfaction with their final blog posts, we utilized their Likert-type response to the statement “I am satisfied with the overall quality of the blog post” (Figure 6a,j).
+5.2.1. Lab Study - H1
+To compare responses under the two conditions in the within-subjects lab study, we employed a non-parametric test, given the data’s Likert nature.
+Furthermore, we used a sign test rather than a Wilcoxon signed rank test, which is normally used to analyze non-parametric within-subjects data for two conditions of one factor, due to the violation of that test’s assumption of a symmetric distribution of the paired differences about the median.
+Participants were significantly more satisfied with the final blog post generated with Papers-to-Posts (M=6.00, Q1=5.00, Q3=7.00) as compared to the baseline tool (M=5.00, Q1=5.00, Q3=6.00) (Dependent-Samples Sign Test, S=11, p¡.05).
+The majority of participants (11) were more satisfied in the treatment condition (equally: 7, less: 2).
+Inspecting participants’ Likert responses regarding aspects of their blog post satisfaction, we see that participants particularly demonstrated more satisfaction with the content covered in the Papers-to-Posts blog post (M=6.00, Q1=5.00, Q3=7.00) compared to the baseline blog post (M=5.50, Q1=4.75, Q3=6.00) (Figure 6b).
+The majority of participants (14) were more satisfied with the content under the treatment condition (equally: 2, less: 4).
+Although it is not apparent from the medians, participants were overall more satisfied with the blog post structure using Papers-to-Posts (M=6.00, Q1=5.00, Q3=6.25) than the baseline (M=6.00, Q1=5.00, Q3=6.00) (Figure 6d).
+The plurality of participants (9) were more satisfied with the structure under the treatment (equally: 6, less: 5).
+Participants did not show increased satisfaction with the blog post’s tone under the treatment (M=6.00, Q1=5.00, Q3=6.00) compared to the baseline (M=5.50, Q1=5.00, Q3=6.00) (Figure 6c).
+A plurality of participants were not more satisfied with the tone under the treatment (more: 7, equally: 6, less: 7).
+Looking at participants’ perceived failure to complete their task (Figure 7c), we observe that participants had lower perceived failure with Papers-to-Posts (M=2.00, Q1=2.00, Q3=3.00) than with the baseline (M=3.00, Q1=2.00, Q3=4.00).
+The plurality of participants (9) thought that they were more successful in writing a research-paper blog post with Papers-to-Posts compared to the baseline (equally: 7, less: 4).
+This further corroborates participants’ blog post satisfaction under the treatment condition.
+5.2.2. Deployment Study - RQ2a
+In the between-subjects deployment study, we observed no difference in the median satisfaction under Papers-to-Posts (M=6.00, Q1=5.00, Q3=6.00) versus the baseline (M=6.00, Q1=5.25, Q3=6.00) in blog post satisfaction (Figure 6j).
+Again with respect to median satisfaction, participants were equally satisfied with the blog post content (treatment: M=6.00, Q1=5.00, Q3=6.50; baseline: M=6.00, Q1=3.50, Q3=6.00; Figure 6k), tone (treatment: M=6.00, Q1=5.00, Q3=6.00; baseline: M=6.00, Q1=5.00, Q3=7.00; Figure 6l), and structure (treatment: M=6.00, Q1=6.00, Q3=6.00; baseline: M=6.00, Q1=5.25, Q3=7.00; Figure 6m) when using Papers-to-Posts versus the baseline.
+In terms of perceived failure to complete their task (Figure 7h), participants had equal median perceived failure with Papers-to-Posts (M=3.00, Q1=2.00, Q3=3.00) and the baseline (M=3.00, Q1=2.00, Q3=3.00).
+5.2.3. Summary
+We observed a significant increase in blog post satisfaction under the treatment condition during the lab study and found that participants were particularly more satisfied with the blog post content, suggesting that Papers-to-Posts best addressed its design goal to support planning blog post content (DG1).
+We did not observe the same increases in blog post satisfaction in the deployment study.
+This discrepancy may be related to the fact that participants were not writing under a time constraint in the deployment study.
+With sufficient time to write the blog post to their liking (only three blog posts were not shared with others), participants may not have noticed a pronounced difference in blog post quality between conditions.
+5.3. Tool Satisfaction
+For measuring tool satisfaction, we used the five-question Likert scale on tool acceptance from Kocielnik et al.
+(Kocielnik et al., 2019) (Figure 6e-i).
+The five responses were averaged to obtain one score.
+5.3.1. Lab Study - H2
+We used the non-parametric sign test to evaluate the lab study data, as it was Likert data and violated the Wilcoxon signed rank test’s assumption of a symmetric distribution of the paired differences about the median.
+While participants had a higher median satisfaction with Papers-to-Posts (M=5.90, Q1=5.20, Q3=6.50) than the baseline tool (M=5.60, Q1=4.80, Q3=6.00), this difference was not significant (Dependent-Samples Sign Test, S=13, p=n.s.).
+Based on median results, participants under the treatment condition more often found the tool helpful (treatment: M=6.00, Q1=5.75, Q3=7.00; baseline: M=5.50, Q1=4.75, Q3=6.00) and useful for productivity (treatment: M=6.00, Q1=5.00, Q3=7.00; baseline: M=5.50, Q1=5.00, Q3=6.00).
+Meanwhile, participants were equally likely under both conditions to use the tool if it were available (treatment: M=6.00, Q1=5.00, Q3=7.00; baseline: M=6.00, Q1=5.00, Q3=6.00), recommend the tool (treatment: M=5.50, Q1=4.75, Q3=7.00; baseline: M=5.50, Q1=4.00, Q3=6.00), and find the tool annoying or distracting (treatment: M=2.00, Q1=2.00, Q3=3.00; baseline: M=2.00, Q1=2.00, Q3=3.00).
+For each of the first four questions, at least four more participants increased rather than decreased their Likert rating from the baseline to the treatment condition.
+However, only two more participants indicated decreased rather than increased annoyance with Papers-to-Posts (less: 8, more: 6, equal: 6).
+5.3.2. Deployment Study - RQ2b
 Similarly to the lab study, we observed slightly greater satisfaction with the tool under the treatment condition (M=5.60, Q1=5.00, Q3=6.40) than the baseline condition (M=5.40, Q1=4.05, Q3=5.80) (Figure 6n-r).
-In alignment, four participants who participated twice expressed a preference for Papers-to-Posts, whereas no one shared a preference for the baseline tool.
-P19 remarked, ”I mean, honestly, I like the first tool better than the second tool… the second tool is a bit simplistic.” However, the complexity and larger associated learning curve of Papers-to-Posts may have decreased its chances of being widely preferred over the baseline.
-For example, P15 commented, ”It was a lot of information on the interface to deal with.
-I would prefer less moving things in front of me.” Nevertheless, we once again found slightly lower cognitive load under the treatment condition (M=15.00,Q1=9.50,Q3=17.50) in comparison to the baseline condition (M=15.50,Q1=10.75,Q3=18.75) (Figure 6e-h).
-This may be related to the fact that, while Papers-to-Posts required understanding a more complex interface, the baseline tool placed greater demands on participants in terms of writing.
-P24 summarized, ”In this [baseline] version, you have greater autonomy regarding what needs to be done.
-While this is beneficial, it also requires more effort and planning.”
-As shown in Figure 7e-h, participants indicated higher satisfaction with control (treatment: M=5.00,Q1=3.00,Q3=6.00, baseline: M=3.00,Q1=3.00,Q3=4.75) in the treatment condition like in the lab study, but they showed no difference in median satisfaction with respect to perceived control (treatment: M=5.00,Q1=5.00,Q3=6.00, baseline: M=5.00,Q1=3.00,Q3=5.00), perceived ownership (treatment: M=6.00,Q1=4.50,Q3=6.00, baseline: M=6.00,Q1=6.00,Q3=6.00) or satisfaction with ownership (treatment: M=6.00,Q1=4.00,Q3=6.00, baseline: M=6.00,Q1=5.00,Q3=6.00) in the treatment condition.
-Participants may have been more satisfied with their control of the tool without perceiving a higher control using Papers-to-Posts in the sense that they were comfortable with letting the tool have a certain amount of control in exchange for less effort to generate text.
-Overall, we observed benefits from Papers-to-Posts in fulfilling Papers-to-Posts’s design goals.
-Regarding the first design goal (DG1), participants indicated only a slight improvement in feeling confident that all essential information was covered with Papers-to-Posts (M=5.00,Q1=4.50,Q3=6.00) as opposed to the baseline (M=5.00,Q1=4.25,Q3=6.00).
-Given that the automatic method for parsing the paper PDF in the deployment study was likely to make more errors than the partially Wizard-of-Oz method for parsing papers in the lab study, it makes sense that participants would now see a reduced benefit from Papers-to-Posts’s paper outline in terms of ensuring all relevant information is included.
+Based on median results, participants under the treatment condition were more likely to find the tool helpful (treatment: M=6.00, Q1=5.00, Q3=6.50; baseline: M=5.00, Q1=5.00, Q3=6.00), use it if it were available (treatment: M=6.00, Q1=5.00, Q3=7.00; baseline: M=5.00, Q1=4.00, Q3=6.00), and recommend the tool (treatment: M=6.00, Q1=5.00, Q3=6.50; baseline: M=5.00, Q1=3.25, Q3=6.00).
+They were also less likely to find the tool annoying (treatment: M=2.00, Q1=2.00, Q3=4.00; baseline: M=2.50, Q1=2.00, Q3=4.00).
+On the other hand, they were equally likely under both conditions to find the tool useful for productivity (treatment: M=6.00, Q1=4.50, Q3=7.00; baseline: M=6.00, Q1=4.00, Q3=6.00).
+5.3.3. Summary
+Across both studies, we observed higher satisfaction with Papers-to-Posts than the baseline tool, though this difference was not significant according to the lab study results.
+In particular, participants found Papers-to-Posts more helpful.
+5.4. Cognitive Load
+To measure cognitive load, we used the NASA TLX Index questionnaire (Hart and Staveland, 1988) and summed the responses for the mental demand, temporal demand, effort, and frustration questions (Figure 7a-b,d-e)212121See 5.2 for discussion of the responses to the performance question.
+5.4.1. Lab Study - H3
+We utilized the non-parametric sign test to evaluate the lab study data, as it was Likert data and violated the Wilcoxon signed rank test’s assumption of a symmetric distribution of the paired differences about the median.
+We saw no significant difference between the baseline and treatment conditions (Dependent-Samples Sign Test, S=8, p=n.s.), although participants reported slightly lower cognitive load with Papers-to-Posts (M=13.00, Q1=11.50, Q3=16.30) than the baseline tool (M=14.50, Q1=12.00, Q3=18.30).
+We saw the largest difference in median results with respect to perceived effort (treatment: M=3.50, Q1=3.00, Q3=5.00; baseline: M=5.00, Q1=3.75, Q3=5.00).
+The majority of participants (11) reported using less effort with Papers-to-Posts than with the baseline tool (more: 4, equal: 5).
+Meanwhile, Papers-to-Posts led to an increase in perceived mental demand (treatment: M=5.00, Q1=3.00, Q3=5.00; baseline: M=4.00, Q1=3.75, Q3=5.25).
+The small plurality of participants (9) perceived more mental demand with Papers-to-Posts (less: 8, equal: 3).
+We saw no change in the median result for perceived temporal demand (treatment: M=3.00, Q1=2.00, Q3=4.25; baseline: M=3.00, Q1=2.00, Q3=4.25) or frustration (treatment: M=2.00, Q1=1.00, Q3=4.00; baseline: M=2.00, Q1=2.00, Q3=3.00).
+An equal number of participants found the treatment more (9) versus less (9) temporally demanding, and the small plurality of participants (9) found the treatment less frustrating (more: 7, equal: 4).
+5.4.2. Deployment Study - RQ2c
+Once again, we found slightly lower cognitive load under the treatment condition (M=15.00, Q1=9.50, Q3=17.50) in comparison to the baseline condition (M=15.50, Q1=10.75, Q3=18.75) (Figure 7f-g,i-j).
+Looking at the breakdown of questions, the treatment led to less perceived mental demand (M=4.00, Q1=2.00, Q3=5.50) than the baseline (M=5.00, Q1=3.25, Q3=5.75) but higher perceived effort (treatment: M=5.00, Q1=3.00, Q3=6.00; baseline: M=4.50, Q1=3.00, Q3=5.00).
+We found no change in the median result for perceived temporal demand (treatment: M=3.00, Q1=2.00, Q3=4.00; baseline: M=3.00, Q1=2.00, Q3=4.00) or frustration (treatment: M=3.00, Q1=2.00, Q3=5.00; baseline: M=3.00, Q1=2.00, Q3=4.75).
+5.4.3. Summary
+In both studies, we observed a small decrease in cognitive load from the baseline to the treatment condition, but this difference was not significant according to the lab study results.
+Participants consistently perceived little difference in the tools with respect to temporal demand and frustration.
+However, their perceptions varied across the studies on whether Papers-to-Posts was more or less mentally demanding and whether it required more or less effort.
+5.5. Task Completion Time
+5.5.1. Lab Study - H4
+We recorded two metrics for task completion time.
+One was the amount of time participants needed during the session before they would be comfortable publicly publishing the blog post.
+The other was the perceived amount of time participants would need after the session before they would be comfortable publicly publishing the blog post.
+As participants reached that point of comfort during most sessions, we focused on the former metric.
+Because this data was numeric within-subjects data for two conditions of one factor, we used a paired-samples t-test.
+Participants worked slightly faster with the baseline tool (M=44.32 minutes, SD=12.51) in comparison to Papers-to-Posts (M=45.99 minutes, SD=13.88) (Appendix 26(a)), but this difference was not significant (Paired-Samples t-Test, t(19)=0.55, p=n.s.).
+5.6. Fulfillment of Planning Design Goal (DG1)
+We analyzed participants’ Likert-type responses to questions related to fulfillment of our first design goal, related to planning the blog post (Figure 8a-b,e-f).
+5.6.1. Lab Study - Exploratory
+Based on the median results, participants were more confident that all essential information was in the blog post (treatment: M=6.00, Q1=3.00, Q3=6.25; baseline: M=5.00, Q1=3.00, Q3=6.00) and found it easier to incorporate information from the paper (treatment: M=6.00, Q1=5.00, Q3=6.00; baseline: M=5.00, Q1=4.75, Q3=6.00) using Papers-to-Posts (Figure 8a-b).
+In addition, the interaction logs showed that participants utilized Papers-to-Posts’s DG1-inspired affordance to select relevant paper bullet points and paragraphs to incorporate in the blog post.
+All participants adjusted the LLM’s pre-selected bullet points to some degree.
+On average, participants added or removed 28 bullet points and 13 paragraphs.
+We note that some added bullet points come from automatic selection of bullet points used to generate a new section rather than manual selection.
+Looking at participants’ logs, when selecting points to generate text, only a few exclusively selected original-text paragraphs rather than bullet points.
+Participants also provided custom bullet points an average of 1.70 times.
+5.6.2. Deployment Study - Exploratory
+Looking at the median results, participants did not indicate an improvement in feeling confident that all essential information was covered with Papers-to-Posts (M=5.00, Q1=4.50, Q3=6.00) as opposed to the baseline (M=5.00, Q1=4.25, Q3=6.00).
+Given that the automatic method for parsing the paper PDF in the deployment study was likely to make more errors than the partially manual method for parsing papers in the lab study, it makes sense that participants would now see a reduced benefit from Papers-to-Posts’s paper outline in terms of ensuring all relevant information is included.
 Multiple participants under both conditions noted that the parsing of their papers was incomplete or flawed.
 However, participants still indicated an increased ease in incorporating information from the paper in the blog post with Papers-to-Posts (M=6.00, Q1=5.00, Q3=6.00) compared to the baseline (M=4.50, Q1=3.25, Q3=5.75).
-For instance, P14 commented, ”I liked being able to check the boxes of the portions of the paper I wanted to include in each section.”
-With respect to our second design goal (DG2), there was no difference in the median for how well Papers-to-Posts (M=5.00,Q1=4.00, Q3=6.00) versus the baseline (M=5.00,Q1=5.00,Q3=6.00) facilitated production of appropriate text, unlike in the lab study (Figure 8e-h).
-Nevertheless, in their qualitative feedback, several participants noted how Papers-to-Posts’s generation and modification features helped them to create more appropriate text.
-For instance, P24 explained that the ”preset custom instructions was very useful because can need to be very specific using these tools,” and P14 noted that the ”provided buttons for making text more or less dramatic was nice.” Furthermore, in their survey responses, participants indicated that Papers-to-Posts (M=6.00,Q1=4.50,Q3=6.00) facilitated iterating on the blog post more so than the baseline (M=5.00,Q1=3.50,Q3=6.00), which is reflected in the lower number of writing actions participants took for a unit change in their initial drafts that we observed in Figure 9(c).
-7. Discussion and Design Implications
-Based on the results of the lab and deployment studies, we reflect on design implications for other human-LLM scientific long-document summarization tools with long-form outputs.
+For instance, P14 commented, “I liked being able to check the boxes of the portions of the paper I wanted to include in each section.”
+5.6.3. Summary
+In both studies, participants indicated increased ease in incorporating paper content in the blog post with Papers-to-Posts.
+However, the two studies had different outcomes with respect to participants’ confidence in all essential information being covered in the blog post.
+While lab study participants indicated higher confidence with Papers-to-Posts the deployment study participants did not.
+This may be because, in the deployment study, we did not include manual corrections to the paper parsing.
+5.7. Fulfillment of Drafting and Revising Design Goals (DG2, DG3)
+We analyzed participants’ Likert-type responses to questions related to fulfillment of our second and third design goals, related to drafting and revising the blog post (Figure 8c-d,g-h).
+5.7.1. Lab Study - Exploratory
+Participants’ Likert responses regarding DG2 and DG3 indicate that they found it easier to produce appropriate blog post text (treatment: M=6.00, Q1=5.00, Q3=6.00; baseline: M=5.00, Q1=4.00, Q3=6.00) and iterate on the blog post text (treatment: M=6.00, Q1=5.00, Q3=6.00; baseline: M=5.00, Q1=4.00, Q3=6.00) in the treatment condition (Figure 8c-d).
+Furthermore, the interaction logs for the treatment condition indicate that participants took advantage of DG2-based affordances for generating text and DG3-based affordances for modifying text (Appendix 25).
+For an average of 6.85 generations per treatment condition, participants often utilized custom instructions (5.70 times), selected paragraphs or bullet points (6.35 times), and included the context of prior blog post sections (6.00 times), though we note that these features were pre-filled or toggled on by default.
+In addition, they modified text 6.85 times on average, with the “condense” (1.85 times), “custom” (1.80 times), and “simpler terms” (1.40 times) modifications being the most common.
+5.7.2. Deployment Study - Exploratory
+There was no difference in the median for how well Papers-to-Posts (M=5.00, Q1=4.00, Q3=6.00) versus the baseline (M=5.00, Q1=5.00, Q3=6.00) facilitated production of appropriate text, unlike in the lab study (Figure 8e-h).
+Nevertheless, participants indicated that Papers-to-Posts (M=6.00, Q1=4.50, Q3=6.00) facilitated iterating on the blog post more so than the baseline (M=5.00, Q1=3.50, Q3=6.00).
+5.7.3. Summary
+In both studies, participants using Papers-to-Posts found it easier to iterate on the blog post.
+However, while lab study participants thought Papers-to-Posts improved their ability to produce appropriate text for the blog post, deployment study participants did not.
+Perhaps without time constraints, deployment study participants had more time to determine how to prompt the baseline tool or adjust its text in order to produce appropriate text.
+5.8. Qualitative Analysis
+Across the two studies, participants’ survey and interview comments reflected the difference in editing power between the two tools.
+5.8.1. Baseline
+Overall, participants using the baseline most commonly cited the initial draft (lab: 6/20, deployment: 8/18) as a helpful aspect of the tool.
+Thus, the most valued baseline feature was not helpful for editing power.
+With respect to the initial draft, multiple participants noted editing or adding to the initial draft rather than replacing it.
+For example, P20-dep explained, “It’s easier to critique than it is to generate,” and P26-dep shared, “It was very useful to have a straw-person of what the blog post could look like (generated by the LLM).
+I mostly used the draft as a place to edit and steer in the direction I wanted the post to take.” A number of participants also noted the tool’s flexibility as a benefit (lab: 6/20, deployment: 4/18).
+For example, P18-lab shared, “I was able to provide instructions that fit my mental model of constructing a blog post and it being there as a ‘copilot’ is just what I need.” Flexibility in drafting text may help writers obtain the edits they want.
+However, the writer holds more responsibility in steering the LLM appropriately; P7-lab reflected, “As a CS student, I like this way more [than the treatment’s way of supporting writing], but I think it needs more thinking to use it creatively.”
+5.8.2. Papers-to-Posts’ Planning Features
+Participants often noted the selectable paper bullet points and paragraphs as a helpful feature of Papers-to-Posts (lab: 10/20, deployment: 14/19).
+Several described the outline as helpful for selecting relevant content to generate text.
+For instance, P15-lab noted, “I could rewrite entire sections or direct it with just a few bullets.” Seven participants mentioned that the outline was useful even just for reviewing the paper’s content or structure.
+P4-lab commented, “The outline was really helpful in providing a quick overview of what content I had written.” For two lab participants, the bullet points provided transparency around the origins of the initial LLM draft and generations.
+P12-lab explained, “I liked the bullet point features to see what was being fed in vs not.” By facilitating review and selection of content to include in the blog post, the paper outline may have contributed to participants’ increased editing power.
+That said, participants also specified difficulties related to the paper outline.
+For instance, two lab participants were overwhelmed with the number of bullet points.
+P9-lab described the outline as “really bloated and hard to navigate.” Meanwhile, three lab participants were confused or encumbered by aspects of utilizing the bullet points.
+For example, P3-lab “felt stressed by choosing between the entire paragraph at hand and the (presumably) AI-generated summary bullets.”
+5.8.3. Papers-to-Posts’ Drafting and Revising Features
+Participants frequently mentioned the utility of Papers-to-Posts’s text modification abilities (lab: 12/20, deployment: 9/19).
+For example, P18-dep reflected, “The ‘MODIFY WRITING’ section is helpful.
+I like that the interface contains this particular section to support iterating on existing text.” All of the preset modification buttons (“simpler terms”, “condense”, “expand”, “more dramatic”, “less dramatic”) were specifically mentioned as helpful affordances at least once in the lab study alone.
+The most commonly noted one was the “condense” button.
+For example, P12-lab commented, “I sometimes put the modify text through multiple rounds of expand and contract, focusing on different parts.
+I found that helpful.” Three participants commented that they desired more preset modification buttons.
+For instance, P16-lab wanted a “more technical” button, while P14-dep wanted “more or less casual” options.
+P7-lab then added, “Or even let user to add their favorite rewrite button.” Interestingly, P2-lab found the preset modification buttons useful for understanding the tool’s capabilities: “It was nice to know what options the tool already knew of, and would perhaps have success with.
+For example, I would probably not have gone to a blank-box tool and said, ’please make the following more dramatic’.
+I would have stuck to more basic things like ‘please summarize the following’.” Several participants (lab: 5/20, deployment: 10/19) also noted the usefulness of at least one aspect of the customizable instructions for either generating or modifying text using Papers-to-Posts.
+P9-shared, “…being able to include explicit instructions for the model to generate text from was helpful in being able to control the information in the text that it generated.” Thus, Papers-to-Posts’s affordances for text drafting and revision may have contributed to participants’ increased editing power.
+Participants also discussed difficulties that they encountered in drafting and revising writing with Papers-to-Posts.
+Four lab participants expressed wanting to take specific actions that were not explicitly supported by Papers-to-Posts, contrasting with the baseline’s flexibility that participants appreciated.
+As an example, two wanted assistance with checking for redundancies in the blog post.
+The other two wanted feedback on the blog post writing.
+Regarding the affordances for modifying text, 8/20 lab and 5/19 deployment participants noted a difficulty.
+Four lab participants raised confusion about how or when to use certain affordances.
+As an example, P6-lab was unsure of “when to use which one of the six [modification] options.” Another recurrent issue noted by five lab participants was a modification affordance not working as desired.
+For P20-lab, setting the desired length of the modified writing to one paragraph still resulted in three, as the back-end prompts are not guaranteed to work as planned.
+For P12-lab, modifications were sometimes too exaggerated, while for P16-lab, a custom modification to make writing more technical was too subdued.
+A couple deployment participants also noted that modifying text sometimes introduced errors.
+Future work may address these points of restriction and confusion noted by the participants using Papers-to-Posts.
+6. Discussion
+Based on our results, we reflect on design implications for other mixed-initiative tools for research-paper summarization and long-form scientific communication.
 6.1. Post-Editing is Not Enough
-Both Papers-to-Posts and the baseline tool included initial LLM-generated drafts generated with powerful LLMs, GPT-4 and GPT-4-32k respectively.
-Thus, the existence of an initial LLM draft cannot be the primary reason why participants preferred the blog post generated with Papers-to-Posts as compared to the baseline tool.
-While the participants had the option to simply post-edit the initial draft and refrain from interacting further with the LLM, only two baseline and two treatment participants across the lab and deployment studies did so.
-This suggests that additional user interfaces and interaction designs can provide better scaffolding supports to further improve upon LLM-generated text compared to simple post-editing.
-7.2. Long-Document Outlines
-While blog post writing has additional goals and constraints compared to general document summarization, our work also has implications that may be useful for the task of general human-AI long document summarization.
-Specifically, a main challenge here is the high costs of human summarizer reviewing, locating, and utilizing different essential information in a long document.
-Prior work in human-LLM summarization has attempted to address this challenge in various manners, including presenting visualizations for the document to be summarized (Zhang et al., 2023b, 2011), dividing the document into chunks (Cheng et al., 2023), and providing document highlights for information already included in the summary (Norkute et al., 2021).
-However, to the best of our knowledge, no prior work has explored the use of generated outlines to help users review the document to be summarized.
-In this work, participants found it easier to find and incorporate essential information from the document using Papers-to-Posts and participants consistently noted the selectable bullet points as one of the most useful features.
-We found that, given this tool, participants were able to make more change to their initial drafts for a given amount of time or writing actions.
-We therefore note that presenting LLM-generated outlines of scientific long documents may make the planning step of writing their summaries easier.
-As participants faced confusion about the connection between the outline and the paper text, we also recommend working to make this connection as clear as possible.
-How detailed the outline should be remains uncertain.
+Both Papers-to-Posts and the baseline tool included initial LLM-generated drafts generated with powerful LLMs.
+In the lab study, they were GPT-4 and GPT-4-32k respectively; in the deployment study, both conditions used GPT-4-0125-preview.
+Nevertheless, while the participants had the option to simply post-edit the initial draft and refrain from interacting further with the LLM, only two baseline and two treatment participants across the lab and deployment studies did so.
+This suggests that writers may benefit from additional support other than an initial draft when writing research-paper blog posts.
+6.2. Increasing Editing Power with Plan-Draft-Revise Workflow
+With additional features to support planning, drafting, and revising the blog post, Papers-to-Posts increased writers’ editing power (i.e., change in the blog post per minute or writing action).
+One reason participants may have demonstrated increased editing power is that we provided an interactive research-paper outline to help them plan content to include in their blog post.
+Indeed, when using Papers-to-Posts versus the baseline tool, the median participant found it easier to incorporate information from the paper in the blog post.
+In their qualitative responses, participants also consistently noted the interactive paper outline as one of the most useful features.
+We therefore note that presenting LLM-generated interactive outlines of research papers shows promise in making the planning step of writing their summaries easier.
+Another reason participants may have demonstrated increased editing power is that we provided pre-filled yet customizable LLM instructions for generating and modifying the blog post text.
+In alignment, the median participant found it easier to iterate on the blog post while using Papers-to-Posts.
+Several participants commented on the usefulness of different aspects of Papers-to-Posts’s preset yet customizable LLM instructions for drafting and revising the blog post text.
+Thus, we observe that preset LLM instructions scaffolded for customization may support users in drafting and revising long-form scientific communication.
+Lowering the temporal cost and writing cost of making an edit, Papers-to-Posts facilitates writers’ iteration on their blog posts, which we saw may be helpful for creating a blog post that satisfies them in a time-constrained setting like the lab study.
+Iteration may also help the writers to reflect on the blog post text (Zhou and Sterman, 2024).
+On the other hand, participants often described the baseline as more flexible and allowing them more freedom.
+Thus, utilizing the two tools in combination may prove useful.
+Papers-to-Posts may be used when the writer desires suggestions for how to edit their blog post, and the baseline tool (or no LLM tool) could be used when the writer has a specific edit in mind or wants to give themselves creative space to contemplate their own edits (Buschek, 2024).
+7. Limitations and Future Work
+Most participants were PhD students in computer science, who may benefit from raising awareness of their early-career work through blog posts; future work may investigate how more senior researchers utilize a tool like Papers-to-Posts.
+Future studies can also examine how researchers in other fields, who are generally less familiar with LLMs and may have different traditional paper formats, may benefit from a tool like Papers-to-Posts.
+Furthermore, Papers-to-Posts was designed to support blog posts that summarize a paper, but there are a variety of scientific blog post formats that future work could support (e.g., narrative or opinion-based formats).
+Future work may additionally investigate how different aspects of the input paper such as length, recency, and author position affect the user experience and outcome.
+Papers-to-Posts itself has several limitations.
+For one, multiple deployment study participants commented on how the automatic paper parsing makes mistakes and misses parts of the paper text.
+The tool also does not support incorporating figures and tables in the blog post.
+In addition, future research could work to make the connection between the outline and paper text clearer (e.g., highlight blog post text to reveal related paper text).
+Moreover, how detailed the paper outline should be remains uncertain.
 While several participants appreciated the bullet points, a couple mentioned that the outline was overwhelming.
-As one participant proposed, a hierarchy of bullet points may be more useful.
-7.3. Transparent Affordances for Long-form Writing Sub-Tasks
-In contrast with shorter forms of scientific writing such as tweetorials (Long et al., 2023) and extended metaphors (Kim et al., 2023) explored in prior work on human-LLM scientific writing, longer scientific texts such as research-paper blog posts require completing a number of sub-tasks common to long-form science communication.
-For example, because research-paper blog posts are often for broader audiences than the paper itself, Papers-to-Posts includes a ”simpler terms” option to make large swaths of text more digestible for such readers.
-As another example, because different sections of a research-paper blog post may serve different purposes, we pre-populated the custom instructions for each of the initial draft sections (e.g., for the introduction section, we ask the tool to present a hook, a high-level description of the problem and why it is interesting, etc.).
-These transparent, section-specific instructions could then be verified or adjusted by the user.
-Several participants commented on the usefulness of Papers-to-Posts’s different affordances for streamlining different generation and modification sub-tasks such as these.
-Thus, we observe that similar tools may assist users by including affordances for sub-tasks related to generating and modifying long-form output.
-8. Limitations and Future Work
-The majority of the participants were PhD students in computer science.
-Although many of the takeaways discussed in this paper may generalize to other groups of researchers, it would be interesting to see how more diverse populations surface unique needs from the tool.
-For example, researchers in the field of computer science tend to be better versed in artificial intelligence than those in other fields, so they may not need as much guidance in utilizing an LLM writing tool.
-Different areas of research may also have very different traditional paper and blog post structures compared to computer science, which may benefit from different tool affordances.
-Furthermore, PhD students are prime candidates for writing blog posts, which can help them raise awareness of their work as they are starting their research careers.
-However, more senior researchers may desire specific features for writing their blog posts, such as a feature to coordinate discussion of multiple works or a feature to connect an older work to the current research landscape.
-We also leave it to future work to investigate how different aspects of the input paper such as length, recency, and author position may affect the support that users want from the tool.
+Lastly, future work may look into how writers’ expertise may be utilized in supporting their blog post writing; a tool for Q-and-A research-paper summarization, for example, utilizes the researcher’s background to personalize its output (Lim et al., 2024).
 8. Ethical Considerations
-There are a few ethical considerations that we would like to note regarding this work.
-For one, it is well-known that LLMs may hallucinate information included in their generated text (Bubeck et al., 2023).
-We mitigate this risk by focusing on researchers summarizing their own work, which they should know well; for example, multiple participants noted inaccuracies or exaggerations in generated text, but most were able to correct those issues.
-A couple participants raised concerns that they had not caught all the errors by the end of their session, but as one of them noted, with more time they could be thorough in ensuring that the text is correct.
-We believe that LLM usage will increase in the future, and it is important to study effective design.
-Another ethical concern is whether using an LLM tool to write research-paper blog posts could lead to the loss of important details from the paper.
-P3 in the baseline condition raised this point: ”…it seems like the easier you make it, i.e., the better the initial summary, the less likely I am to myself remember those things that happen to be missing.” That said, participants overall felt more confident that all essential information was covered in their blog posts with Papers-to-Posts, so this work provided one step in the direction of support for reducing missing information.
-Furthermore, an LLM tool for writing research-paper blog posts may contribute to making science more accessible in society.
-This may raise concerns of reducing paper readership.
-However, blog posts do not provide the same amount of information as papers and do not go through peer review, research papers still hold their own value, and blog posts can act as catalysts for paper reading.
-Finally, researchers may not acknowledge the use of AI in writing their blog posts, which could mislead readers, particularly if the researcher has not verified the blog post contents.
-For the participants who shared their blog posts, we found that many chose to disclose its use, but to ensure better transparency around its use, future versions of Papers-to-Posts may highlight the provenance of the output writing (e.g., by visualizing the writer’s interaction with the LLM-powered tool (Naimul Hoque et al., 2023)).
+It is well-known that LLMs may hallucinate information included in their generated text (Bubeck et al., 2023).
+We mitigate this risk by focusing on researchers summarizing their own work, which they should know well.
+Multiple participants identified inaccuracies or exaggerations in generated text, which they could then correct.
+However, a few participants raised concerns about catching errors, suggesting a need for better support in detecting them.
+Another ethical concern is whether using an LLM-powered tool to write paper blog posts could lead to the loss of important details.
+Reflecting on the baseline, P3-lab raised this point: “…it seems like the easier you make it, i.e., the better the initial summary, the less likely I am to myself remember those things that happen to be missing.” Nevertheless, the median participant found it easier to incorporate information from the paper with Papers-to-Posts, so this work provided one step in the direction of reducing loss of important details.
+Furthermore, while an LLM-powered tool for writing research-paper blog posts may contribute to making science more accessible, it may raise concerns about decreasing paper readership.
+However, blog posts neither provide the same amount of information nor go through peer review.
+Papers still hold their own value, and blog posts can act as catalysts for paper reading.
+Finally, researchers may not acknowledge the use of AI in writing their blog posts, which could mislead readers, particularly if the writer has not properly verified the content.
+Many participants who shared their blog posts chose to disclose its use, but to provide better transparency, future versions of Papers-to-Posts may highlight the provenance of the output writing (e.g., by visualizing the writer’s interaction with the tool (Naimul Hoque et al., 2023)).
 9. Conclusion
-We introduce Papers-to-Posts, an LLM-powered research-paper blog post writing tool, which effectively helps researchers plan, draft, and revise their blog posts, Through a formative study (N=6) with an LLM technology probe for this task, we identified the following challenges faced by writers: they 1) appreciated a paper outline for finding and selecting relevant information from the paper but faced difficulties in navigating it and 2) tried to provide recurrent instructions for different kinds of sub-tasks to the LLM without specific affordances to support these sub-tasks.
-We formulated design goals to address each of these challenges, namely to help writers 1) plan their writing by facilitating review of the long document’s contents and selection of essential information to include in the writing and 2) draft and revise their writing by streamlining commonly desired instructions to the LLM for generating and modifying text.
-We validated that our LLM-powered tool Papers-to-Posts addresses these goals and helps authors write blog posts for their own papers through two user studies comparing Papers-to-Posts against a baseline providing an LLM blog post draft and free-form LLM prompting.
-In a within-subjects, time-controlled lab study (N=20 participants), we found that Papers-to-Posts significantly increased writer satisfaction with the output blog post, without increasing task completion time or cognitive load.
-We also observe that Papers-to-Posts led to significantly greater change to the initial LLM draft blog post, indicating that this increased satisfaction may be due to the tool facilitating iteration on the blog post.
-A field deployment study (N=37 blog posts, 26 participants) showed that in a more realistic setting Papers-to-Posts again resulted in greater change in the blog post for a given amount of editing time or number of editing actions.
-We hope that this work may inform the design of future human-LLM tools for scientific long-document summarization into long-form output.
+We introduce Papers-to-Posts, an LLM-powered tool for writing research-paper blog posts that follows the Plan-Draft-Revise workflow.
+We validated that Papers-to-Posts supports these steps through two user studies comparing how researchers write about their own papers with Papers-to-Posts versus a strong baseline (LLM blog post draft and free-form LLM prompting).
+In a within-subjects lab study (N=20 participants), Papers-to-Posts significantly increased writers’ editing power (i.e., change in initial blog post within a given amount of time), significantly increased satisfaction with the output blog post, and did not increase cognitive load.
+In a more realistic setting, a deployment study (N=37 blog posts, 26 participants) also showed that Papers-to-Posts resulted in greater editing power in terms of both time and writing actions, without decreasing blog post satisfaction or increasing cognitive load.
+We hope that this work informs future research and tools related to research-paper summarization and long-form scientific communication.
 Acknowledgements.
@@
 - Burns et al. (2003) Terry W Burns, D John O’Connor, and Susan M Stocklmayer. 2003. Science communication: a contemporary definition. Public understanding of science 12, 2 (2003), 183–202.
+- Buschek (2024) Daniel Buschek.
+2024.
+Collage is the New Writing: Exploring the Fragmentation of Text and User Interfaces in AI Tools.
+In Proceedings of the 2024 ACM Designing Interactive Systems Conference.
+2719–2737.
 - Cai et al.
@@
 1–13.
+- Dhillon et al.
+(2024) Paramveer S Dhillon, Somayeh Molaei, Jiaqi Li, Maximilian Golub, Shaochun Zheng, and Lionel Peter Robert.
+2024.
+Shaping Human-AI Collaboration: Varied Scaffolding Levels in Co-writing with Language Models.
+In Proceedings of the CHI Conference on Human Factors in Computing Systems.
+1–18.
 - Ding et al.
@@
 - Gero et al. (2022) Katy Ilonka Gero, Vivian Liu, and Lydia Chilton. 2022. Sparks: Inspiration for science writing using language models. In Designing interactive systems conference. 1002–1019.
-- Ghazvininejad et al.
-(2017) Marjan Ghazvininejad, Xing Shi, Jay Priyadarshi, and Kevin Knight.
-2017.
-Hafez: an interactive poetry generation system.
-In Proceedings of ACL 2017, System Demonstrations.
-43–48.
 - Goldfarb-Tarrant et al.
@@
 arXiv preprint arXiv:1904.02357 (2019).
-- Google ([n. d.]) Google. [n. d.]. BARD. https://bard.google.com/chat.
+- Google ([n. d.]) Google. [n. d.]. Gemini. https://gemini.google.com/.
+- Google ([n. d.]) Google. [n. d.]. NotebookLM. https://notebooklm.google/.
 - Hart and Staveland (1988) Sandra G Hart and Lowell E Staveland.
@@
 arXiv preprint arXiv:2206.06383 (2022).
+- Lim et al.
+(2024) Hyunseung Lim, Ji Yong Cho, Taewan Kim, Jeongeon Park, Hyungyu Shin, Seulgi Choi, Sunghyun Park, Kyungjae Lee, Juho Kim, Moontae Lee, et al.
+2024.
+Co-Creating Question-and-Answer Style Articles with Large Language Models for Research Promotion.
+In Proceedings of the 2024 ACM Designing Interactive Systems Conference.
+975–994.
 - Long et al.
@@
 - Lu et al. (2023) Guang Lu, Sylvia B Larcher, and Tu Tran. 2023. Hybrid Long Document Summarization using C2F-FAR and ChatGPT: A Practical Study. arXiv preprint arXiv:2306.01169 (2023).
+- Lu et al.
+(2024) Zhuoran Lu, Sheshera Mysore, Tara Safavi, Jennifer Neville, Longqi Yang, and Mengting Wan.
+2024.
+Corporate Communication Companion (CCC): An LLM-empowered Writing Assistant for Workplace Social Media.
+arXiv preprint arXiv:2405.04656 (2024).
 - Luc et al.
@@
 Journal of science Communication 13, 03 (2014), A05.
-- Medium ([n. d.]) Medium. [n. d.]. Medium. https://blog.medium.com/best-practices-for-writing-on-medium-386506ae62b9.
 - Mehlenbacher (2019) Ashley Mehlenbacher. 2019. Science communication online: Engaging experts and publics on the internet. The Ohio State University Press.
@@
 arXiv e-prints (2023), arXiv–2311.
-- Norkute et al.
-(2021) Milda Norkute, Nadja Herger, Leszek Michalak, Andrew Mulder, and Sally Gao.
-2021.
-Towards explainable AI: Assessing the usefulness and impact of added explainability features in legal document summarization.
-In Extended Abstracts of the 2021 CHI Conference on Human Factors in Computing Systems.
-1–7.
 - NotionAI ([n. d.]) NotionAI. [n. d.]. NotionAI. https://www.notion.so/product.
-- OpenAI ([n. d.]) OpenAI. [n. d.]. chatGPT. https://chat.openai.com/.
 - Patrick Dunleavy ([n. d.]) Patrick Dunleavy. [n. d.]. LSE. https://blogs.lse.ac.uk/impactofsocialsciences/2016/01/25/how-to-write-a-blogpost-from-your-journal-article/.
+- Reza et al.
+(2024) Mohi Reza, Nathan M Laundry, Ilya Musabirov, Peter Dushniku, Zhi Yuan “Michael” Yu, Kashish Mittal, Tovi Grossman, Michael Liut, Anastasia Kuzminykh, and Joseph Jay Williams.
+2024.
+ABScribe: Rapid Exploration & Organization of Multiple Writing Variations in Human-AI Co-Writing Tasks using Large Language Models.
+In Proceedings of the CHI Conference on Human Factors in Computing Systems.
+1–18.
+- Shamsabadi and D’Souza (2024) Mahsa Shamsabadi and Jennifer D’Souza.
+2024.
+A fair and free prompt-based research assistant.
+In International Conference on Applications of Natural Language to Information Systems.
+Springer, 218–224.
 - Singh et al.
@@
 - UWaterloo ([n. d.]) UWaterloo. [n. d.]. UWaterloo. https://uwaterloo.ca/writing-and-communication-centre/writing-academic-blog-posts.
-- Van Someren et al. (1994) Maarten Van Someren, Yvonne F Barnard, and J Sandberg. 1994. The think aloud method: a practical approach to modelling cognitive. London: AcademicPress 11, 6 (1994).
 - Wang et al.
@@
 1–30.
+- Zhou and Sterman (2024) David Zhou and Sarah Sterman.
+2024.
+Ai.llude: Investigating Rewriting AI-Generated Text to Support Creative Expression.
+In Proceedings of the 16th Conference on Creativity & Cognition.
+241–254.
 - Zou and Hyland (2019) Hang Zou and Ken Hyland. 2019. Reworking research: Interactions in academic articles and blogs. Discourse studies 21, 6 (2019), 713–733.
-Appendix A Prompts
-A.1. Prompt for Initial LLM Draft for Baseline Tool
-Figure 10 shows the prompt for the initial LLM draft for the baseline tool.
-A.2. Prompts for Initial LLM Draft for Papers-to-Posts
-A.3. Prompts for Working with Papers-to-Posts
-Appendix B Additional Log Analysis Plots
-Figures 25 and 26 show additional log analysis plots. Figure 27 shows the distribution of behavioral actions taken.
-Appendix C Formative Study Technology Probe
-Figure 28 shows the formative study technology probe.
+Appendix A Exploratory Analysis of Perceived Control and Ownership
+We analyzed participants’ Likert-type responses to questions related to control and ownership (Figure 9).
+A.1. Lab Study
+The median participant did not indicate higher perceived control (treatment: M=5.00, Q1=4.50, Q3=6.00; baseline: M=5.00, Q1=3.00, Q3=6.00) or satisfaction with their control (treatment: M=5.00, Q1=2.75, Q3=6.00; baseline: M=5.00, Q1=3.00, Q3=5.25) under the treatment condition (Figure 9a-b).
+However, if we look within-subjects, 11 participants had higher perceived control with Papers-to-Posts than the baseline tool (equal: 2, lower: 7), and 10 participants had higher satisfaction with their control using Papers-to-Posts (equal: 4, lower: 6).
+On the other hand, participants did not demonstrate increased perceived ownership (treatment: M=6.00, Q1=5.75, Q3=6.00; baseline: M=6.00, Q1=5.00, Q3=6.00) or satisfaction with ownership (treatment: M=6.00, Q1=4.00, Q3=6.25; baseline: M=6.00, Q1=5.00, Q3=6.00) under the treatment condition (Figure 9c-d).
+Although they felt that they had more control over Papers-to-Posts, participants may not necessarily have felt that the tool contributed more or less to the blog post’s creation.
+A.2. Deployment Study
+As shown in Figure 9e-h, participants indicated higher median satisfaction with control in the treatment condition (treatment: M=5.00, Q1=3.00, Q3=6.00; baseline: M=3.00, Q1=3.00, Q3=4.75), but they showed no difference in median perceived control (treatment: M=5.00, Q1=5.00, Q3=6.00; baseline: M=5.00, Q1=3.00, Q3=5.00), perceived ownership (treatment: M=6.00, Q1=4.50, Q3=6.00; baseline: M=6.00, Q1=6.00, Q3=6.00) or satisfaction with ownership (treatment: M=6.00, Q1=4.00, Q3=6.00; baseline: M=6.00, Q1=5.00, Q3=6.00) in the treatment condition.
+A.3. Summary
+While lab study participants had higher median perceived control, deployment study participants did not.
+However, we saw some evidence in both studies of higher perceived satisfaction with control under the treatment condition.
+Participants may have been more satisfied with their control without perceiving a higher control in the sense that they were more comfortable with letting the tool have a certain amount of control in exchange for more editing power.
+Across the two studies, we did not see a change in median perceived ownership or satisfaction with ownership.
+Appendix B Prompts
+B.1. Prompt for Initial LLM Draft for Baseline Tool
+B.2. Prompts for Start-Up Step of Papers-to-Posts
+B.3. Prompts for Drafting Step of Papers-to-Posts
+B.4. Prompts for Revising Step of Papers-to-Posts
+Appendix C Additional Log Analysis Plots
+Figures 25 and 26 show additional log analysis plots for the lab study, while Figure 27 shows additional log analysis plots for the deployment study.
 Appendix D Updated Tool for Deployment Study
 Figure 28 shows the updated tool for the deployment study.
+Appendix E Example Baseline Initial Draft
+Below is example output from the baseline tool for the paper “ConceptEVA: Concept-Based Interactive Exploration and Customization of Document Summaries” (Zhang et al., 2023b).
+Note that this output is for the deployment study version of Papers-to-Posts, which used automatic paper parsing without manual fixes.
+This example was generated in October 2024.
+E.1. Introduction
+In the realm of academic research, distilling the essence of long, multi-topic documents into concise summaries is a formidable challenge.
+This task becomes even more daunting when considering the diverse backgrounds of readers who may seek to understand these documents.
+Traditional summarization techniques, both extractive and abstractive, often fall short in addressing the nuances of such complex texts.
+Recognizing this gap, our research introduces ConceptEVA, a novel system designed to revolutionize the way we approach the summarization of academic papers.
+ConceptEVA leverages the latest advancements in natural language processing (NLP) and artificial intelligence (AI) to offer a mixed-initiative approach, enabling users to generate, evaluate, and customize summaries tailored to their specific interests.
+This system not only promises to enhance the accessibility of academic literature but also to enrich the reader’s engagement with the material.
+E.2. Methods
+ConceptEVA is built upon a multi-task Longformer Encoder Decoder (LED), fine-tuned to adeptly handle the summarization of lengthy, multi-domain documents.
+This innovative system employs attention mechanisms to efficiently process extensive texts while maintaining high information fidelity.
+A key feature of ConceptEVA is its interactive visualization of document concepts, presented as a network that illustrates semantic relatedness and co-occurrence.
+Users can interact with this visualization to select concepts of interest, which ConceptEVA then uses to automatically update the summary, emphasizing the chosen concepts.
+Our evaluation of ConceptEVA involved two iterations: an initial expert review and a subsequent within-subjects study, comparing ConceptEVA’s performance against manually-generated summaries.
+E.3. Results
+The findings from our study indicate a clear preference among participants for ConceptEVA’s customized summaries over their manually-generated counterparts.
+Participants appreciated the ability to tailor summaries to their interests, resulting in higher satisfaction levels with the summaries produced by ConceptEVA.
+However, incorporating critique into summaries remained a challenge, highlighting an area for future improvement.
+Despite this, the overall response to ConceptEVA was positive, with users finding the system helpful for exploring and verifying concepts within documents, especially those outside their domain of expertise.
+E.4. Conclusion
+ConceptEVA represents a significant step forward in the field of document summarization, particularly for academic literature that spans multiple topics and domains.
+By integrating interactive visual analysis with advanced NLP techniques, ConceptEVA offers a user-friendly platform for generating, evaluating, and customizing summaries.
+This mixed-initiative approach not only improves the summarization process but also enhances the user’s engagement with the document.
+While challenges remain, particularly in integrating critique into summaries, the positive feedback from our study participants underscores ConceptEVA’s potential as a valuable tool for researchers, educators, and anyone seeking to navigate the complexities of academic literature.
+Future developments will focus on refining ConceptEVA’s capabilities, further bridging the gap between comprehensive document analysis and user-centric summarization.
+Appendix F Example Papers-to-Posts Output
+Below is example output from Papers-to-Posts for the paper “ConceptEVA: Concept-Based Interactive Exploration and Customization of Document Summaries” (Zhang et al., 2023b).
+Note that this output is for the deployment study version of Papers-to-Posts, which used automatic paper parsing without manual fixes.
+These examples were generated in October 2024.
+F.1. Initial Draft
+F.1.1. Introduction
+In the rapidly evolving field of natural language processing and artificial intelligence, the challenge of effectively summarizing long, complex documents that span multiple topics remains a significant hurdle.
+This is particularly true for academic papers, which often contain dense, multi-disciplinary content that can be difficult to distill into a concise summary.
+Traditional methods of extractive summarization, which rely on pulling key sentences directly from the text, fall short in capturing the nuanced essence of such documents.
+The advent of abstractive summarization, which paraphrases the original text to produce a more human-like summary, offers a promising solution but is still plagued by challenges, especially when dealing with documents that cover a broad range of knowledge domains.
+Enter ConceptEVA, a novel mixed-initiative approach designed to tackle the complexities of summarizing long, multi-topic academic documents.
+By integrating advanced machine learning techniques with interactive visual analysis, ConceptEVA allows users to generate, evaluate, and customize summaries in a way that is both efficient and tailored to their specific interests.
+This system represents a significant leap forward in automated text summarization, offering a more dynamic and user-centric solution to the problem of digesting extensive academic literature.
+The significance of ConceptEVA lies not only in its ability to produce concise summaries of complex documents but also in its interactive nature, which empowers users to steer the summarization process.
+This human-in-the-loop approach ensures that the generated summaries are relevant and valuable to the reader, addressing the long-standing challenge of creating universally applicable summaries for documents that traverse multiple domains.
+As we continue to push the boundaries of what’s possible in automated text summarization, ConceptEVA stands out as a beacon of innovation, promising to transform how we engage with and understand academic literature.
+F.1.2. Methods
+In developing ConceptEVA, we employed a multi-faceted approach to address the challenges of summarizing long, multi-topic documents.
+At the core of ConceptEVA is a custom multi-task Longformer encoder-decoder, specifically designed for the summarization of extensive documents.
+This model leverages the power of attention mechanisms from transformers to efficiently manage memory usage while ensuring high fidelity in information preservation across longer texts.
+The attention mechanisms operate at both local and global levels, facilitating a nuanced understanding and processing of the document’s content.
+To enhance user interaction and customization, ConceptEVA integrates interactive visualizations of document concepts as a network.
+This visualization reflects semantic relatedness and co-occurrence, allowing users to focus on and select concepts of interest.
+These selected concepts can then automatically update the summary to emphasize them, supported by a ”focus-on” function that surfaces and highlights related concepts.
+This interactive component is underpinned by a multi-domain ontology for concept identification and a force-directed layout for visual representation, ensuring a user-friendly interface for customizing summaries.
+The iterative development of ConceptEVA was informed by an initial survey of eight research practitioners and underwent two stages of refinement based on expert reviews and a within-subjects study.
+This process included fine-tuning a pre-trained Longformer encoder-decoder for scientific document summarization, incorporating paraphrasing and semantic sentence embedding techniques, and developing an interactive summary editor to facilitate better human-AI collaboration.
+Additionally, we utilized Faiss, a similarity search library for dense vectors, to enhance the summarizer’s context understanding, thereby improving the quality and relevance of the generated summaries.
+F.1.3. Results
+Our study revealed that ConceptEVA significantly enhances user satisfaction with the summarization process, particularly when compared to manually-generated summaries.
+Specifically, 8 out of 12 participants preferred the summaries generated using ConceptEVA’s human-in-the-loop approach over the baseline, indicating a clear preference for the customized summaries provided by our system.
+This preference was especially pronounced among participants reviewing documents outside their domain of expertise, underscoring ConceptEVA’s utility in helping users explore unfamiliar topics through efficient and focused summarization.
+Moreover, the integration of ConceptEVA’s ”focus-on” function, which allows users to select and emphasize specific concepts of interest, was instrumental in achieving higher satisfaction levels.
+This feature not only facilitated a more interactive and personalized summarization experience but also highlighted the system’s capability to adapt summaries based on user input.
+However, it was noted that participants with less experience in the subject matter found it challenging to incorporate critique, such as limitations and implications, into their summaries, suggesting an area for further improvement in guiding users through more nuanced aspects of summarization.
+A significant association was found between the summarization approach employed by ConceptEVA and the summary satisfaction rating, as evidenced by a chi-squared test of independence.
+This statistical validation, along with user feedback, underscores the effectiveness of ConceptEVA in enhancing the summarization process through mixed-initiative interactions and customized content, paving the way for future developments in automated text summarization technologies.
+F.1.4. Conclusion
+In conclusion, ConceptEVA represents a significant advancement in the field of automated text summarization, particularly for academic documents.
+By leveraging a human-in-the-loop approach, ConceptEVA not only enhances the relevance and quality of summaries but also fosters a more collaborative interaction between users and AI.
+This system’s ability to integrate interactive visual analysis with natural language processing techniques allows for a more nuanced and customizable summarization process, addressing the critical need for summaries that cater to individual user interests and requirements.
+The positive feedback from our user study, especially regarding the system’s content-specific assistance and the preference for ConceptEVA-generated summaries over manual efforts, underscores the potential of this approach to significantly improve how we engage with and digest complex academic literature.
+Looking forward, we aim to expand ConceptEVA’s capabilities by testing it on a broader range of documents and incorporating additional features based on user feedback.
+This will include enhancing the system’s flexibility to adapt summaries more dynamically to user priorities and further refining the AI’s ability to maintain logical and narrative coherence throughout the summary.
+By continuing to focus on the synergy between human expertise and AI efficiency, ConceptEVA is poised to set new standards in the automation of document summarization, making it an invaluable tool for researchers, scholars, and anyone looking to efficiently extract insights from extensive academic writings.
+F.2. Initial Draft Bullet Point Selection
+F.2.1. Introduction Section’s Selected Bullet Points
+ConceptEVA is a mixed-initiative approach for summarizing long and multi-topic documents.
+CHI 2023 conference in Hamburg, Germany featured a multi-disciplinary research paper with an auto-generated summary
+Automated text summarization aims to present key concepts in a concise manner and reduce repetition.
+Challenges in abstractive summarization include summarizing long, complex documents spanning multiple knowledge domains.
+A human-in-the-loop approach is needed for interactive generation of relevant summaries.
+ConceptEVA integrates visual analysis and NLP for evaluating and customizing long document summaries.
+ConceptEVA introduces a human-in-the loop, mixed-initiative approach to evaluate and customize document summary generation
+Our work focuses on interactively exploring concept focus in generated summaries and emphasizing concepts of interest in academic publications.
+ConceptEVA customizes summaries by updating them with user-selected concepts of interest.
+ConceptEVA is an interactive document summarization system for long, multi-domain documents.
+F.2.2. Methods Section’s Selected Bullet Points
+It incorporates a custom multi-task longformer encoder decoder for summarizing longer documents.
+Interactive visualizations of document concepts as a network in ConceptEVA help users focus on concepts of interest
+Attention mechanisms from transformers used at local levels to reduce memory usage
+ConceptEVA design informed by initial survey of eight research practitioners and refined through two stages of development and evaluation
+ConceptEVA includes fine-tuning an LED model, identifying concepts using an ontology, and providing interactive visualization for customization.
+Maintain the user's mental map of the original document by preserving its layout
+We use Faiss to implement the approach, a similarity search library for dense vectors in large scale.
+Extended ConceptEVA with an interactive summary editor for better human-AI collaboration
+We constructed a set of guidelines for participants to follow when generating a summary manually or using ConceptEVA.
+The second iteration includes a longformer encoder-decoder pre-trained for scientific documents, fine-tuned for paraphrasing and sentence embedding, and concepts visualized using a force-directed network.
+F.2.3. Results Section’s Selected Bullet Points
+Participants' satisfaction with ConceptEVA customized summaries is higher than their own manually-generated summary
+ConceptEVA is helpful for examining and verifying ideas
+ConceptEVA is more useful for participants outside their domain of interest
+Inexperienced participants found it difficult to incorporate critique into the summary
+Text embeddings are pre-computed for each sentence in the source document and retrieved when users select concepts.
+8 out of 12 participants rated the summary generated using ConceptEVA's human-in-the-loop approach (task T3) higher than the baseline
+Study used 6 participants to summarize ConceptScope and 6 participants to summarize BodyVis
+A significant association was found between summarization approach and summary satisfaction rating
+Higher rating for human-in-the-loop approach led to more efficient location and focus on concepts
+Participants preferred summaries created through ConceptEVA's human-in-the-loop approach
+F.2.4. Conclusion Section’s Selected Bullet Points
+ConceptEVA is a mixed-initiative system for academic document readers and writers to generate, evaluate, and customize automated summaries.
+ConceptEVA integrates visual analysis and NLP for evaluating and customizing long document summaries.
+ConceptEVA's approach better exploits human-AI collaboration and iteratively improves the summary
+Saliency, fluency, novelty, and coherence are key factors in evaluating automated summarization techniques
+The technique should be flexible enough to change the summaries based on the priority of the user.
+Research shows that regular exercise can improve mental health by reducing symptoms of depression and anxiety.
+The AI-generated summarization approach prioritizes content over the flow of the summary, compromising logical and narrative connections between sentences
+Participants examining their own summaries can reveal insights into the effectiveness of the human-in-the-loop approach.
+We will eliminate the limitation by testing ConceptEVA on more diverse papers in the future
+ConceptEVA was found helpful for content-specific aspects of summarization
+F.3. Example Modifications
+We present modifications of the first sentence in the initial draft, using the preset modification macros and default settings for the LLM instructions (except we toggled off “Check if want modified writing to be based on selected paragraphs and bullet points”).
+F.3.1. Input Text to be Modified
+In the rapidly evolving field of natural language processing and artificial intelligence, the challenge of effectively summarizing long, complex documents that span multiple topics remains a significant hurdle.
+F.3.2. Expanded Version
+In the rapidly evolving field of natural language processing (NLP) and artificial intelligence (AI), the challenge of effectively summarizing long, complex documents that span multiple topics remains a significant hurdle.
+This is particularly true for academic papers, which often contain dense, multi-disciplinary content that can be difficult to distill into a concise summary.
+Traditional summarization techniques, which may work well for shorter, single-topic documents, often fall short when applied to these more complex texts.
+F.3.3. Condensed Version
+In the fast-paced domains of natural language processing and artificial intelligence, summarizing lengthy, multi-topic documents is a major challenge.
+F.3.4. Simpler-Terms Version
+In the fast-changing world of computer programs that understand and use human language, finding a way to briefly explain long, complicated texts that cover many subjects is still a big challenge.
+F.3.5. Less Dramatic Version
+In the field of natural language processing and artificial intelligence, summarizing lengthy, multi-topic documents is a notable challenge.
+F.3.6. More Dramatic Version
+Title: Tackling the Complexity of Summarizing Multi-Topic Documents in AI
+In the dynamic world of AI and natural language processing, we’re constantly facing the tough challenge of summarizing lengthy, intricate documents covering various topics.